### v1

# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

Abstract
We consider maximal operators acting on vector valued functions, that is functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts.
1 Introduction
This paper deals with matrix weighted extensions of the Fefferman–Stein vector valued inequalities for the Hardy–Littlewood maximal function and the closely connected topic of matrix weighted extensions of the product BMO spaces. Before stating our main results, we briefly recall some historical background on each area.
The classical Fefferman–Stein vector valued inequalities, first proved by C. Fefferman and E. Stein [Fefferman_Stein_1971], state that for all $1<p,q<\infty$ one has
$\left\|\left(\sum_{k=1}^{\infty}|Mf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}\leq C(n,p,q)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})},$ (1.1)
for any sequence $\{f_{k}\}_{k=1}^{\infty}$ of (say) locally integrable complex valued functions on $\mathbb{R}^{n},$ where $M$ is the usual (uncentered) Hardy–Littlewood maximal function. That means
$Mf(x)\coloneq\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|f(x)|\,\mathrm{dm}(x),\quad x\in\mathbb{R}^{n},$
where the supremum ranges over all cubes $Q$ in $\mathbb{R}^{n}$ (with faces parallel to the coordinate hyperplanes), $|E|$ is the Lebesgue measure of a Lebesgue measurable set $E\subseteq\mathbb{R}^{n},$ and $\,\mathrm{dm}(x)$ denotes integration with respect to Lebesgue measure. The constant $C(n,p,q)$ in (1.1) depends only on $n,$ $p$ and $q.$ In [Fefferman_Stein_1971] also a weak type version of (1.1) is proved.
The estimate (1.1) is actually a special case of more general bounds for vector valued extensions of operators. That is, given some (not necessarily linear) operator $T$ acting boundedly on $\mathrm{L}^{p}(\mathbb{R}^{n})$ for some $1<p<\infty,$ we seek to find those $1<q<\infty$ satisfying an estimate of the form
$\left\|\left(\sum_{k=1}^{\infty}|Tf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}\leq C(T,n,p,q)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}$ (1.2)
for any sequence $\{f_{k}\}^{\infty}_{k=1}.$ Such inequalities seem to have been studied for the first time systematically by J.-L. Rubio de Francia [Rubio_de_Francia_1985]. For a thorough modern exposition of the methods in [Rubio_de_Francia_1985] we refer to [Tao_notes_5_247A] as well as [Grafakos_Book_1].
The ideas in [Rubio_de_Francia_1985] already hinted at an intimate connection between extrapolation and bounds for vector valued extensions as in (1.2). By an extrapolation problem one understands the following. Given an operator $T$ acting on (suitable) functions on $\mathbb{R},$ we assume that for some $1<p<\infty$ it is already known or given that for all weights (that means, a.e. positive locally integrable functions) $w$ on $\mathbb{R}^{n}$ that belong to some class $C(p),$ one has the estimate $\|Tf\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,w)\|f\|_{\mathrm{L}^{p}(w)}.$ Given this information, find all $1<q<\infty$ and as well as associated classes $C(q)$ of weights on $\mathbb{R}^{n},$ such that for any $w\in C(q)$ one has an estimate of the form $\|Tf\|_{\mathrm{L}^{q}(w)}\leq C(T,n,q,w)\|f\|_{\mathrm{L}^{q}(w)}.$ Rubio de Francia [Rubio_de_Francia_1984] solved completely the extrapolation problem in the case that $C(p)$ coincides with the Muckenhoupt $A_{p}$ class, that is $w\in A_{p}$ if and only if
$[w]_{A_{p}}\coloneq\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}w(x)\,\mathrm{dm}(x)\right)\left(\frac{1}{|Q|}\int_{Q}w(x)^{-1/(p-1)}\,\mathrm{dm}(x)\right)^{p-1}<\infty,$
where the supremum ranges again over all cubes $Q\subseteq\mathbb{R}^{n}.$ In this case, beginning with any fixed $1<p<\infty$ and an extrapolation hypothesis holding for all weights $w\in A_{p},$ the extrapolation problem is solvable for any $1<q<\infty$ and all weights $w\in A_{q}.$ The extrapolation theorem of Rubio de Francia was subsequently further refined by various authors, until a sharp quantitative version of it was proved by J. Duoandikoetxea [Duoandikoetxea_2011]. A very thorough treatment of various forms of extrapolation with extensive historical background can be found in [Book_Extrapolation]. In fact, this method is so powerful, that it naturally yields weighted estimates for vector valued extensions, that is
$\left\|\left(\sum_{k=1}^{\infty}|Tf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,q,w)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(w)},$ (1.3)
as explained in detail in [Book_Extrapolation].
Inequalities of the form (1.3) are a major tool for estimating operators arising naturally when decomposing biparameter operators or bicommutators in simpler, localized pieces. Such a decomposition for the so called Journeé operators was established by H. Martikainen [Martikainen_2012] (generalizing an analogous decomposition proved earlier by T. Hytönen [Hytonen_2012] for Calderón–Zygmound operators in the context of the solution of the $A_{2}$ problem). I. Holmes, S. Petermichl and B. Wick [Holmes_Petermichl_Wick_2018] showed that these localized pieces can be estimated in terms of a weighted product BMO space. In the following, we recall the relevant definitions.
The classical space $\text{BMO}(\mathbb{R}^{n})$ consists of all locally integrable functions $b$ on $\mathbb{R}^{n}$ such that
$\|b\|_{\text{BMO}(\mathbb{R}^{n})}\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}|b(x)-\langle b\rangle_{Q}|\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all cubes $Q\subseteq\mathbb{R}^{n},$ and we denote $\langle b\rangle_{Q}\coloneq\frac{1}{|Q|}\int_{Q}b(x)\,\mathrm{dm}(x).$ The importance of this space is two-fold. First, it is the dual space to the (real variable) Hardy space $\mathrm{H}^{1}(\mathbb{R}^{n}).$ Second, the norm $\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}$ is the “correct” quantity controlling the boundedness of commutators $[T,b]=[T,M_{b}],$ where $M_{b}$ denotes (pointwise) multiplication by $b$ (called symbol of the commutator) and $T$ is a Calderón–Zygmound operator. This was proved for the Hilbert transform by Z. Nehari [Nehari_1957] and in full generality by R. R. Coifmann, Rochberg and G. Weiss [Coifman_Rochberg_Weiss_1976]. Moreover, the John–Nirenberg inequalities are an important intrinsic property of the space $\mathrm{BMO}(\mathbb{R}^{n}).$
B. Muckenhoupt and R. L. Wheeden [Muckenhoupt_Wheeden_BMO_1976] considered and studied the weighted BMO norm
$\|b\|_{\mathrm{BMO}(\nu)}\coloneq\sup_{Q}\frac{1}{\nu(Q)}\int_{Q}|b(x)-\langle b\rangle_{Q}|\,\mathrm{dm}(x),$ (1.4)
where the supremum ranges over all cubes $Q\subseteq\mathbb{R}^{n}$ and $\nu$ is a $A_{2}$ weight on $\mathbb{R}^{n}.$ A characterization of the two weighted boundedness of commutators in terms of a weighted BMO norm of the symbol was established by S. Bloom [Bloom_1985] for the Hilbert transform and later for arbitrary Calderón–Zygmund operators by I. Holmes, M. Lacey and B. Wick [Holmes_Lacey_Wick_2016]. In the latter work two weighted versions of (1.4) were introduced and associated John–Nirenberg inequalities were established. These played an important role in the commutator estimates in [Holmes_Lacey_Wick_2016].
The study of biparameter BMO spaces on product spaces $\mathbb{R}^{n}\times\mathbb{R}^{m}$ was initiated by S. Y. A. Chang [Chang_1979] and R. Fefferman [RFefferman_1979]. Here, “biparameter” refers to invariance of the considered function spaces under rescaling each coordinate variable of the domain of definition separately. Works [Chang_1979] and [RFefferman_1979] introduced and investigated the biparameter product BMO space $\text{BMO}(\mathbb{R}\times\mathbb{R})$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{2}$ (considered as the product space $\mathbb{R}\times\mathbb{R}$) such that
$\|b\|_{\text{BMO}(\mathbb{R}\times\mathbb{R})}\coloneq\sup_{\Omega}\bigg{(}\frac{1}{|\Omega|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ R\subseteq\Omega\end{subarray}}\left|(b,w_{R})\right|^{2}\bigg{)}^{1/2}<\infty,$
where the supremum reanges over all (say) bounded Borel subsets $\Omega$ of $\mathbb{R}^{2}$ with nonzero measure, $\bm{\mathcal{D}}$ stands for the family of all dyadic rectangles of $\mathbb{R}^{2},$ and $(w_{R})_{R\in\bm{\mathcal{D}}}$ is some (mildly regular) wavelet system adapted to $\bm{\mathcal{D}}.$ Here and below we denote
$(b,w_{R})\coloneq\int_{\mathbb{R}^{2}}b(x)w_{R}(x)\,\mathrm{dm}(x).$
The aforementioned works [Chang_1979] and [RFefferman_1979] established in particular that $\text{BMO}(\mathbb{R}\times\mathbb{R})$ is the dual to the biparameter Hardy space $\mathrm{H}^{1}(\mathbb{R}\times\mathbb{R}).$ Moreover, a dyadic version of this product BMO space is the correct space for characterizing the boundedness of bicommutators $[T_{1},[T_{2},b]],$ where $T_{1},T_{2}$ are Haar multipliers, as proved by [Blasco_Pott_2005].
A weighted version of the Chang–Fefferman product BMO space was introduced and studied by Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018] in the context of proving two weight upper bounds for biparameter praproducts. Namely, given a biparameter dyadic grid $\bm{\mathcal{D}}$ in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and a biparameter $\bm{\mathcal{D}}$-dyadic $A_{2}$ weight $w$ on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ [Holmes_Petermichl_Wick_2018] considers the dyadic Bloom type product space $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with the property
$\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)}\coloneq\sup_{\Omega}\bigg{(}\frac{1}{\nu(\Omega)}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|b_{R}^{\varepsilon}|^{2}\langle\nu^{-1}\rangle_{R}\bigg{)}^{1/2}<\infty,$ (1.5)
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ of nonzero finite measure. We refer to Section 2 for a detailed explanation of the notation and the terminology. In [Holmes_Petermichl_Wick_2018] an $\mathrm{H}^{1}$-BMO duality type result was established in this setting, which played a crucial role in the proofs of the upper bounds there. More recently, a two-weight version of (1.5) was defined in [Kakaroumpas_Soler_2022] and associated John–Nirenberg inequalities were established. These played an important role in [Kakaroumpas_Soler_2022] for characterizing the two weight boundedness of bicommutators with Haar multipliers, extending the aforementioned result of [Blasco_Pott_2005] to the two weight setting.
1.1 Main results
One of the main goals of this paper is to prove matrix weighted bounds for vector valued extensions of the Christ–Goldberg maximal function, which can be understood as a matrix weighted extension of the classical Fefferman–Stein vector valued inequalities for the Hardy–Littlewood maximal function.
Theorem 1.
Let $1<p<\infty.$ Consider a $(d\times d)$ matrix $A_{p}$ weight $W$ and a sequence of vector valued functions $\{f_{n}\}^{\infty}_{n=1}.$ Then, for each $1<q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}},$
and $M_{W}$ denotes the Christ–Goldberg maximal function corresponding to the weight $W$ and the exponent $p.$
We refer to Subsection 3.1 for a detailed explanation of the notation in Theorem 1. Theorem 1 yields readily a similar estimate for the so called modified Christ–Goldberg maximal function, as explained in Theorem 7 below.
We deduce Theorem 1 from a general principle for establishing matrix weighted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 5 below. Our method for deducing such bounds is inspired from [Book_Extrapolation]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [Cruz_Uribe_Bownik_Extrapolation], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the very recent limited range extrapolation theorem for matrix weights proved in [Kakaroumpas_Nguyen_Vardakis_2024] yields similar bounds as in Theorem 5 that are valid only for a limited range of exponents.
Theorem 1 and its counterpart for the modified Christ–Goldberg maximal function, Theorem 7, allow one to complete the proof of matrix weighted bounds for general (not necessarily paraproduct free) Journé operators given in [DKPS2024] for any $1<p<\infty$ (not necessarily $p=2$), as explained in [DKPS2024, Section 8] (thus avoiding the use of extrapolation for biparameter matrix weights as in [Vuorinen_2023]). In this paper we focus on the application of Theorem 1 for setting up the foundations of a theory of two matrix weighted product BMO. Namely, let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2},$
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure. This definition is an extension of one of the equivalent definitions for the space of two matrix weighted one-parameter BMO, whose study was initiated in [Isralowitz_Kwon_Pott2017], [Isralowitz_2017] and culminated in [Isralowitz_Pott_Treil_2022]. Moreover, for every sequence $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C}),$ we define
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\coloneq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{1}(\mathbb{R}^{n+m})},$
which is the direct biparameter analog of the one parameter two matrix weighted $\mathrm{H}^{1}$ norm from [Isralowitz_2017]. In this context, we prove the following result.
Theorem 2.
Let $U,V,p,\bm{\mathcal{D}}$ be as above. For any $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p),$ the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C}$ given by
$\ell_{B}(\Phi)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$
is well-defined and bounded with $\|\ell\|\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$ where the implied constants depend only on $n,m,p,d$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Conversely, for every bounded linear functional $\ell$ on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ there is $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with $\ell=\ell_{B}.$
We note that our proof of Theorem 2 trivially works also in the one-parameter setting, thus answering to the positive the question posed in [Isralowitz_2017] about the extension of the one parameter two matrix weighted $\mathrm{H}^{1}$-BMO duality proved there from $p=2$ to arbitrary exponents $1<p<\infty.$
The duality result of Theorem 2 coupled with Theorem 1 yields readily two-matrix weighted bounds for biparameter paraproducts, see Proposition 11 and Proposition 14 below. These yield in turn already some two matrix weighted upper bounds for bicommutators, see Subsection 4.3 below.
Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper.
2 Background
In this section we review the central definitions for our work. We also recall some basic facts about the various objects we deal with, in most cases without proof. References with further details are provided in this case.
2.1 Dyadic grids
For definiteness, in what follows intervals in $\mathbb{R}$ will always be assumed to be left-closed, right-open and bounded. A cube in $\mathbb{R}^{n}$ will be a set of the form $Q=I_{1}\times\ldots\times I_{n}$, where $I_{k},\leavevmode\nobreak\ k=1,\ldots,n$ are intervals in $\mathbb{R}$ of the same length, which we denote by $\ell(Q)\coloneq|I_{1}|.$ A rectangle in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ (with sides parallel to the coordinate axes) will be a set of the form $R=R_{1}\times R_{2}$, where $R_{1}$ is a cube in $\mathbb{R}^{n}$ and $R_{2}$ is a cube in $\mathbb{R}^{m}$.
A collection $\mathcal{D}$ of intervals in $\mathbb{R}$ will be said to be a dyadic grid in $\mathbb{R}$ if one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, such that the following hold:
  • for all $k\in\mathbb{Z}$, $\mathcal{D}_{k}$ forms a partition of $\mathbb{R}$, and all intervals in $\mathcal{D}_{k}$ have the same length $2^{-k}.$
  • for all $k\in\mathbb{Z}$, every $J\in\mathcal{D}_{k}$ can be written as a union of exactly $2$ intervals in $\mathcal{D}_{k+1}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}$ if
$\mathcal{D}\coloneq\left\{[m2^{k},(m+1)2^{k}):\leavevmode\nobreak\ k,m\in\mathbb{Z}\right\}.$
A collection $\mathcal{D}$ of cubes in $\mathbb{R}^{n}$ will be said to be a dyadic grid in $\mathbb{R}^{n}$ if for some dyadic grids $\mathcal{D}^{1},\ldots,\mathcal{D}^{n}$ in $\mathbb{R}$ one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, where
$\mathcal{D}_{k}=\{I_{1}\times\ldots\times I_{n}:\leavevmode\nobreak\ I_{i}\in\mathcal{D}^{i},\leavevmode\nobreak\ |I_{i}|=2^{-k},\leavevmode\nobreak\ i=1,\ldots,n\}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}^{n}$ if
$\mathcal{D}\coloneq\{[m_{1}2^{k},(m_{1}+1)2^{k})\times\dots\times[m_{n}2^{k},(m_{n}+1)2^{k}):\leavevmode\nobreak\ k,m_{1},\ldots,m_{n}\in\mathbb{Z}\}.$
If $\mathcal{D}$ is a dyadic grid in $\mathbb{R}^{n}$, then we denote
$\mathrm{ch}_{i}(Q)\coloneq\{K\in\mathcal{D}:\leavevmode\nobreak\ K\subseteq Q,\leavevmode\nobreak\ |K|=2^{-in}|Q|\},\qquad Q\in\mathcal{D},\leavevmode\nobreak\ i=0,1,2,\ldots.$
A collection $\bm{\mathcal{D}}$ is said to be a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ if for a dyadic grid $\mathcal{D}^{1}$ in $\mathbb{R}^{n}$ and a dyadic grid $\mathcal{D}^{2}$ in $\mathbb{R}^{m}$ we have
$\bm{\mathcal{D}}\coloneq\{R_{1}\times R_{2}:\leavevmode\nobreak\ R_{i}\in\mathcal{D}_{i},\leavevmode\nobreak\ i=1,2\},$
and in this case we write (slightly abusing the notation) $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$.
If $\bm{\mathcal{D}}$ is a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then we denote
$\mathrm{ch}_{i}(R)\coloneq\{Q_{1}\times Q_{2}:\leavevmode\nobreak\ Q_{i}\in\mathrm{ch}_{i_{j}}(R_{j}),\leavevmode\nobreak\ j=1,2\},$
where
$R\in\bm{\mathcal{D}},\leavevmode\nobreak\ i=(i_{1},i_{2}),\leavevmode\nobreak\ i_{1},i_{2}=0,1,2,\ldots.$
We say that $R$ is the $i$-th ancestor of $P$ in $\bm{\mathcal{D}}$ if $P\in\mathrm{ch}_{i}(R)$.
2.2 Haar systems
2.2.1 Haar system on $\mathbb{R}$
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}$. For any interval $I\in\mathcal{D}$ we denote by $h^{0}_{I},h^{1}_{I}$ the $\mathrm{L}^{2}$-normalized cancellative and noncancellative respectively Haar functions over the interval $I\in\mathcal{D}$, that is
$h^{0}_{I}\coloneq\frac{\mathbf{1}_{I_{+}}-\mathbf{1}_{I_{-}}}{\sqrt{|I|}},\qquad h^{1}_{I}\coloneq\frac{\mathbf{1}_{I}}{\sqrt{|I|}}$
(so $h^{0}_{I}$ has mean 0). For simplicity we denote $h_{I}\coloneq h_{I}^{0}$. For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R})$, we denote $f_{I}\coloneq(f,h_{I})$, $I\in\mathcal{D}$. It is well-known that the system $\{h_{I}\}_{I\in\mathcal{D}}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R})$. Of course, all these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.2 Haar system on $\mathbb{R}^{n}$
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}^{n}$. We denote
$\mathcal{E}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of one-parameter signatures. For a cube $Q=I_{1}\times\dots\times I_{n}\in\mathcal{D}$ and $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{n})\in\mathcal{E},$ we denote by $h^{\varepsilon}_{Q}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the cube $Q$ corresponding to the signature $\varepsilon$ defined by
$h^{\varepsilon}_{Q}(x)\coloneq h^{\varepsilon_{1}}_{I_{1}}(x_{1})\dots h^{\varepsilon_{n}}_{I_{n}}(x_{n}),\qquad x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}.$
For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n})$ we denote $f^{\varepsilon}_{Q}\coloneq(f,h^{\varepsilon}_{Q})$. It is well-known that the system $\{h^{\varepsilon}_{Q}\colon\leavevmode\nobreak\ Q\in\mathcal{D},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n})$. All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.3 Haar system on the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ We denote $\mathcal{E}\coloneq\mathcal{E}^{1}\times\mathcal{E}^{2}$, where
$\mathcal{E}^{1}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\},\leavevmode\nobreak\ \mathcal{E}^{2}\coloneq\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of biparameter signatures. For $R=R_{1}\times R_{2}\in\bm{\mathcal{D}}$ and $\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\mathcal{E},$ we denote by $h_{R}^{\varepsilon}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the rectangle $R$ corresponding to the signature $\varepsilon$ defined by
$h_{R}^{\varepsilon}=h_{R_{1}}^{\varepsilon_{1}}\otimes h_{R_{2}}^{\varepsilon_{2}},$
that is
$h_{R}^{\varepsilon}(x_{1},x_{2})=h_{I}^{\varepsilon_{1}}(x_{1})h_{J}^{\varepsilon_{2}}(x_{2}),\qquad(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$
For any function $f\in\mathrm{L}_{\mathrm{loc}}^{1}(\mathbb{R}^{n+m})$ we denote $f_{R}^{\varepsilon}\coloneq(f,h_{R}^{\varepsilon})$. From the corresponding one-parameter facts, we immediately deduce that the system $\{h_{R}^{\varepsilon}\colon R\in\bm{\mathcal{D}},\varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n+m}).$ For $P\in\mathcal{D}^{1},$ $Q\in\mathcal{D}^{2}$, $\varepsilon_{1}\in\mathcal{E}^{1}$ and $\varepsilon_{2}\in\mathcal{E}^{2}$ we denote
$f^{\varepsilon_{1},1}_{P}(x_{2})\coloneq(f(\,\cdot\,,x_{2}),h_{P}^{\varepsilon_{1}}),\quad x_{2}\in\mathbb{R}^{m},$
$f^{\varepsilon_{2},2}_{Q}(x_{1})\coloneq(f(x_{1},\,\cdot\,),h_{Q}^{\varepsilon_{2}}),\quad x_{1}\in\mathbb{R}^{n}.$
All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions entrywise.
Finally, we remark that the dimensions $n,m$ will always be clear from the context, as well as whether $\mathcal{E}$ refers to the set of one-parameter or biparameter signatures.
2.3 Matrix norms in terms of column norms
In the sequel we denote by $\{e_{1},\ldots,e_{d}\}$ the standard basis of $\mathbb{C}^{d}.$ We will be often using the fact that
$|A|\sim_{d}\sum_{k=1}^{d}|Ae_{k}|,\quad\forall A\in\mathrm{M}_{d}(\mathbb{C}),$ (2.1)
without explicitly mentioning it. Note also that for all $0<p<\infty$ and for nonnegative numbers $x_{1},\ldots,x_{d}$ we have the estimate
$\min(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}\leq\left(\sum_{i=1}^{d}x_{i}\right)^{p}\leq\max(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}.$
In particular, if $F_{k}:\mathbb{R}^{n}\rightarrow\mathrm{M}_{d}(\mathbb{C})$ is a sequence of Lebesgue-measurable functions and $0<p,q,r<\infty,$ then we have
$\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\left(\sum_{i=1}^{d}|F_{k}(x)e_{i}|\right)^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\sum_{i=1}^{d}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q}\left(\int_{\mathbb{R}^{n}}\sum_{i=1}^{d}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,q}\sum_{i=1}^{d}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}.$ (2.2)
2.4 Matrix weighted matrix-valued Lebesgue spaces
A function $W$ on $\mathbb{R}^{n}$ is said to be a $d\times d$-matrix valued weight, or just matrix weight, if it is a locally integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function such that $W(x)$ is a positive-definite matrix for a.e. $x\in\mathbb{R}^{n}.$ Here, by locally integrable matrix valued function $W(x)$ we mean that the scalar valued function $|W(x)|,$ that is the matrix norm, is locally integrable.
Given a $d\times d$-matrix weight $W$ on $\mathbb{R}^{n}$ and $1<p<\infty,$ we define the norm
$\|F\|_{\mathrm{L}^{p}(W)}\coloneq\left(\int_{\mathbb{R}^{n}}|W(x)^{1/p}F(x)|^{p}\,\mathrm{dm}(x)\right)^{1/p},$
for all $\mathrm{M}_{d}(\mathbb{C})$-valued measurable functions $F$ on $\mathbb{R}^{n}.$ This norm defines the matrix weighted matrix-valued Lebesgue space $\mathrm{L}^{p}(W).$
2.5 Reducing operators
Let $1<p<\infty$. Let $E$ be a bounded measurable subset of $\mathbb{R}^{n}$ of nonzero measure. Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ that is integrable over $E$ (meaning that its matrix norm is integrable, so $\int_{E}|W(x)|\,\mathrm{dm}(x)<\infty$) and that takes a.e. values in the set of positive-definite $(d\times d)$ matrices. It is proved in [Goldberg_2003, Proposition 1.2] that there exists a (not necessarily unique) positive-definite matrix $\mathcal{W}_{E}\in\mathrm{M}_{d}(\mathbb{C})$, called reducing operator of $W$ over $E$ with respect to the exponent $p$, such that
$\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}\leq|\mathcal{W}_{E}\,e|\leq\sqrt{d}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}$
for all $e\in\mathbb{C}^{d}.$ If $d=1$, i.e. $W$ is scalar-valued, then one can clearly take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/p},$ where we denote the average of $W$ over $E$ by
$\langle W\rangle_{E}=\frac{1}{|E|}\int_{E}W(x)\,\mathrm{dm}(x).$
Moreover, if $p=2$, then
$\frac{1}{|E|}\int_{E}|W(x)^{1/2}e|^{2}\,\mathrm{dm}(x)=\frac{1}{|E|}\int_{E}\langle W(x)e,e\rangle\,\mathrm{dm}(x),$
which by linearity is equal to
$\langle\langle W\rangle_{E}\,e,e\rangle=|\langle W\rangle_{E}^{1/2}e|^{2}$
for all $e\in\mathbb{C}^{d}.$ Thus in this special case one can take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/2}.$
Assume in addition now that the function $W^{\prime}\coloneq W^{-1/(p-1)}$ is also integrable over $E$. Then we let $\mathcal{W}^{\prime}_{E}$ be the reducing matrix of $W^{\prime}$ over $E$ corresponding to the exponent $p^{\prime}\coloneq p/(p-1)$, so that
$|\mathcal{W}^{\prime}_{E}\,e|\sim_{d}\left(\frac{1}{|E|}\int_{E}|W^{\prime}(x)^{1/p^{\prime}}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}=\left(\frac{1}{|E|}\int_{E}|W(x)^{-1/p}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}$
for all $e\in\mathbb{C}^{d}.$ Note that one can take $\mathcal{W}^{\prime\prime}_{E}=\mathcal{W}_{E}$. Observe that
$|\mathcal{W}_{E}\mathcal{W}^{\prime}_{E}|\sim_{p,d}\left(\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p}.$
For a detailed exposition of reducing operators we refer for example to [DKPS2024]. Here we just state the following estimates for later convenience.
Lemma 3.
Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ taking a.e. positive-definite values, and such that $W$ and $W^{\prime}=W^{-1/(p-1)}$ are integrable over $E$ for some $1<p<\infty.$ Set
$C_{E}\coloneq\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x).$
We consider reducing operators of $W$ with respect to exponent $p$ and of $W^{\prime}$ with respect to exponent $p^{\prime}.$
  1. (1)
    There holds
    $|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{1/p}{\langle|W^{1/p}e|\rangle}_{E}$
    and
    $|\mathcal{W}_{E}^{-1}e|\leq|\mathcal{W}_{E}^{\prime}e|\lesssim_{p,d}C_{E}^{1/p}|\mathcal{W}_{E}^{-1}e|,$
    for all $e\in\mathbb{C}^{d}.$
  2. (2)
    There holds
    $|{\langle W^{1/p}\rangle}_{E}e|\leq|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{d/p}|{\langle W^{1/p}\rangle}_{E}e|,$
    for all $e\in\mathbb{C}^{d}.$
A proof of part (1) can be found, for example, in [DKPS2024]. A proof of part (2) can be found in [Isralowitz_Kwon_Pott2017].
Let $E,F$ be measurable subsets of $\mathbb{R}^{n},\mathbb{R}^{m}$ respectively with $0<|E|,|F|<\infty.$ Let $1<p<\infty.$ Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued integrable function on $E\times F$ taking a.e. values in the set of positive-definite $d\times d$-matrices. For all $x_{1}\in E,$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in F.$ For a.e. $x_{1}\in E,$ denote by $\mathcal{W}_{x_{1},F}$ the reducing operator of $W_{x_{1}}$ over $F$ with respect to the exponent $p.$ It is proved in [DKPS2024] (see also [Cruz_Uribe_Bownik_Extrapolation]) that one can choose the reducing operator $\mathcal{W}_{x_{1},F}$ in a way that is measurable in $x_{1}.$
Set $W_{F}(x_{1})\coloneq\mathcal{W}_{x_{1},F}^{p},$ for a.e. $x_{1}\in E.$ Then $W_{F}\in\mathrm{L}^{1}(E;\mathrm{M}_{d}(\mathbb{C})).$ It is proved in [DKPS2024] that
$|\mathcal{W}_{F,E}e|\sim_{p,d}|\mathcal{W}_{E\times F}e|,\qquad\forall e\in\mathbb{C}^{d},$ (2.3)
where $\mathcal{W}_{F,E}$ is the reducing operator of $W_{F}$ over $E$ with respect to the exponent $p,$ and $\mathcal{W}_{E\times F}$ is the reducing operator of $W$ over $E\times F$ with respect to the exponent $p.$
2.6 Matrix $A_{p}$ weights
2.6.1 One-parameter matrix $A_{p}$ weights
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}$. We say that $W$ is a one-parameter $d\times d$-matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n})}\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all cubes $Q$ in $\mathbb{R}^{n}$. Note that if $W$ is a $(d\times d)$ matrix valued $A_{p}$ weight on $\mathbb{R}^{n}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n})}^{1/p^{\prime}}\sim_{p,d}[W]_{A_{p}(\mathbb{R}^{n})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n})}\sim_{p,d}\sup_{Q}|\mathcal{W}^{\prime}_{Q}\mathcal{W}_{Q}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\mathcal{D}$ is any dyadic grid in $\mathbb{R}^{n}$, we define
$[W]_{A_{p},\mathcal{D}}\coloneq\sup_{Q\in\mathcal{D}}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x),$
and we say that $W$ is a one-parameter $(d\times d)$ matrix valued $\mathcal{D}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\mathcal{D}}<\infty$.
2.6.2 Biparameter matrix $A_{p}$ weights
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$. We say that $W$ is a biparameter $(d\times d)$ matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\coloneq\sup_{R}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all rectangles $R$ in $\mathbb{R}^{n}$ (with sides parallel to the coordinate axes). Note that if $W$ is a $(d\times d)$ matrix valued biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued biparameter $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p^{\prime}}\sim_{d,p}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\sim_{p,d}\sup_{R}|\mathcal{W}^{\prime}_{R}\mathcal{W}_{R}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\bm{\mathcal{D}}$ is any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, we define
$[W]_{A_{p},\bm{\mathcal{D}}}\coloneq\sup_{R\in\bm{\mathcal{D}}}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x);$
we say that $W$ is a biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\bm{\mathcal{D}}}<\infty$.
2.6.3 One-parameter restriction of biparameter matrix $A_{p}$ weights
Let $1<p<\infty.$ Let $W$ be a $(d\times d)$ matrix biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}.$ It is proved in [DKPS2024] that
$[W_{x_{1}}]_{A_{p}(\mathbb{R}^{m})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
Fix any cube $Q$ in $\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ let $\mathcal{W}_{x_{1},Q}$ be the reducing operator of $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}$ over $Q$ with respect to the exponent $p.$ Set $W_{Q}(x_{1})\coloneq\mathcal{W}_{x_{1},Q}^{p},$ for a.e. $x_{1}\in\mathbb{R}^{n}.$ It is shown in [DKPS2024] that
$[W_{Q}]_{A_{p}(\mathbb{R}^{n})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$ (2.4)
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
2.7 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite, we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [Cruz_Uribe_Bownik_Extrapolation, Cruz_Uribe_Extrapolation] and for a detailed exposition see [Aubin_2009].
Since through the article we will consider both functions taking values in $\mathbb{C}^{d}$ and functions taking values in $\mathcal{K}_{\mathrm{bcs}},$ we will use a typographic convention to avoid confusion. We will denote functions taking values in $\mathbb{C}^{d}$ with lowercase letters $f,g,h,\ldots$ On the other hand, we will use uppercase letters $F,G,H,\ldots$ to denote functions taking values in $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ In any case, we will also explicitly state the target space of the functions we use.
Given a set $K\subseteq\mathbb{C}^{d},$ let us define its norm by
$|K|\coloneq\sup\{|v|\colon v\in K\}.$
For a matrix $A$ we will denote its usual norm (given by its largest singular value) by $|A|.$ The action of matrix weights on convex body valued functions is given by the next definition. Given a convex body $K$ and a positive definite matrix $A,$ define the product
$AK\coloneq\{Au\colon u\in K\}$
and observe that $AK$ will also be a convex body.
We will say that a function $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ is measurable if for every open set $U\subseteq\mathbb{C}^{d}$ it holds that the set
$F^{-1}(U)\coloneq\{x\in\mathbb{R}^{n}\colon F(x)\cap U\neq\emptyset\}$
is measurable (in the sense of Lebesgue). A convex body valued function $F$ is measurable if and only if there exists a sequence $\{f_{k}\}_{k\geq 1}$ of measurable functions $f_{k}\colon\mathbb{R}^{n}\rightarrow\mathbb{C}^{d}$ such that
$F(x)=\overline{\{f_{k}(x)\colon k\geq 1\}}$
for every $x\in\mathbb{R}^{n}$ (see [Aubin_2009, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9]).
Lemma 4.
Consider a measurable function $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ Then there exists $f\in S^{0}(F)$ such that
$|f(x)|=|F(x)|$
for all $x\in\mathbb{R}^{n}.$
In order to define integrals of convex body valued functions, consider first the set
$S^{1}(F)\coloneq\{f\in S^{0}(F)\colon f\in\mathrm{L}^{1}(\mathbb{R}^{n})\}$
of integrable selection functions for $F.$ The Aumann integral of $F$ is then defined as
$\int_{\mathbb{R}^{n}}F(x)\,\,\mathrm{dm}(x)=\left\{\int_{\mathbb{R}^{n}}f(x)\,\,\mathrm{dm}(x)\colon f\in S^{1}(F)\right\}.$
In this work, we will restrict to integrably bounded convex body valued functions, that is to functions $F$ such that $|F(x)|$ is integrable. In particular, in this case $S^{0}(F)=S^{1}(F).$ If $|F(x)|$ is only locally integrable, we will say that $F$ is locally integrably bounded. For further convenience, given a cube $Q\subseteq\mathbb{R}^{n},$ we define the averaging operator $A_{Q}$ by
$A_{Q}F(x)\coloneq\frac{\chi_{Q}(x)}{|Q|}\int_{\mathbb{R}^{n}}F(y)\,\,\mathrm{dm}(y),$
where $F$ is a locally integrably bounded convex body valued function.
Given $1\leq p<\infty,$ we define the Lebesgue space of convex body valued functions $\mathrm{L}^{p}(\mathbb{R}^{n},\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})),$ or just $\mathrm{L}^{p}$ when there is no ambiguity, as the set of functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that
$\left\|{F}\right\|_{{\mathrm{L}^{p}}}\coloneq\left(\int_{\mathbb{R}^{n}}|F(x)|^{p}\,\,\mathrm{dm}(x)\right)^{1/p}<\infty.$
The space of convex body valued functions $\mathrm{L}^{\infty}(\mathbb{R}^{n},\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})),$ or just $\mathrm{L}^{\infty},$ is defined as the set of convex body valued functions $F$ for which
$\left\|{F}\right\|_{{\mathrm{L}^{\infty}}}\coloneq\operatorname*{ess\,sup}\{|F(x)|\colon x\in\mathbb{R}^{n}\}<\infty.$
Given a $(d\times d)$-matrix weight $W,$ we define the weighted Lebesgue space $\mathrm{L}^{p}(W)$ of convex body valued functions as the set of functions $F$ for which $W(x)^{1/p}F(x)\in\mathrm{L}^{p}.$
3 Vector valued extensions of operators on convex body valued functions
The aim of this section is to prove vector valued estimates for convex body valued functions with matrix weights, as well as to apply it to vector valued estimates for the matrix weighted maximal operators for vector valued functions (see Subsection 3.1 for the definitions and details). This will be a consequence of an extrapolation result on spaces of functions with values on $\mathcal{K}_{\mathrm{bcs}}$ and it is a minor modification of the analogous result for matrix weights and vector valued functions due to Bownik and Cruz-Uribe (see [Cruz_Uribe_Bownik_Extrapolation, Section 9]). Before stating the extrapolation theorem, we introduce the concept of families of extrapolation pairs following the convention in [Book_Extrapolation]. A family of extrapolation pairs $\mathcal{F}$ is a collection of pairs $(F,G)$ of measurable functions taking values in $\mathcal{K}_{\mathrm{bcs}}$ such that neither $F$ nor $G$ is equal to $\{0\}$ almost everywhere. In addition, given such a family $\mathcal{F},$ we call each element $(F,G)\in\mathcal{F}$ an extrapolation pair. We are interested in inequalities of the form
$\left\|{F}\right\|_{{\mathrm{L}^{p}(W)}}\leq C\left\|{G}\right\|_{{\mathrm{L}^{p}(W)}},\qquad(F,G)\in\mathcal{F},$
in the sense that this holds for all pairs $(F,G)\in\mathcal{F}$ for which the left-hand side is finite and with the constant $C$ depending on the characteristic $[W]_{A_{p}}$ but not on the particular weight $W.$
Theorem A (Bownik, Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation]).
Consider a family of extrapolation pairs $\mathcal{F}.$ Suppose that for some $p_{0},$ $1\leq p_{0}\leq\infty,$ there exists an increasing function $C_{p_{0}}$ such that for every matrix weight $W\in A_{p_{0}}$ it holds that
$\left\|{F}\right\|_{{\mathrm{L}^{p_{0}}(W)}}\leq C_{p_{0}}([W]_{A_{p_{0}}})\left\|{G}\right\|_{{\mathrm{L}^{p_{0}}(W)}},\qquad(F,G)\in\mathcal{F}.$ (3.1)
Then, for all $1<p<\infty$ and for every matrix weight $W\in A_{p}$ it holds that
$\left\|{F}\right\|_{{\mathrm{L}^{p}(W)}}\leq C_{p}(p_{0},n,d,[W]_{A_{p}})\left\|{G}\right\|_{{\mathrm{L}^{p}(W)}},\qquad(F,G)\in\mathcal{F},$
where
$C_{p}(p_{0},n,d,[W]_{A_{p}})=C(p,p_{0})C_{p_{0}}\left(C(n,d,p,p_{0})[W]_{A_{p}}^{\max\left\{1,\frac{p_{0}-1}{p-1}\right\}}\right)$
if $p_{0}<\infty,$ and
$C_{p}(p_{0},n,d,[W]_{A_{p}})=C(p,p_{0})C_{p_{0}}\left(C(n,d,p,p_{0})[W]_{A_{p}}^{\frac{1}{p-1}}\right)$
if $p_{0}=\infty.$
The proof of this result is exactly the same as in [Cruz_Uribe_Bownik_Extrapolation, Theorem 9.1], up to the point that the calculations are done directly with the given convex body valued functions $F$ and $G,$ instead of taking $F(x)=\{\lambda f(x)\colon|\lambda|\leq 1\}$ and $G(x)=\{\lambda g(x)\colon|\lambda|\leq 1\}$ for given vector valued functions $f$ and $g$ (also note that in [Cruz_Uribe_Bownik_Extrapolation] $F(x)$ and $G(x)$ are segments with endpoints ±$\pm f(x)$ and ±$\pm g(x)$ because in that article real symmetric convex bodies are used, instead of complex symmetric). Since the proof in [Cruz_Uribe_Bownik_Extrapolation] follows closely that of the scalar case, we also refer the reader to [Book_Extrapolation, Chapter I.3] for a detailed exposition on the topic in the classical setting.
The main result of this section follows from Theorem A. This is a vector valued estimate for sequences $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ of a given family of extrapolation pairs $\mathcal{F}.$
Theorem 5 (Vector valued estimates for families of extrapolation pairs).
Consider a family of extrapolation pairs $\mathcal{F}.$ Suppose that for some $1\leq p_{0}\leq\infty$ there exists an increasing function $C_{p_{0}}$ such that for every matrix weight $W\in A_{p_{0}}$ inequalities (3.1) hold. Then, for any $p$ and $q,$ $1<p,q<\infty,$ and for every matrix weight $W\in A_{p}$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,p_{0},q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}G_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ and
$C(n,d,p_{0},p,q,[W]_{A_{p}})=$ $C(d,p,q)$
$C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q-1}\right\}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)$
if $p_{0}<\infty,$ and
$C(n,d,p_{0},p,q,[W]_{A_{p}})=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\frac{1}{q-1}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)$
if $p_{0}=\infty.$
Recall that one can define the Minkowski addition of $A,B\in\mathcal{K}_{\mathrm{bcs}}$ (and for subsets of $\mathbb{C}^{d}$ in general) by
$A+B\coloneq\{a+b\colon a\in A,b\in B\}.$
It is easy to check that if $A,B\in\mathcal{K}_{\mathrm{bcs}},$ then also $A+B\in\mathcal{K}_{\mathrm{bcs}}.$ One can extend the Minkowski addition of two convex body to that of $N$ convex bodies by induction, with $N\geq 2.$ We can also define the scalar multiplication, given $A\in\mathcal{K}_{\mathrm{bcs}}$ and $\lambda\in\mathbb{C},$ by
$\lambda A\coloneq\{\lambda a\colon a\in A\},$
and also $\lambda A\in\mathcal{K}_{\mathrm{bcs}}.$ Keep in mind that these operations do not define a vector space structure, since the Minkowski addition has no inverse.
Given $1\leq q<\infty$ and a sequence $\{K_{n}\}_{n=1}^{\infty}\subseteq\mathcal{K}_{\mathrm{bcs}}$ such that
$\sum_{n=1}^{\infty}|K_{n}|^{q}<\infty,$ (3.2)
we define the infinite $\ell^{q}$ Minkowski addition of $\{K_{n}\},$ denoted by $\Sigma_{q}(\{K_{n}\}),$ as
$\Sigma_{q}(\{K_{n}\})\coloneq\bigcup\left\{\sum_{n=1}^{\infty}a_{n}v_{n}\colon v_{n}\in K_{n}\text{ for }n\geq 1,\{a_{n}\}\in\ell^{q^{\prime}},\left\|{\{a_{n}\}}\right\|_{{\ell^{q^{\prime}}}}\leq 1\right\}.$ (3.3)
The sums in (3.3) are convergent due to the summability condition (3.2) and Hölder’s inequality. In particular, it holds that
$\left|\Sigma_{q}(\{K_{n}\})\right|\leq\left(\sum_{n=1}^{\infty}|K_{n}|^{q}\right)^{1/q}$
for any $1\leq q<\infty$ whenever the right-hand side of this expression is finite. Furthermore, given a $(d\times d)$ matrix $A,$ it also holds by linearity that
$A\Sigma_{q}(\{K_{n}\})=\Sigma_{q}(\{AK_{n}\}).$
For $1\leq q<\infty$ it also happens that $\Sigma_{q}(\{K_{n}\})$ is a convex symmetric bounded closed set, which is proved in next lemma.
Lemma 6.
Let $1\leq q<\infty$ and let $\{K_{m}\}^{\infty}_{m=1}$ be a sequence of sets in $\mathcal{K}_{\mathrm{bcs}}$ with
$\sum_{m=1}^{\infty}|K_{m}|^{q}<\infty.$
Then $K=\Sigma_{q}(\{K_{n}\})$ is a well-defined set in $\mathcal{K}_{\mathrm{bcs}}$ with
$|K|\sim_{d,q}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
Proof.
First of all, for all sequences $\{v_{m}\}_{m=1}^{\infty}$ with $v_{m}\in K_{m},$ for all $m=1,2,\ldots$ and for all complex numbers $a_{1},a_{2},\ldots$ with $\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\leq 1$ we have by Hölder’s inequality
$\sum_{m=1}^{\infty}|a_{m}v_{m}|\leq\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\left(\sum_{m=1}^{\infty}|v_{m}|^{q}\right)^{1/q}\leq\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}<\infty.$
This shows that $K$ is a well-defined bounded set with $|K|\leq\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
It is obvious that $K$ is symmetric. Moreover, a standard weak star compactness argument yields that $K$ is closed. Let us check that $K$ is convex. For any sequences $\{v_{m}\}_{m=1}^{\infty},$$\{u_{m}\}_{m=1}^{\infty}$ with $v_{m},u_{m}\in K_{m},$ for all $m=1,2,\ldots,$ any complex numbers $a_{1},a_{2},\ldots$ and $b_{1},b_{2},\ldots$ with $\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\leq 1$ and $\left\|{\{b_{m}\}}\right\|_{{q^{\prime}}}\leq 1,$ and any $t\in(0,1),$ we can write
$t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}$
$=\sum_{m=1}^{\infty}(t|a_{m}|+(1-t)|b_{m}|)$
$\cdot\left(\frac{t|a_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{a_{m}}{|a_{m}|}v_{m}+\frac{(1-t)|b_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{b_{m}}{|b_{m}|}u_{m}\right),$
with the obvious modifications if $a_{m}=0$ or $b_{m}=0$ for some $m\geq 1,$ thus it is clear that $t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}\in K.$
We will now show that
$|K|\gtrsim_{d,q}\frac{1}{\sqrt{d}}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
We can pick a sequence $\{v_{m}\}_{m=1}^{\infty}$ with $v_{m}\in K_{m}$ and $|v_{m}|\geq\frac{1}{2}|K_{m}|,$ for all $m=1,2,\ldots.$ It is clear that there exists a positive integer $N$ with
$\sum_{m=1}^{N}|v_{m}|^{q}\geq\frac{1}{2}\sum_{m=1}^{\infty}|v_{m}|^{q}.$
For any vector $x\in\mathbb{C}^{d},$ we denote by $x^{1},\ldots,x^{d}$ its coordinates. Notice that
$\sum_{m=1}^{N}|v_{m}|^{q}\sim_{d,q}\sum_{m=1}^{N}\sum_{j=1}^{d}|v_{m}^{j}|^{q}=\sum_{j=1}^{d}\sum_{m=1}^{N}|v_{m}^{j}|^{q},$
therefore there exists $j\in\{1,\ldots,d\}$ with
$\sum_{m=1}^{N}|v_{m}^{j}|^{q}\gtrsim_{d,q}\sum_{m=1}^{N}|v_{m}|^{q}.$
Clearly, one can find complex numbers $a_{1},\ldots,a_{N}$ with $\left\|{\{a_{m}\}_{m=1}^{N}}\right\|_{{q^{\prime}}}\leq 1$ such that
$\sum_{m=1}^{N}a_{m}v_{m}^{j}=\left(\sum_{m=1}^{N}|v_{m}^{j}|^{q}\right)^{1/q}.$
It follows that
$\left|\sum_{m=1}^{N}a_{m}v_{m}\right|$ $\geq\left|\sum_{m=1}^{N}a_{m}v_{m}^{j}\right|=\left(\sum_{m=1}^{N}|v_{m}^{j}|^{q}\right)^{1/q}$
$\gtrsim_{d,q}\left(\sum_{m=1}^{N}|v_{m}|^{q}\right)^{1/q}\gtrsim_{q}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
Since $\sum_{m=1}^{N}a_{m}v_{m}=\sum_{m=1}^{N}a_{m}v_{m}+\sum_{m=N+1}^{\infty}0\cdot 0\in K,$ the proof is complete. ∎
We are now ready to prove Theorem 5.
Proof of Theorem 5.
Fix first $1<q<\infty.$ We construct a new family $\mathcal{F}_{q}$ of convex body valued functions as follows. For each sequence $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ such that
$\sum_{n=1}^{\infty}|F_{n}(x)|^{q}<\infty\quad\text{ and }\quad\sum_{n=1}^{\infty}|G_{n}(x)|^{q}<\infty$ (3.4)
at almost every $x\in\mathbb{R}^{n},$ we define
$F(x)=\Sigma_{q}(\{F_{n}(x)\}),\qquad G(x)=\Sigma_{q}(\{G_{n}(x)\}).$
Observe that by lemma 6, both $F$ and $G$ are convex body valued functions with
$|W(x)^{1/p}F(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}F_{n}(x)|^{q}\right)^{1/q}$
and
$|W(x)^{1/p}G(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}G_{n}(x)|^{q}\right)^{1/q}$
at almost every $x\in\mathbb{R}^{n},$ for any $1<p<\infty.$ It is also easy to see that the convex body valued functions $F,G$ are measurable.
Next, note that the family of extrapolation pairs $\mathcal{F}_{q}$ that we have just constructed satisfies inequalities (3.1) with exponent $q$ for any matrix weight $W\in A_{q}.$ Indeed, for any given matrix weight $W\in A_{q},$ due to Lemma 6 it holds that
$\begin{split}\left\|{F}\right\|_{{\mathrm{L}^{q}(W)}}^{q}&=\int_{\mathbb{R}^{n}}|W^{1/q}F(x)|^{q}\,\,\mathrm{dm}(x)\\ &=\int_{\mathbb{R}^{n}}\left|\Sigma_{q}(\{W(x)^{1/q}F_{n}(x)\})\right|^{q}\,\,\mathrm{dm}(x)\\ &\leq C(d,q)\sum_{n=1}^{\infty}\int_{\mathbb{R}^{n}}|W(x)^{1/q}F_{n}(x)|^{q}\,\,\mathrm{dm}(x)\\ &\leq C_{q}([W]_{q})^{q}\sum_{n=1}^{\infty}\int_{\mathbb{R}^{n}}|W(x)^{1/q}G_{n}(x)|^{q}\,\,\mathrm{dm}(x)\\ &\leq C(d,q)C_{q}([W]_{q})^{q}\int_{\mathbb{R}^{n}}\left|\Sigma_{q}(\{W(x)^{1/q}G_{n}(x)\})\right|^{q}\,\,\mathrm{dm}(x)\\ &=C_{q}([W]_{q})^{q}\left\|{G}\right\|_{{\mathrm{L}^{q}(W)}}^{q}.\end{split}$
Here we have used Theorem A with the extrapolation paris $(F_{n},G_{n})\in\mathcal{F}$ and with exponent $q,$ so that we have
$C_{q}([W]_{A_{q}})=C(d,q,p_{0})C_{p_{0}}\left(C(n,d,q,p_{0})[W]_{A_{q}}^{\max\left\{1,\frac{p_{0}-1}{q-1}\right\}}\right).$
We have just seen that the family of extrapolation pairs $\mathcal{F}_{q}$ satisfies (3.1), therefore we can apply Theorem A and Lemma 6 to get that for any $1<p<\infty$ and any matrix weight $W\in A_{p}$ it holds that
$\begin{split}\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}F_{n}(x)|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}&\leq C(d,p)\left\|{W^{1/p}F}\right\|_{{\mathrm{L}^{p}}}\\ \leq C_{p,q}(&[W]_{A_{p}})\left\|{W^{1/p}G}\right\|_{{\mathrm{L}^{p}}}\\ \leq C_{p,q}(&[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}G_{n}(x)|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\end{split}$
for every sequence $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ satisfying (3.4) and where
$\begin{split}C_{p,q}([W]_{A_{p}})&=C(d,p)C_{p,q}\left(C(n,d,p,q)[W]_{A_{p}}^{\max\left\{1,\frac{q-1}{p-1}\right\}}\right)\\ &=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q-1}\right\}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)\end{split}$
if $p_{0}<\infty,$ and
$\begin{split}C_{p,q}([W]_{A_{p}})&=C(d,p)C_{p,q}\left(C(n,d,p,q)[W]_{A_{p}}^{\max\left\{1,\frac{q-1}{p-1}\right\}}\right)\\ &=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\frac{1}{q-1}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)\end{split}$
if $p_{0}=\infty,$ with the additional factor $C(d,p,q)$ being in both cases due to the various applications of Lemma 6. ∎
3.1 Fefferman–Stein vector valued inequalities for weighted maximal operators
Next, we give an application of Theorem 5. Fix $1<p<\infty.$ Given a $(d\times d)$ matrix valued weight $W$ on $\mathbb{R}^{n},$ one can define the pointwise matrix weighted maximal operator for vector valued functions by
$M_{W}f(x)\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}f(y)|\,\,\mathrm{dm}(y)\mathbf{1}_{Q}(x),$
where $f$ is a locally integrable function taking values on $\mathbb{C}^{d}$ and where the supremum is taken over all cubes with sides parallel to the coordinate axes. The maximal operator $M_{W}$ is also called Christ–Goldberg maximal operator, since it was studied by these two authors (see [Christ_Goldberg_2001] and [Goldberg_2003]). One can also define a weighted maximal operator with reducing operators as
$\widetilde{M}_{W}f(x)\coloneq\sup_{Q}\int_{Q}|\mathcal{W}_{Q}f(y)|\,\,\mathrm{dm}(y)\mathbf{1}_{Q}(x),$
where $f$ is a locally integrable function with values on $\mathbb{C}^{d},$ the supremum is taken over bounded cubes with sides parallel to the axes and where $\mathcal{W}_{Q}$ is the reducing operator of $W$ over $Q$ with exponent $p.$ We do not make the dependence on the exponent $p$ explicit since it will always be clear from the context.
Both weighted maximal operators play an important role in the theory of matrix weighted norm inequalities. In addition, whenever $W$ is a matrix $A_{p}$ weight, both maximal operators are bounded from weighted $\mathrm{L}^{p}(W)$ to unweighted $\mathrm{L}^{p}.$ Isralowitz and Moen [Isralowitz_Moen_2019] proved that
$\left\|{M_{W}}\right\|_{{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}}}\lesssim_{n,d,p}[W]_{A_{p}}^{1/(p-1)},$ (3.5)
while Isralowitz, Kwon and Pott [Isralowitz_Kwon_Pott2017] showed that
$\left\|{\widetilde{M}_{W}}\right\|_{{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}}}\lesssim_{n,d,p}[W]_{A_{p}}^{1/(p-1)}.$
We will use a trick due to Bownik and Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation] so as to use Theorem 5 (which is a statement about convex body valued functions) to show statements about $\mathbb{C}^{d}$ vector valued functions. In order to apply the previous results for convex body valued operators to these weighted maximal operators, we will also need to define the convex body valued analogue of $M_{W}.$ Both definitions will be related by the following correspondence between vector valued and convex body valued functions. Given a vector valued function $f,$ that is taking values on $\mathbb{C}^{d},$ we define the convex body valued function $F$ by
$F(x)\coloneq\{\lambda f(x)\colon|\lambda|\leq 1\}.$ (3.6)
By construction, $F$ takes values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and also $|F(x)|=|f(x)|.$ Thus, any estimate for operators depending on $|f(x)|$ can be studied through estimates for an analogous operator acting on $F(x)$ and depending as well on $|F(x)|$ uniquely. It is straightforward to check that, for a locally integrable vector valued function $f$ and a cube $Q,$ it holds that the averaging operator $A_{Q}$ applied to $F$ given by (3.6) can be computed as
$A_{Q}F(x)=\left\{\frac{1}{|Q|}\int_{Q}k(y)f(y)\,\,\mathrm{dm}(y)\colon k\in\mathrm{L}^{\infty},\left\|{k}\right\|_{{\mathrm{L}^{\infty}}}\leq 1\right\}\mathbf{1}_{Q}(x).$ (3.7)
For this reason, we will abuse notation and denote $A_{Q}f(x)=A_{Q}F(x).$ Also note that, by an appropriate choice of the function $k$ in (3.7), one can see that
$|A_{Q}f(x)|\sim_{d}\frac{1}{|Q|}\int_{Q}|f(y)|\,\,\mathrm{dm}(y),$ (3.8)
and more generally
$|A_{Q}F(x)|\sim_{d}\frac{1}{|Q|}\int_{Q}|F(y)|\,\,\mathrm{dm}(y)$
for convex body valued functions in general, which can be seen by choosing an appropriate selection function. For fixed $1<p<\infty$ and given a $(d\times d)$-matrix weight $W,$ the convex body valued analogue of $M_{W}$ acting on a locally integrably bounded convex body valued function $F$ is defined by
$M_{W}^{\mathcal{K}}F(x)\coloneq\overline{\mathrm{conv}\left(\bigcup_{Q}A_{Q}(W(x)^{1/p}F(x))\right)},$
where the union is taken over all bounded cubes with sides parallel to the axes. Moreover, given a locally integrable vector valued function $f,$ we denote $M_{W}^{\mathcal{K}}f(x)=M_{W}^{\mathcal{K}}F(x),$ where $F$ is the convex body valued function given by (3.6). In addition, if $W$ is a matrix weight and $f$ a vector valued function, linearity and (3.8) yield that
$|M_{W}^{\mathcal{K}}f(x)|=|W(x)^{1/p}M^{\mathcal{K}}f(x)|\sim_{d}M_{W}f(x),$ (3.9)
where $M^{\mathcal{K}}$ denotes the maximal operator weighted by the $(d\times d)$-identity matrix.
Theorem 1.
Consider a $(d\times d)$ matrix weight $W$ and a sequence of vector valued functions $\{f_{n}\}.$ Then, for each $1<p,q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Proof.
In order to apply Theorem 5, we need to construct a family $\mathcal{F}$ of extrapolation pairs for which the left-hand side of (3.1) is finite. To this end, we will restrict ourselves to vector valued functions $f\in\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ (compactly supported essentially bounded functions), and a density argument will yield the conclusion for $f\in\mathrm{L}^{p}(\mathbb{R}^{n};\mathbb{C}^{d}).$ For each vector valued function $f\in\mathrm{L}^{\infty}_{\mathrm{c}},$ we consider the extrapolation pair $(M^{\mathcal{K}}F,F),$ where $F$ is the convex body valued function defined by (3.6).
Consider the given $1<q<\infty.$ It is clear that the left-hand side of (3.1) is finite for every pair $(M^{\mathcal{K}}F,F)\in\mathcal{F}$ for $p_{0}=q,$ and that (3.1) is satisfied if we take $C_{p_{0}}(t)=C_{q}(t)=C(n,d,q)t^{\frac{1}{q-1}}.$ Indeed, because of (3.9), for any matrix $A_{q}$ weight $W$ we have that
$\left\|{M^{\mathcal{K}}F}\right\|_{{\mathrm{L}^{q}(W)}}\sim_{d}\left\|{M_{W}f}\right\|_{{\mathrm{L}^{q}}}\leq C(n,d,q)[W]_{A_{q}}^{1/(q-1)}\left\|{f}\right\|_{{\mathrm{L}^{q}(W)}}<\infty,$
where we have used estimate (3.5) and that $\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})\subseteq\mathrm{L}^{q}(\mathbb{R}^{n};\mathbb{C}^{d}).$
Now, Theorem 5 gives that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}M^{\mathcal{K}}F_{n}|^{q_{1}}\right)^{1/{q_{1}}}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p_{0},p,q_{1},[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q_{1}}\right)^{1/{q_{1}}}}\right\|_{{\mathrm{L}^{p}}}$ (3.10)
for any $1<p,q_{1}<\infty$ (with $q_{1}$ possibly different of $q=p_{0}$) and any matrix weight $W\in A_{p},$ where $\{(M^{\mathcal{K}}F_{n},F_{n})\}$ is a sequence in $\mathcal{F}$ and where
$C(n,d,p_{0},p,q_{1},[W]_{A_{p}})$
$=C(d,p)C_{p_{0}}\left(C(n,d,p_{0},p,q_{1})[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q_{1}-1}\right\}\max\left\{1,\frac{q_{1}-1}{p-1}\right\}\right)}\right).$
In particular, if we restrict our attention to $q_{1}=p_{0}=q,$ we can take
$C(n,d,q,p,[W]_{A_{p}})=C(n,d,q,p)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Also observe that for each sequence $\{(M^{\mathcal{K}}F_{n},F_{n})\}\subseteq\mathcal{F},$ we can choose a sequence $\{f_{n}\}\subseteq\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ such that $F_{n}$ is obtained from $f_{n}$ using (3.6) for every $n\geq 1.$ On one hand observe that using that $|W^{1/p}F_{n}|=|W^{1/p}f_{n}|$ for the sequence $\{f_{n}\}\subseteq\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ that we chose previously, we get that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}=\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$ (3.11)
On the other hand, using (3.9), we see that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}M^{\mathcal{K}}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\sim_{d}\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$ (3.12)
Finally, replacing (3.11) and (3.12) into (3.10), we get the desired vector valued estimates for the pointwise weighted maximal operator, as we wanted to show. ∎
The analogous result for the operator $\widetilde{M}_{W}$ is a consequence of Theorem 1.
Theorem 7.
Consider a $(d\times d)$ matrix weight $W$ and a sequence of vector valued functions $\{f_{n}\}.$ Then, for each $1<p,q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|\widetilde{M}_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\frac{1}{p}+\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Proof.
Just observe that
$\begin{split}\widetilde{M}_{W}f(x)&=\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|\mathcal{W}_{Q}f(y)|\,\,\mathrm{dm}(y)\\ &\sim_{d}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(z)^{1/p}f(y)|^{p}\,\,\mathrm{dm}(z)\right)^{1/p}\,\,\mathrm{dm}(y)\\ &\leq C(d,p)[W]_{A_{p}}^{1/p}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}\frac{1}{|Q|}\int_{Q}|W(z)^{1/p}f(y)|\,\,\mathrm{dm}(z)\,\,\mathrm{dm}(y)\\ &\leq C(d,p)[W]_{A_{p}}^{1/p}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}M_{W}f(z)\,\,\mathrm{dm}(z)\\ &=C(d,p)[W]_{A_{p}}^{1/p}M(M_{W}f)(x).\end{split}$
Therefore, we get that
$\left\|{\left(\sum_{n=1}^{\infty}|\widetilde{M}_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(d,p)[W]_{A_{p}}^{1/p}\left\|{\left(\sum_{n=1}^{\infty}|M(M_{W}f_{n})|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$
An application of the classical Fefferman–Stein vector valued inequalities for the maximal operator followed by the use of Proposition 1 yields the desired result. ∎
4 Two matrix weighted biparameter product BMO
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$
Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We emphasize that $\mathcal{E}$ stands for the set of biparameter signatures, $\mathcal{E}=\mathcal{E}^{1}\times\mathcal{E}^{2},$ where $\mathcal{E}^{1}=\{0,1\}^{n}\setminus\{(1,\ldots,1)\}$ and $\mathcal{E}^{2}=\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$ We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2},$
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure, and all reducing operators are taken with respect to exponent $p.$ Note that
$|\mathcal{V}_{R}P\mathcal{U}^{-1}_{R}|=|\mathcal{U}^{-1}_{R}P^{\ast}\mathcal{V}_{R}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P^{\ast}(\mathcal{V}^{\prime}_{R})^{-1}|,\qquad\forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall P\in\mathrm{M}_{d}(\mathbb{C}).$
Therefore
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sim_{p,d}\|B^{\ast}\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(V^{\prime},U^{\prime},p^{\prime})}.$
4.1 $\mathrm{H}^{1}$-BMO duality
The main goal of this subsection is to prove Theorem 2. We split the proof in Propositions 8 and 9 below, each treating one direction of Theorem 2.
We define $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ as the set of all sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ such that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\coloneq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}<\infty.$
This is the direct biparameter analog of the one-parameter two matrix weighted $\mathrm{H}^{1}$ norm defined in [Isralowitz_2017]. It is easy to check that $(\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),\|\,\cdot\,\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)})$ is a Banach space.
Proposition 8.
Let $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p).$ Then, the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C},$
$\ell_{B}(\Phi)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$
is well-defined bounded with $\|\ell_{B}\|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}.$
Proof.
We adapt the first half of the proof of [Isralowitz_2017, Theorem 1.3]. Let $\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ be arbitrary. We show that the sum
$(\Phi,B)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})$
converges absolutely with
$|(\Phi,B)|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}.$
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})|=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathrm{tr}(\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R})|$
$\lesssim_{d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}|\leq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}|$
$\lesssim_{p,d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|.$
Set now
$F\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}$
and define
$\Omega_{k}\coloneq\{F>2^{k}\},\qquad k\in\mathbb{Z},$
$\mathcal{B}_{k}\coloneq\left\{R\in\bm{\mathcal{D}}:\leavevmode\nobreak\ |R\cap\Omega_{k+1}|\leq\frac{1}{2}|R|<|R\cap\Omega_{k}|\right\},\qquad k\in\mathbb{Z},$
$\widetilde{\Omega}_{k}\coloneq\left\{M_{\bm{\mathcal{D}}}(\mathbf{1}_{\Omega_{k}})>\frac{1}{2}\right\},\qquad k\in\mathbb{Z}.$
Clearly $\Omega_{k}\subseteq\widetilde{\Omega}_{k}$ up to a set of zero measure, and in fact $|\widetilde{\Omega}_{k}|\sim|\Omega_{k}|,$ for all $k\in\mathbb{Z}.$ It is also obvious that $R\subseteq\widetilde{\Omega_{k}},$ for all $R\in\mathcal{B}_{k},$ for all $k\in\mathbb{Z}.$ If $R\in\bm{\mathcal{D}}$ with $\Phi_{R}^{\varepsilon}\neq 0$ for some $\varepsilon\in\mathcal{E}$ and $|R|\geq 2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow-\infty$ we deduce
$|R|\geq 2|R\cap\{F>0\}|=2|R|,$
since $F>0$ a. e. on $R,$ contradiction. If $R\in\bm{\mathcal{D}}$ with $|R|<2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow\infty$ we deduce
$|R|\leq 2|R\cap\{F=\infty\}|=0,$
since by assumption $F\in\mathrm{L}^{1}(\mathbb{R}^{n+m}),$ contradiction. It follows that for every $R\in\bm{\mathcal{D}}$ with $\Phi_{R}\neq 0$ for some $\varepsilon\in\mathcal{E}$ there exists $k\in\mathbb{Z}$ such that $R\in\mathcal{B}_{k}.$ Thus, we have
$|(\Phi,B)|$ $\lesssim_{p,d}\sum_{k\in\mathbb{Z}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|$
$\leq\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2}$
$\leq\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\text{sh}(\mathcal{B}_{k})|^{1/2}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\widetilde{\Omega}_{k}|^{1/2}.$
We will now prove that
$\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{4/p}2^{2k}|\widetilde{\Omega}_{k}|,\qquad\forall k\in\mathbb{Z}.$ (4.1)
This will be enough to conclude the proof, because assuming it we will get
$\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\widetilde{\Omega}_{k}|^{1/2}\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\widetilde{\Omega}_{k}|$
$\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\Omega_{k}|\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|F\|_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}.$
Fix now $k\in\mathbb{Z}.$ We begin by noticing that
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\leq 2^{2k+2}|\widetilde{\Omega}_{k}\setminus\Omega_{k+1}|\leq 2^{2k+2}|\widetilde{\Omega}_{k}|,$
by the definition of $\Omega_{k+1}.$ Moreover, we have
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\geq\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\,\mathrm{dm}(x)$
$=\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\,\mathrm{dm}(x)$
$\sim_{d}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\,\mathrm{dm}(x)$
$\geq\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{|R\setminus\Omega_{k+1}|}{|R|}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}.$
Let $v\in\mathbb{C}^{d}$ be arbitrary. Consider the function
$w\coloneq|V^{-1/p}v|^{p^{\prime}}.$
Then, $w$ is a scalar $\bm{\mathcal{D}}$-dyadic biparameter $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with
$[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}\lesssim_{p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{p^{\prime}/p},$
because $V^{-1/p}$ is a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight (see for example Lemma 3.2 in [DKPS2024]). It is then well-known that
$[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}$
and
$\langle w\rangle_{R}^{1/p^{\prime}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}\langle w^{1/p^{\prime}}\rangle_{R},\quad\forall R\in\bm{\mathcal{D}}$
(see for example Subsection 2.3.3 in [Kakaroumpas_Soler_2022]). Using Jensen’s inequality and the definition of the $A_{2}$ characteristic, it follows that
$\frac{w^{1/p^{\prime}}(R\setminus\Omega_{k+1})}{w^{1/p^{\prime}}(R)}\geq[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}^{-1}\frac{|R\setminus\Omega_{k+1}|}{|R|}\geq[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}^{-1}\cdot\frac{1}{2},$
so
$\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}v|\,\mathrm{dm}(x)\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-1/p}\int_{R}|V(x)^{-1/p}v|\,\mathrm{dm}(x).$
Thus
$\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-2/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{2/p^{\prime}}$
$\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\sim_{d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2},$
concluding the proof. ∎
Before we proceed to the second half of Theorem 2, recall that the strong dyadic Christ–Goldberg maximal function corresponding to a weight $W$ on $\mathbb{R}^{n+m}$ (and exponent $p$) is defined as
$M_{\bm{\mathcal{D}},W}f(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|W(x)^{1/p}f|\rangle_{R}\mathbf{1}_{R}(x),\qquad x\in\mathbb{R}^{n+m},\leavevmode\nobreak\ f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n+m};\mathbb{C}^{d}).$
Because of [Vuorinen_2023, Theorem 1.3] we have that the operator $M_{\bm{\mathcal{D}},W}$ is bounded when acting on $L^{p}(W)\rightarrow L^{p}(\mathbb{R}^{n+m})$ (note that the target space is unweighted) provided that $[W]_{A_{p},\bm{\mathcal{D}}}<\infty,$ specifically one has the bound
$\|M_{\bm{\mathcal{D}},W}\|_{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p}}^{2/(p-1)},$ (4.2)
for $1<p<\infty.$
Proposition 9.
Let $\ell$ be any bounded linear functional on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$ Then, there exists a unique $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with
$\ell(\Phi)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\forall\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$
Moreover, there holds
$\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}\|\ell\|.$
Proof.
We partially adapt the second half of the proof of [Isralowitz_2017, Theorem 1.3]. We denote by $\mathcal{H}$ the set of all $\mathrm{M}_{d}(\mathbb{C})$-valued $\mathrm{L}^{2}$ functions on $\mathbb{R}^{n+m}$ with finitely many nonzero biparameter cancellative Haar coefficients, and we identify such functions with finitely supported sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ in the obvious way. Then, for all $\Phi\in\mathcal{H},$ we have
$\Phi=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}$
and therefore
$\ell(\Phi)$ $=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)E_{ij})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)\ell(E_{ij})$
$=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})=(\Phi,B),$
where $E_{ij}$ is the $d\times d$-matrix with 1 at the $(i,j)$-entry and 0 at all other entries, $A(i,j)$ is the $(i,j)$-entry of a matrix $A,$ and
$B_{R}^{\varepsilon}(i,j)\coloneq\overline{\ell(h_{R}^{\varepsilon}E_{ji})},\qquad\forall i,j=1,\ldots,d,\leavevmode\nobreak\ \forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall\varepsilon\in\mathcal{E}.$
By Proposition 8 and since $\mathcal{H}$ is dense in $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),$ it suffices only to prove that the defined sequence $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ is in $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)$ with
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}\|\ell\|.$ (4.3)
Note that by the scalar, unweighted BMO equivalences we have
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}$
$\sim_{n,m,p,d}\sup_{\Omega}\frac{1}{|\Omega|^{1/p^{\prime}}}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m})}=:\mathcal{C},$
so it suffices only to prove that $\mathcal{C}\lesssim_{n,m,p,d}\|\ell\|.$ By the Monotone Convergence Theorem, we can without loss of generality assume that $B$ has only finitely many nonzero terms.
[…]
$\langle A,B\rangle\coloneq\mathrm{tr}(B^{\ast}A).$
[…]
$\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
$\sim_{n,m,p,d}\left\{\left|\int_{\mathbb{R}^{n+m}}\langle F(x),G(x)\rangle\,\mathrm{dm}(x)\right|:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\right\}$
=$=\bigg{\{}\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((G^{\varepsilon}_{R})^{\ast}F_{R}^{\varepsilon})\bigg{|}:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\bigg{\}},$ (4.4)
for all $F\in\mathcal{H}.$ We recall also the usual (unweighted) dyadic Littlewood–Paley estimates
$\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
[…]
Fix now any Lebesgue-measurable subset $\Omega$ of $\mathbb{R}^{n+m}$ with finite nonzero measure. Then, we have
$\bigg{(}\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p^{\prime}/2}\,\mathrm{dm}(x)\bigg{)}^{1/p^{\prime}}$
(4.1)${\sim}_{n,m,p,d}\bigg{\|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}h_{R}^{\varepsilon}\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
[…]
$=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}}\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}\left((B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\right)\bigg{|}$
$=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|(\hat{A},B)|=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|\ell(\hat{A})|$
$\leq\|\ell\|\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\hat{A}^{\varepsilon}_{R}\coloneq\begin{cases}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1},\text{ if }R\in\bm{\mathcal{D}}(\Omega)\\ \\ 0,\text{ otherwise}\end{cases},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))},$
for all $A\in\mathcal{H}.$ We set
$N_{\Omega}(x)\coloneq\sup_{R\in\bm{\mathcal{D}}(\Omega)}|V(x)^{-1/p}\mathcal{V}_{R}|\mathbf{1}_{R}(x),\quad x\in\mathbb{R}^{n+m}.$
By testing the strong dyadic Christ–Goldberg maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter matrix $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$), and because this operator is bounded (4.2), we obtain
$\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}.$
Thus, for all $A\in\mathcal{H}$ we have
$\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}$ $=\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$=\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\mathcal{V}_{R}A_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$\leq\int_{\Omega}N_{\Omega}(x)\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|A_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$\lesssim_{n,m,p,d}\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
$\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))},$
concluding the proof of (4.3).
Uniqueness of $B$ follows immediately by testing $\ell$ on sequences in $\mathcal{H}.$
Remark 10.
It is obvious that the proofs of Propositions 8, 9 work also in the one-parameter setting.
4.2 Two-matrix weighted bounds for paraproducts
Let $B:\mathbb{R}^{n+m}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|\bigg{)}^{1/2},$
where $\{B_{R}^{\varepsilon}\}$ is the sequence of biparameter Haar coefficients of $B$. Following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “pure” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
$\Pi_{\bm{\mathcal{D}},B}^{(11)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}h^{\varepsilon}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R},\quad\Pi_{\bm{\mathcal{D}},B}^{(00)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{\mathbf{1}_{R}}{|R|}B^{\varepsilon}_{R}f^{\varepsilon}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{1\oplus\varepsilon\oplus\delta}B_{R}^{\varepsilon}f_{R}^{\delta}.$
Here we denote
$1\oplus 1=0\oplus 0=0,\quad 1\oplus 0=0\oplus 1=1$
and extend these operations in the obvious entrywise fashion to $\mathcal{E}^{1},$ $\mathcal{E}^{2}$ and $\mathcal{E}.$ Moreover, abusing notation we denote $(1,\ldots,1)$ (where the number of entries is always clear from the context) by just $1$ Notice that $1\oplus\varepsilon_{i}\oplus\delta_{i}\neq 1,$ for all $\varepsilon,\delta\in\mathcal{E}$ with $\varepsilon_{i}\neq\delta_{i},$ $i=1,2.$
Clearly $(\Pi_{\bm{\mathcal{D}},B}^{(00)})^{\ast}=\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense. Observe also that a change of summation variables yields
$\Gamma_{\bm{\mathcal{D}},B}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{\varepsilon}B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\delta},$
therefore $(\Gamma_{\bm{\mathcal{D}},B})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense.
Proposition 11.
Let $d,p,U,V$ and $B$ be as above.
  1. (a)
    There holds
    $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\sim\|\Pi_{\bm{\mathcal{D}},B}^{(00)}f\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
  2. (b)
    There holds
    $\|\Gamma_{\bm{\mathcal{D}},B}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
Proof.
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,p,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt part of the proof of [Isralowitz_2017, Theorem 2.2]. Note that by the John–Nirenberg inequalities for (unweighted) scalar dyadic product BMO we have
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}$
$\sim_{p,d,n,m}\sup_{\Omega}\frac{1}{|\Omega|^{1/p}}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\coloneq\mathcal{C},$
where the supremum is taken over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{2}$ of nonzero finite measure. Therefore, it suffices to prove that
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\gtrsim\mathcal{C}$
and
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
Let us first see that $\mathcal{C}\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$ Let $\Omega$ be any Lebesgue-measurable subset of $\mathbb{R}^{2}$ of nonzero finite measure. Let also $e\in\mathbb{C}^{d}\setminus\{0\}$ be arbitrary. We test $\Pi_{\bm{\mathcal{D}},B}^{(11)}$ on the function $f\coloneq\mathbf{1}_{\Omega}U^{-1/p}e.$ Using Lemma 5.3 of [DKPS2024] we obtain
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\gtrsim\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\geq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}.$
Note also that $\|f\|_{\mathrm{L}^{p}(U)}=|\Omega|^{1/p}\cdot|e|.$ Therefore, we see that
$\frac{1}{|\Omega|^{1/p}}\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}|e|,$
for all $e\in\mathbb{C}^{d}.$ In view of (2.3) we deduce
$\frac{1}{|\Omega|^{1/p}}\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$
By part (2) of Lemma 3 we have
$|P\langle U^{-1/p}\rangle_{R}|=|\langle U^{-1/p}\rangle_{R}P^{\ast}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P|\sim_{p,d}|\mathcal{U}^{-1}_{R}P|=|P\mathcal{U}^{-1}_{R}|,\quad\forall P\in\mathrm{M}_{d}(\mathbb{C}),$
yielding the desired result.
We now prove that $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ Using Lemma 5.3 of [DKPS2024] we have
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\sim\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\leq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}|\mathcal{U}_{R}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\leq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{R\in\bm{\mathcal{D}}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}(\langle\widetilde{M}_{\bm{\mathcal{D}},U}f\rangle_{R})^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\sim_{p,n,m}\|\Pi_{\bm{\mathcal{D}},b}^{(11)}(\widetilde{M}_{\bm{\mathcal{D}},U}f)\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})},$
where $b=(b_{R}^{\varepsilon})_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}$ is the sequence given by
$b_{R}^{\varepsilon}\coloneq|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|,\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E},$
and
$\widetilde{M}_{\bm{\mathcal{D}},U}f\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|\mathcal{U}_{R}f|\rangle_{R}\mathbf{1}_{R}.$
By the well-known unweighted bounds for paraproducts in the scalar setting (see e. g. [Blasco_Pott_2005]) we have
$\|\Pi_{\bm{\mathcal{D}},b}^{(11)}\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{p,n,m}\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}.$
By definition, $\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}=\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ From [DKPS2024, Proposition 4.1] we also have
$\|\widetilde{M}_{\bm{\mathcal{D}},U}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim 1.$
It follows that
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\widetilde{M}_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|f\|_{\mathrm{L}^{p}(U)}.$
Finally, by duality we obtain
$\|\Pi_{\bm{\mathcal{D}},B}^{(00)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}$ $=\|\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}\|_{L^{p^{\prime}}(V^{\prime})\rightarrow L^{p^{\prime}}(U^{\prime})}\sim\|B^{\ast}\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(V^{\prime},U^{\prime},p^{\prime})}$
$\sim_{p,d}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We have
$(\Gamma_{\bm{\mathcal{D}},B}f,g)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\langle B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\varepsilon},g_{R}^{\delta}\rangle.$
Observe that if $A\in\mathrm{M}_{d}(\mathbb{C})$ and $x,y\in\mathbb{C}^{d},$ then
$\langle Ax,y\rangle$ $=x^{T}A^{T}\bar{y}=\mathrm{tr}(x^{T}A^{T}\bar{y})=\overline{\mathrm{tr}(A^{\ast}y\bar{x}^{T})}.$
It follows that
$|(\Gamma_{\bm{\mathcal{D}},B}f,g)|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}\text{tr}((B_{R}^{\varepsilon})^{\ast}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T})\bigg{|}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R|}}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\bigg{(}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\bigg{)}\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}|^{2}\cdot|\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$=2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
$\leq 2^{n+m}S_{\bm{\mathcal{D}},V^{\prime}}g(x)\widetilde{S}_{\bm{\mathcal{D}},U}f(x),$
for all $x\in\mathbb{R}^{n+m},$ where
$S_{\bm{\mathcal{D}},V^{\prime}}g\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2},\quad\widetilde{S}_{\bm{\mathcal{D}},U}f\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \delta\in\mathcal{E}\end{subarray}}|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}.$
Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq 2^{n+m}\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemma 5.2 and Corollary 5.4] that
$\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},\qquad\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})},$
yielding the desired bound. ∎
Next, following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “mixed” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
$\Pi_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(h_{R_{1}}^{\varepsilon_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}},\quad\Pi_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\varepsilon_{2}}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}h_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{1\oplus\varepsilon_{2}\oplus\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\\ \\ delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}h_{R}^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{1\oplus\varepsilon_{1}\oplus\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{\delta_{1},\varepsilon_{2}}_{R}.$
It is clear that $(\Pi_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(10)}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense. Observe also that a change of summation variables yields
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}_{R},$
therefore $(\Gamma_{\bm{\mathcal{D}},B}^{(10)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(10),\ast}$ as well as $(\Gamma_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(01),\ast}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense.
Two-matrix weighted bounds for the above mixed paraproducts can be easily deduced from two-matrix weighted bounds for the mixed type operators considered in [DKPS2024, Section 8]. Although at the time of [DKPS2024] only a rather incomplete treatment of the latter bounds was possible, they all follow now readily from Theorem 1 and Theorem 7. We give only one example below, the proof for the other operators being similar. Note that the parts of the proof that do not rely on the matrix weighted extension of the Fefferman–Stein vector valued inequalities were already carried out in [DKPS2024, Section 8]. Nevertheless, for the reader’s convenience we include full details.
Lemma 12.
Let $1<p<\infty$ and let $W$ be a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ For (suitable) functions $f:\mathbb{R}^{n+m}\to\mathbb{C},$ let
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
for all $x=(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Then, we have
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\beta}\|f\|_{\mathrm{L}^{p}(W)},$
where
$\beta=\begin{cases}1+\frac{2}{p}+\frac{1}{p-1},\text{ if }p\leq 2\\ \\ \frac{1}{2}+\frac{1}{p}+\frac{2}{p-1},\text{ if }p>2\end{cases}.$
To prove Lemma 12, we need a technical observation already present implicitly in the proof of [DKPS2024, Lemma 5.3].
Lemma 13.
Let $1<p<\infty$ and let $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ be a family of nonnegative measurable functions on $\mathbb{R}^{m}.$ Then, it holds
$\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\lesssim_{m,p}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}.$
Proof.
We use duality. Note that through an application of the monotone convergence theorem we may assume without loss of generality that only finitely many of the functions $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ are not identically equal to zero. Then, by the dyadic (unweighted) Littlewood–Payley estimates we have
$\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\sim_{m,p}\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{m})},$
where
$F\coloneq\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}h_{P}^{\varepsilon}.$
Let $g\in L^{p^{\prime}}(\mathbb{R}^{m})$ be arbitrary. Then, we have
$\int_{\mathbb{R}^{m}}|F(x)g(x)|\,\mathrm{dm}(x)\leq\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle\cdot|g_{P}^{\varepsilon}|$
$=\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\int_{\mathbb{R}^{m}}f_{P,\varepsilon}(x)h_{P}^{\varepsilon}(x)g_{P}^{\varepsilon}h_{P}^{\varepsilon}(x)\,\mathrm{dm}(x)$
$\leq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|g_{P}^{\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{m})}$
$\sim_{m,p}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\|g\|_{L^{p^{\prime}}(\mathbb{R}^{m})}.$
An appeal to the Riesz representation theorem concludes the proof. ∎
We now prove Lemma 12.
Proof of Lemma 12.
First of all, we have
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}^{p}=\int_{\mathbb{R}^{n}}A(x_{1})\,\mathrm{dm}(x_{1}),$
where
$A(x_{1})\coloneq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{2}}|\mathcal{W}_{R_{1}\times R_{2}}(f_{R_{2}}^{\varepsilon_{2},2})_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}),$
for all $x_{1}\in\mathbb{R}^{m}.$
Fix $x_{1}\in\mathbb{R}^{m}.$ For a.e. $x_{2}\in\mathbb{R}^{n},$ we denote by $\mathcal{W}_{x_{2},R_{1}}$ the reducing operator of the weight $W_{x_{2}}(y)\coloneq W(y,x_{2}),$ $y\in\mathbb{R}^{n}$ over any $R_{1}\in\mathcal{D}^{1}$ with respect to the exponent $p.$ For fixed $R_{1}\in\mathcal{D}^{1},$ we define $W_{R_{1}}(x_{2})\coloneq\mathcal{W}_{x_{2},R_{1}}^{p},$ for a.e. $x_{2}\in\mathbb{R},$ and denote by $\mathcal{W}_{R_{1},R_{2}}$ the reducing operator of $W_{R_{1}}$ over any $R_{2}\in\mathcal{D}^{2}$ with respect to the exponent $p.$ Applying now first (2.3), then part (1) of Lemma 3, and finally (2.4), we obtain
$\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R_{1}\times R_{2}}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\lesssim_{p,d}\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R_{1},R_{2}}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})$
$\lesssim_{p,d}\sup_{R_{1}\in\mathcal{D}^{1}}[W_{R_{1}}]_{A_{p},\mathcal{D}^{2}}^{\frac{1}{p}}\left\langle|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\right\rangle_{R_{2}}\mathbf{1}_{R_{1}}(x_{1})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\frac{1}{p}}\bigg{\langle}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{\rangle}_{R_{2}},$
for all $\varepsilon_{2}\in\mathcal{E}^{2}$ and $R_{2}\in\mathcal{D}^{2}.$ Observe that in the last $\lesssim_{p,d},$ we used (2.4), as well as the fact that the supremum of the integrals is dominated by the integral of the supremum. So
$A(x_{1})\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}B(x_{1}),$
where
$B(x_{1})\coloneq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{\langle}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{\rangle}_{R_{2}}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}).$
Since $x_{1}$ is fixed, by Lemma 13 we obtain
$B(x_{1})$ $\leq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})$
$=\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}).$
Thus, by Fubini–Tonelli we have
$\int_{\mathbb{R}^{n}}B(x_{1})\,\mathrm{dm}(x_{1})$
$\leq\int_{\mathbb{R}^{m}}\bigg{(}\int_{\mathbb{R}^{n}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})\bigg{)}\,\mathrm{dm}(x_{2}).$
For a.e. $x_{2}\in\mathbb{R}^{m},$ using Theorem 7 in the first step and (2.6.3) in the second step, we obtain
$\int_{\mathbb{R}^{n}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$\lesssim_{m,p,d}[W_{x_{2}}]_{A_{p},\mathcal{D}^{1}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}|W_{x_{2}}^{1/p}(x_{1})f_{R_{2}}^{\varepsilon_{2},2}(x_{1})|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}|W(x_{1},x_{2})^{1/p}f_{R_{2}}^{\varepsilon_{2},2}(x_{1})|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$=[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{1}),$
where we denote $W_{x_{1}}\coloneq W(x_{1},x_{2}),$ $x_{1}\in\mathbb{R}^{n}.$ Thus, applying Fubini–Tonelli again we get
$\int_{\mathbb{R}^{n}}B(x_{1})\,\mathrm{dm}(x_{1})$
$\lesssim_{m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{n}}\bigg{(}\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})\bigg{)}\,\mathrm{dm}(x_{1}).$
For a.e. $x_{1}\in\mathbb{R}^{n},$ applying first the matrix weighted bounds for the one-parameter dyadic square function from [Isralowitz_2020] and then (2.6.3), we get
$\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})$
$\lesssim_{n,p,d}[W_{x_{1}}]_{A_{p},\mathcal{D}_{2}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W_{x_{1}}(x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W(x_{1},x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2}).$
Putting the above estimates together, we finally deduce
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}^{p}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\alpha}\int_{\mathbb{R}^{n+m}}|W(x)^{1/p}f(x)|^{p}\,\mathrm{dm}(x),$
where
$\alpha=1+1+\max\{p,p^{\prime}\}+\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\},$
concluding the proof. ∎
We now prove two-matrix weighted bounds for the mixed paraproducts.
Proposition 14.
Let $d,p,U,V$ and $B$ be as above. If $\Pi$ is any of the above defined mixed biparameter paraproducts, then there holds
$\|\Pi\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
Proof.
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We treat as an example $\Gamma_{\bm{\mathcal{D}},B}^{(10)},$ the proof for the other mixed paraproducts being similar or following by duality. We have
$\left|(\Gamma_{\bm{\mathcal{D}},B}^{(10)}f,g)\right|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\langle B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},g_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}\rangle\bigg{|}$
$=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})\bigg{|}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$ (4.6)
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\left(\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\right)\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R_{1}|\cdot|R_{2}|^{2}}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}$
$\leq 2^{m}S_{\bm{\mathcal{D}},V^{\prime}}g(x)[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x),$
where
$S_{\bm{\mathcal{D}},V^{\prime}}g(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}$
and
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{U}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
for all $x=(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemmas 5.2] that
$\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
Moreover, Lemma 12 yields
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},$
proving (4.6). ∎
4.3 Two matrix weighted upper bounds for bicommutators
Two matrix weighted upper bounds for biparameter paraproducts lead naturally to two matrix weighed upper bounds for bicommutators. Here we overview very briefly the simplest case: that of bicommutators with Haar multipliers on $\mathbb{R}.$ The argument is a straightforward adaptation of the one in the (unweighted) $\mathrm{L}^{2}$ scalar valued case from [Blasco_Pott_2005].
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be a biparameter dyadic grid on $\mathbb{R}^{2},$ and let $\sigma_{1}=\{\sigma_{1}(I)\}_{I\in\mathcal{D}^{1}}$ and $\sigma_{2}=\{\sigma_{2}(J)\}_{J\in\mathcal{D}^{2}}$ be (finitely supported) sequences in $\{-1,0,1\}.$ For any function $f\in\mathrm{L}^{2}(\mathbb{R};\mathbb{C}^{d})$ we define
$T_{\sigma_{1}}f\coloneq\sum_{I\in\mathcal{D}^{1}}\sigma_{1}(I)h_{I}f_{I},\quad T_{\sigma_{2}}f\coloneq\sum_{J\in\mathcal{D}^{2}}\sigma_{2}(J)h_{J}f_{J}.$
We can then consider the operators $T^{1}_{\sigma_{1}}$ and $T^{2}_{\sigma_{2}}$ acting on functions $f\in\mathrm{L}^{2}(\mathbb{R}^{2};\mathbb{C}^{d})$ by
$T^{1}_{\sigma_{1}}f(x_{1},x_{2})\coloneq T_{\sigma_{1}}(f(\,\cdot\,,x_{2}))(x_{1}),\quad T^{2}_{\sigma_{2}}f(x_{1},x_{2})\coloneq T_{\sigma_{2}}(f(x_{1},\,\cdot\,))(x_{2}),$
for a.e. $(x_{1},x_{2})\in\mathbb{R}^{2}.$
Let $B:\mathbb{R}^{2}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. For scalar valued locally integrable functions $b$ on $\mathbb{R}^{2}$ it is shown in [Blasco_Pott_2005] that
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},b]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{b}]],$ (4.7)
where $\Lambda_{b}$ is the so-called symmetrized paraproduct given by
$\Lambda_{b}f\coloneq\Pi^{(11)}_{b}f+\Pi^{(10)}_{b}f+\Pi^{(01)}_{b}f+\Pi^{(00)}_{b}f.$
Applying (4.7) entrywise we deduce
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},B]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{B}]],$
where the symmetrized paraproduct $\Lambda_{B}$ is given by
$\Lambda_{B}f\coloneq\Pi^{(11)}_{B}f+\Pi^{(10)}_{B}f+\Pi^{(01)}_{B}f+\Pi^{(00)}_{B}f.$
Let now $1<p<\infty$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}\times\mathbb{R}.$ In the following estimates, all implied constants depend only on $d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Using (2.6.3) and the well-known two matrix weighted bounds in the one parameter setting we deduce
$\|T^{j}_{\sigma_{j}}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1,\quad j=1,2.$
Observe also that Proposition 11 and Proposition 14 immediately yield
$\|\Lambda_{B}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1.$
Thus, we obtain
$\|[T_{1},[T_{2},B]]\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}=\|[T_{1},[T_{2},\Lambda_{B}]]\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}$
$\leq\|T_{1}T_{2}\Lambda_{B}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}+\|T_{1}\Lambda_{B}T_{2}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}$
$+\|T_{2}\Lambda_{B}T_{1}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}+\|\Lambda_{B}T_{1}T_{2}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1.$
[…]
Spyridon Kakaroumpas
Würzburg Mathematics Center of Communication and Interaction,
Julius-Maximilians-Universität Würzburg,
Campus Hubland Nord,
Emil-Fischer-Straße 41,
97074 Würzburg, Germany
E-mail address:
spyridon.kakaroumpas@uni-wuerzburg.de
Odí Soler i Gibert
Universitat Politècnica de Catalunya - BarcelonaTech (UPC),
Pavelló I,
Diagonal 647,
08028 Barcelona, Catalunya
E-mail address:
odi.soler@upc.edu