### v2

# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

Abstract
We consider maximal operators acting on vector valued functions, that is functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts. For the reader’s convenience, we include an appendix explaining how to adapt the extrapolation for real convex body valued functions of Bownik and Cruz-Uribe to the setting of complex convex body valued functions that we treat.
1 Introduction
This paper deals with matrix weighted extensions of the Fefferman–Stein vector valued inequalities for the Hardy–Littlewood maximal function and the closely connected topic of matrix weighted extensions of the product BMO spaces. Before stating our main results, we briefly recall some historical background on each area.
The classical Fefferman–Stein vector valued inequalities, first proved by C. Fefferman and E. Stein [Fefferman_Stein_1971], state that for all $1<p,q<\infty$ one has
$\left\|\left(\sum_{k=1}^{\infty}|Mf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}\leq C(n,p,q)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})},$ (1.1)
for any sequence $\{f_{k}\}_{k=1}^{\infty}$ of (say) locally integrable complex valued functions on $\mathbb{R}^{n},$ where $M$ is the usual (uncentered) Hardy–Littlewood maximal function. That means
$Mf(x)\coloneq\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|f(x)|\,\mathrm{dm}(x),\quad x\in\mathbb{R}^{n},$
where the supremum ranges over all cubes $Q$ in $\mathbb{R}^{n}$ (with faces parallel to the coordinate hyperplanes), $|E|$ is the Lebesgue measure of a Lebesgue measurable set $E\subseteq\mathbb{R}^{n},$ and $\,\mathrm{dm}(x)$ denotes integration with respect to Lebesgue measure. The constant $C(n,p,q)$ in (1.1) depends only on $n,$ $p$ and $q.$ In [Fefferman_Stein_1971] also a weak type version of (1.1) is proved.
The estimate (1.1) is actually a special case of more general bounds for vector valued extensions of operators. That is, given some (not necessarily linear) operator $T$ acting boundedly on $\mathrm{L}^{p}(\mathbb{R}^{n})$ for some $1<p<\infty,$ we seek to find those $1<q<\infty$ satisfying an estimate of the form
$\left\|\left(\sum_{k=1}^{\infty}|Tf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}\leq C(T,n,p,q)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(\mathbb{R}^{n})}$ (1.2)
for any sequence $\{f_{k}\}^{\infty}_{k=1}.$ Such inequalities seem to have been studied for the first time systematically by J.-L. Rubio de Francia [Rubio_de_Francia_1985]. For a thorough modern exposition of the methods in [Rubio_de_Francia_1985] we refer to [Tao_notes_5_247A] as well as [Grafakos_Book_1].
The ideas in [Rubio_de_Francia_1985] already hinted at an intimate connection between extrapolation and bounds for vector valued extensions as in (1.2). By an extrapolation problem one understands the following. Given an operator $T$ acting on (suitable) functions on $\mathbb{R},$ we assume that for some $1<p<\infty$ it is already known or given that for all weights (that means, a.e. positive locally integrable functions) $w$ on $\mathbb{R}^{n}$ that belong to some class $C(p),$ one has the estimate $\|Tf\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,w)\|f\|_{\mathrm{L}^{p}(w)}.$ Given this information, find all $1<q<\infty$ and as well as associated classes $C(q)$ of weights on $\mathbb{R}^{n},$ such that for any $w\in C(q)$ one has an estimate of the form $\|Tf\|_{\mathrm{L}^{q}(w)}\leq C(T,n,q,w)\|f\|_{\mathrm{L}^{q}(w)}.$ Rubio de Francia [Rubio_de_Francia_1984] solved completely the extrapolation problem in the case that $C(p)$ coincides with the Muckenhoupt $A_{p}$ class, that is $w\in A_{p}$ if and only if
$[w]_{A_{p}}\coloneq\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}w(x)\,\mathrm{dm}(x)\right)\left(\frac{1}{|Q|}\int_{Q}w(x)^{-1/(p-1)}\,\mathrm{dm}(x)\right)^{p-1}<\infty,$
where the supremum ranges again over all cubes $Q\subseteq\mathbb{R}^{n}.$ In this case, beginning with any fixed $1<p<\infty$ and an extrapolation hypothesis holding for all weights $w\in A_{p},$ the extrapolation problem is solvable for any $1<q<\infty$ and all weights $w\in A_{q}.$ The extrapolation theorem of Rubio de Francia was subsequently further refined by various authors, until a sharp quantitative version of it was proved by J. Duoandikoetxea [Duoandikoetxea_2011]. A very thorough treatment of various forms of extrapolation with extensive historical background can be found in [Book_Extrapolation]. In fact, this method is so powerful, that it naturally yields weighted estimates for vector valued extensions, that is
$\left\|\left(\sum_{k=1}^{\infty}|Tf_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,q,w)\left\|\left(\sum_{k=1}^{\infty}|f_{k}|^{q}\right)^{1/q}\right\|_{\mathrm{L}^{p}(w)},$ (1.3)
as explained in detail in [Book_Extrapolation].
Inequalities of the form (1.3) are a major tool for estimating operators arising naturally when decomposing biparameter operators or bicommutators in simpler, localized pieces. Such a decomposition for the so called Journeé operators was established by H. Martikainen [Martikainen_2012] (generalizing an analogous decomposition proved earlier by T. Hytönen [Hytonen_2012] for Calderón–Zygmound operators in the context of the solution of the $A_{2}$ problem). I. Holmes, S. Petermichl and B. Wick [Holmes_Petermichl_Wick_2018] showed that these localized pieces can be estimated in terms of a weighted product BMO space. In the following, we recall the relevant definitions.
The classical space $\text{BMO}(\mathbb{R}^{n})$ consists of all locally integrable functions $b$ on $\mathbb{R}^{n}$ such that
$\|b\|_{\text{BMO}(\mathbb{R}^{n})}\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}|b(x)-\langle b\rangle_{Q}|\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all cubes $Q\subseteq\mathbb{R}^{n},$ and we denote $\langle b\rangle_{Q}\coloneq\frac{1}{|Q|}\int_{Q}b(x)\,\mathrm{dm}(x).$ The importance of this space is two-fold. First, it is the dual space to the (real variable) Hardy space $\mathrm{H}^{1}(\mathbb{R}^{n}).$ Second, the norm $\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}$ is the “correct” quantity controlling the boundedness of commutators $[T,b]=[T,M_{b}],$ where $M_{b}$ denotes (pointwise) multiplication by $b$ (called symbol of the commutator) and $T$ is a Calderón–Zygmound operator. This was proved for the Hilbert transform by Z. Nehari [Nehari_1957] and in full generality by R. R. Coifmann, Rochberg and G. Weiss [Coifman_Rochberg_Weiss_1976]. Moreover, the John–Nirenberg inequalities are an important intrinsic property of the space $\mathrm{BMO}(\mathbb{R}^{n}).$
B. Muckenhoupt and R. L. Wheeden [Muckenhoupt_Wheeden_BMO_1976] considered and studied the weighted BMO norm
$\|b\|_{\mathrm{BMO}(\nu)}\coloneq\sup_{Q}\frac{1}{\nu(Q)}\int_{Q}|b(x)-\langle b\rangle_{Q}|\,\mathrm{dm}(x),$ (1.4)
where the supremum ranges over all cubes $Q\subseteq\mathbb{R}^{n}$ and $\nu$ is a $A_{2}$ weight on $\mathbb{R}^{n}.$ A characterization of the two weighted boundedness of commutators in terms of a weighted BMO norm of the symbol was established by S. Bloom [Bloom_1985] for the Hilbert transform and later for arbitrary Calderón–Zygmund operators by I. Holmes, M. Lacey and B. Wick [Holmes_Lacey_Wick_2016]. In the latter work two weighted versions of (1.4) were introduced and associated John–Nirenberg inequalities were established. These played an important role in the commutator estimates in [Holmes_Lacey_Wick_2016].
The study of biparameter BMO spaces on product spaces $\mathbb{R}^{n}\times\mathbb{R}^{m}$ was initiated by S. Y. A. Chang [Chang_1979] and R. Fefferman [RFefferman_1979]. Here, “biparameter” refers to invariance of the considered function spaces under rescaling each coordinate variable of the domain of definition separately. Works [Chang_1979] and [RFefferman_1979] introduced and investigated the biparameter product BMO space $\text{BMO}(\mathbb{R}\times\mathbb{R})$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{2}$ (considered as the product space $\mathbb{R}\times\mathbb{R}$) such that
$\|b\|_{\text{BMO}(\mathbb{R}\times\mathbb{R})}\coloneq\sup_{\Omega}\bigg{(}\frac{1}{|\Omega|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ R\subseteq\Omega\end{subarray}}\left|(b,w_{R})\right|^{2}\bigg{)}^{1/2}<\infty,$
where the supremum reanges over all (say) bounded Borel subsets $\Omega$ of $\mathbb{R}^{2}$ with nonzero measure, $\bm{\mathcal{D}}$ stands for the family of all dyadic rectangles of $\mathbb{R}^{2},$ and $(w_{R})_{R\in\bm{\mathcal{D}}}$ is some (mildly regular) wavelet system adapted to $\bm{\mathcal{D}}.$ Here and below we denote
$(b,w_{R})\coloneq\int_{\mathbb{R}^{2}}b(x)w_{R}(x)\,\mathrm{dm}(x).$
The aforementioned works [Chang_1979] and [RFefferman_1979] established in particular that $\text{BMO}(\mathbb{R}\times\mathbb{R})$ is the dual to the biparameter Hardy space $\mathrm{H}^{1}(\mathbb{R}\times\mathbb{R}).$ Moreover, a dyadic version of this product BMO space is the correct space for characterizing the boundedness of bicommutators $[T_{1},[T_{2},b]],$ where $T_{1},T_{2}$ are Haar multipliers, as proved by [Blasco_Pott_2005].
A weighted version of the Chang–Fefferman product BMO space was introduced and studied by Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018] in the context of proving two weight upper bounds for biparameter praproducts. Namely, given a biparameter dyadic grid $\bm{\mathcal{D}}$ in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and a biparameter $\bm{\mathcal{D}}$-dyadic $A_{2}$ weight $w$ on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ [Holmes_Petermichl_Wick_2018] considers the dyadic Bloom type product space $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with the property
$\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)}\coloneq\sup_{\Omega}\bigg{(}\frac{1}{\nu(\Omega)}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|b_{R}^{\varepsilon}|^{2}\langle\nu^{-1}\rangle_{R}\bigg{)}^{1/2}<\infty,$ (1.5)
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ of nonzero finite measure. We refer to Section 2 for a detailed explanation of the notation and the terminology. In [Holmes_Petermichl_Wick_2018] an $\mathrm{H}^{1}$-BMO duality type result was established in this setting, which played a crucial role in the proofs of the upper bounds there. More recently, a two-weight version of (1.5) was defined in [Kakaroumpas_Soler_2022] and associated John–Nirenberg inequalities were established. These played an important role in [Kakaroumpas_Soler_2022] for characterizing the two weight boundedness of bicommutators with Haar multipliers, extending the aforementioned result of [Blasco_Pott_2005] to the two weight setting.
1.1 Main results
One of the main goals of this paper is to prove matrix weighted bounds for vector valued extensions of the Christ–Goldberg maximal function, which can be understood as a matrix weighted extension of the classical Fefferman–Stein vector valued inequalities for the Hardy–Littlewood maximal function.
Theorem 1.
Let $1<p<\infty.$ Consider a $(d\times d)$ matrix $A_{p}$ weight $W$ and a sequence of vector valued functions $\{f_{n}\}^{\infty}_{n=1}.$ Then, for each $1<q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}},$
and $M_{W}$ denotes the Christ–Goldberg maximal function corresponding to the weight $W$ and the exponent $p.$
We refer to Subsection 3.1 for a detailed explanation of the notation in Theorem 1. Theorem 1 yields readily a similar estimate for the so called modified Christ–Goldberg maximal function, as explained in Theorem 7 below.
We deduce Theorem 1 from a general principle for establishing matrix weighted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 5 below. Our method for deducing such bounds is inspired from [Book_Extrapolation]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [Cruz_Uribe_Bownik_Extrapolation], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the recent limited range extrapolation theorem for matrix weights proved in [Kakaroumpas_Nguyen_Vardakis_2024] yields similar bounds as in Theorem 5 that are valid only for a limited range of exponents. Note that the extrapolation for matrix weights and the other methods and techniques from [Cruz_Uribe_Bownik_Extrapolation] (which also belong to the foundations of the work in [Kakaroumpas_Nguyen_Vardakis_2024]) concern the setting of real convex body valued functions. An extension to complex convex body valued functions presents nevertheless no difficulties thanks to previous work in [DKPS2024]. For the reader’s convenience we supply details in the appendix.
Theorem 1 and its counterpart for the modified Christ–Goldberg maximal function, Theorem 7, allow one to complete the proof of matrix weighted bounds for general (not necessarily paraproduct free) Journé operators given in [DKPS2024] for any $1<p<\infty$ (not necessarily $p=2$), as explained in [DKPS2024, Section 8] (thus avoiding the use of extrapolation for biparameter matrix weights as in [Vuorinen_2023]). In this paper we focus on the application of Theorem 1 for setting up the foundations of a theory of two matrix weighted product BMO. Namely, let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2},$
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure. This definition is an extension of one of the equivalent definitions for the space of two matrix weighted one-parameter BMO, whose study was initiated in [Isralowitz_Kwon_Pott2017], [Isralowitz_2017] and culminated in [Isralowitz_Pott_Treil_2022]. Moreover, for every sequence $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C}),$ we define
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\coloneq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{1}(\mathbb{R}^{n+m})},$
which is the direct biparameter analog of the one parameter two matrix weighted $\mathrm{H}^{1}$ norm from [Isralowitz_2017]. In this context, we prove the following result.
Theorem 2.
Let $U,V,p,\bm{\mathcal{D}}$ be as above. For any $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p),$ the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C}$ given by
$\ell_{B}(\Phi)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$
is well-defined and bounded with $\|\ell\|\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$ where the implied constants depend only on $n,m,p,d$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Conversely, for every bounded linear functional $\ell$ on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ there is $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with $\ell=\ell_{B}.$
We note that our proof of Theorem 2 trivially works also in the one-parameter setting, thus answering to the positive the question posed in [Isralowitz_2017] about the extension of the one parameter two matrix weighted $\mathrm{H}^{1}$-BMO duality proved there from $p=2$ to arbitrary exponents $1<p<\infty.$
The duality result of Theorem 2 coupled with Theorem 1 yields readily two-matrix weighted bounds for biparameter paraproducts, see Proposition 11 and Proposition 14 below. These yield in turn already some two matrix weighted upper bounds for bicommutators, see Subsection 4.3 below.
Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper. In addition, the authors are indebted to both Marcin Bownik and David Cruz-Uribe for carefully reading this article and for providing valuable feedback that improved the clarity and readability of this work.
2 Background
In this section we review the central definitions for our work. We also recall some basic facts about the various objects we deal with, in most cases without proof. References with further details are provided in this case.
2.1 Dyadic grids
For definiteness, in what follows intervals in $\mathbb{R}$ will always be assumed to be left-closed, right-open and bounded. A cube in $\mathbb{R}^{n}$ will be a set of the form $Q=I_{1}\times\ldots\times I_{n}$, where $I_{k},\leavevmode\nobreak\ k=1,\ldots,n$ are intervals in $\mathbb{R}$ of the same length, which we denote by $\ell(Q)\coloneq|I_{1}|.$ A rectangle in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ (with sides parallel to the coordinate axes) will be a set of the form $R=R_{1}\times R_{2}$, where $R_{1}$ is a cube in $\mathbb{R}^{n}$ and $R_{2}$ is a cube in $\mathbb{R}^{m}$.
A collection $\mathcal{D}$ of intervals in $\mathbb{R}$ will be said to be a dyadic grid in $\mathbb{R}$ if one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, such that the following hold:
  • for all $k\in\mathbb{Z}$, $\mathcal{D}_{k}$ forms a partition of $\mathbb{R}$, and all intervals in $\mathcal{D}_{k}$ have the same length $2^{-k}.$
  • for all $k\in\mathbb{Z}$, every $J\in\mathcal{D}_{k}$ can be written as a union of exactly $2$ intervals in $\mathcal{D}_{k+1}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}$ if
$\mathcal{D}\coloneq\left\{[m2^{k},(m+1)2^{k}):\leavevmode\nobreak\ k,m\in\mathbb{Z}\right\}.$
A collection $\mathcal{D}$ of cubes in $\mathbb{R}^{n}$ will be said to be a dyadic grid in $\mathbb{R}^{n}$ if for some dyadic grids $\mathcal{D}^{1},\ldots,\mathcal{D}^{n}$ in $\mathbb{R}$ one can write $\mathcal{D}=\bigcup_{k\in\mathbb{Z}}\mathcal{D}_{k}$, where
$\mathcal{D}_{k}=\{I_{1}\times\ldots\times I_{n}:\leavevmode\nobreak\ I_{i}\in\mathcal{D}^{i},\leavevmode\nobreak\ |I_{i}|=2^{-k},\leavevmode\nobreak\ i=1,\ldots,n\}.$
We say that $\mathcal{D}$ is the standard dyadic grid in $\mathbb{R}^{n}$ if
$\mathcal{D}\coloneq\{[m_{1}2^{k},(m_{1}+1)2^{k})\times\dots\times[m_{n}2^{k},(m_{n}+1)2^{k}):\leavevmode\nobreak\ k,m_{1},\ldots,m_{n}\in\mathbb{Z}\}.$
If $\mathcal{D}$ is a dyadic grid in $\mathbb{R}^{n}$, then we denote
$\mathrm{ch}_{i}(Q)\coloneq\{K\in\mathcal{D}:\leavevmode\nobreak\ K\subseteq Q,\leavevmode\nobreak\ |K|=2^{-in}|Q|\},\qquad Q\in\mathcal{D},\leavevmode\nobreak\ i=0,1,2,\ldots.$
A collection $\bm{\mathcal{D}}$ is said to be a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ if for a dyadic grid $\mathcal{D}^{1}$ in $\mathbb{R}^{n}$ and a dyadic grid $\mathcal{D}^{2}$ in $\mathbb{R}^{m}$ we have
$\bm{\mathcal{D}}\coloneq\{R_{1}\times R_{2}:\leavevmode\nobreak\ R_{i}\in\mathcal{D}_{i},\leavevmode\nobreak\ i=1,2\},$
and in this case we write (slightly abusing the notation) $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$.
If $\bm{\mathcal{D}}$ is a product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then we denote
$\mathrm{ch}_{i}(R)\coloneq\{Q_{1}\times Q_{2}:\leavevmode\nobreak\ Q_{i}\in\mathrm{ch}_{i_{j}}(R_{j}),\leavevmode\nobreak\ j=1,2\},$
where
$R\in\bm{\mathcal{D}},\leavevmode\nobreak\ i=(i_{1},i_{2}),\leavevmode\nobreak\ i_{1},i_{2}=0,1,2,\ldots.$
We say that $R$ is the $i$-th ancestor of $P$ in $\bm{\mathcal{D}}$ if $P\in\mathrm{ch}_{i}(R)$.
2.2 Haar systems
2.2.1 Haar system on $\mathbb{R}$
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}$. For any interval $I\in\mathcal{D}$ we denote by $h^{0}_{I},h^{1}_{I}$ the $\mathrm{L}^{2}$-normalized cancellative and noncancellative respectively Haar functions over the interval $I\in\mathcal{D}$, that is
$h^{0}_{I}\coloneq\frac{\mathbf{1}_{I_{+}}-\mathbf{1}_{I_{-}}}{\sqrt{|I|}},\qquad h^{1}_{I}\coloneq\frac{\mathbf{1}_{I}}{\sqrt{|I|}}$
(so $h^{0}_{I}$ has mean 0). For simplicity we denote $h_{I}\coloneq h_{I}^{0}$. For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R})$, we denote $f_{I}\coloneq(f,h_{I})$, $I\in\mathcal{D}$. It is well-known that the system $\{h_{I}\}_{I\in\mathcal{D}}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R})$. Of course, all these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.2 Haar system on $\mathbb{R}^{n}$
Let $\mathcal{D}$ be a dyadic grid in $\mathbb{R}^{n}$. We denote
$\mathcal{E}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of one-parameter signatures. For a cube $Q=I_{1}\times\dots\times I_{n}\in\mathcal{D}$ and $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{n})\in\mathcal{E},$ we denote by $h^{\varepsilon}_{Q}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the cube $Q$ corresponding to the signature $\varepsilon$ defined by
$h^{\varepsilon}_{Q}(x)\coloneq h^{\varepsilon_{1}}_{I_{1}}(x_{1})\dots h^{\varepsilon_{n}}_{I_{n}}(x_{n}),\qquad x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}.$
For any function $f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n})$ we denote $f^{\varepsilon}_{Q}\coloneq(f,h^{\varepsilon}_{Q})$. It is well-known that the system $\{h^{\varepsilon}_{Q}\colon\leavevmode\nobreak\ Q\in\mathcal{D},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n})$. All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions in the obvious entrywise way.
2.2.3 Haar system on the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ We denote $\mathcal{E}\coloneq\mathcal{E}^{1}\times\mathcal{E}^{2}$, where
$\mathcal{E}^{1}\coloneq\{0,1\}^{n}\setminus\{(1,\ldots,1)\},\leavevmode\nobreak\ \mathcal{E}^{2}\coloneq\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$
We call $\mathcal{E}$ the set of biparameter signatures. For $R=R_{1}\times R_{2}\in\bm{\mathcal{D}}$ and $\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\mathcal{E},$ we denote by $h_{R}^{\varepsilon}$ the $\mathrm{L}^{2}$-normalized cancellative Haar function over the rectangle $R$ corresponding to the signature $\varepsilon$ defined by
$h_{R}^{\varepsilon}=h_{R_{1}}^{\varepsilon_{1}}\otimes h_{R_{2}}^{\varepsilon_{2}},$
that is
$h_{R}^{\varepsilon}(x_{1},x_{2})=h_{I}^{\varepsilon_{1}}(x_{1})h_{J}^{\varepsilon_{2}}(x_{2}),\qquad(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$
For any function $f\in\mathrm{L}_{\mathrm{loc}}^{1}(\mathbb{R}^{n+m})$ we denote $f_{R}^{\varepsilon}\coloneq(f,h_{R}^{\varepsilon})$. From the corresponding one-parameter facts, we immediately deduce that the system $\{h_{R}^{\varepsilon}\colon R\in\bm{\mathcal{D}},\varepsilon\in\mathcal{E}\}$ forms an orthonormal basis for $\mathrm{L}^{2}(\mathbb{R}^{n+m}).$ For $P\in\mathcal{D}^{1},$ $Q\in\mathcal{D}^{2}$, $\varepsilon_{1}\in\mathcal{E}^{1}$ and $\varepsilon_{2}\in\mathcal{E}^{2}$ we denote
$f^{\varepsilon_{1},1}_{P}(x_{2})\coloneq(f(\,\cdot\,,x_{2}),h_{P}^{\varepsilon_{1}}),\quad x_{2}\in\mathbb{R}^{m},$
$f^{\varepsilon_{2},2}_{Q}(x_{1})\coloneq(f(x_{1},\,\cdot\,),h_{Q}^{\varepsilon_{2}}),\quad x_{1}\in\mathbb{R}^{n}.$
All these notations and facts extend to $\mathbb{C}^{d}$-valued and $\mathrm{M}_{d}(\mathbb{C})$-valued functions entrywise.
Finally, we remark that the dimensions $n,m$ will always be clear from the context, as well as whether $\mathcal{E}$ refers to the set of one-parameter or biparameter signatures.
2.3 Matrix norms in terms of column norms
In the sequel we denote by $\{e_{1},\ldots,e_{d}\}$ the standard basis of $\mathbb{C}^{d}.$ We will be often using the fact that
$|A|\sim_{d}\sum_{k=1}^{d}|Ae_{k}|,\quad\forall A\in\mathrm{M}_{d}(\mathbb{C}),$ (2.1)
without explicitly mentioning it. Note also that for all $0<p<\infty$ and for nonnegative numbers $x_{1},\ldots,x_{d}$ we have the estimate
$\min(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}\leq\left(\sum_{i=1}^{d}x_{i}\right)^{p}\leq\max(1,d^{p-1})\sum_{i=1}^{d}x_{i}^{p}.$
In particular, if $F_{k}:\mathbb{R}^{n}\rightarrow\mathrm{M}_{d}(\mathbb{C})$ is a sequence of Lebesgue-measurable functions and $0<p,q,r<\infty,$ then we have
$\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\left(\sum_{i=1}^{d}|F_{k}(x)e_{i}|\right)^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q,r}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}\sum_{i=1}^{d}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,p,q}\left(\int_{\mathbb{R}^{n}}\sum_{i=1}^{d}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}$
$\sim_{d,q}\sum_{i=1}^{d}\left(\int_{\mathbb{R}^{n}}\left(\sum_{k=1}^{\infty}|F_{k}(x)e_{i}|^{r}\right)^{p}\,\mathrm{dm}(x)\right)^{q}.$ (2.2)
2.4 Matrix weighted matrix-valued Lebesgue spaces
A function $W$ on $\mathbb{R}^{n}$ is said to be a $d\times d$-matrix valued weight, or just matrix weight, if it is a locally integrable $\mathrm{M}_{d}(\mathbb{C})$-valued function such that $W(x)$ is a positive-definite matrix for a.e. $x\in\mathbb{R}^{n}.$ Here, by locally integrable matrix valued function $W(x)$ we mean that the scalar valued function $|W(x)|,$ that is the matrix norm, is locally integrable.
Given a $d\times d$-matrix weight $W$ on $\mathbb{R}^{n}$ and $1<p<\infty,$ we define the norm
$\|F\|_{\mathrm{L}^{p}(W)}\coloneq\left(\int_{\mathbb{R}^{n}}|W(x)^{1/p}F(x)|^{p}\,\mathrm{dm}(x)\right)^{1/p},$
for all $\mathrm{M}_{d}(\mathbb{C})$-valued measurable functions $F$ on $\mathbb{R}^{n}.$ This norm defines the matrix weighted matrix-valued Lebesgue space $\mathrm{L}^{p}(W).$
2.5 Reducing operators
Let $1<p<\infty$. Let $E$ be a bounded measurable subset of $\mathbb{R}^{n}$ of nonzero measure. Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ that is integrable over $E$ (meaning that its matrix norm is integrable, so $\int_{E}|W(x)|\,\mathrm{dm}(x)<\infty$) and that takes a.e. values in the set of positive-definite $(d\times d)$ matrices. It is proved in [Goldberg_2003, Proposition 1.2] that there exists a (not necessarily unique) positive-definite matrix $\mathcal{W}_{E}\in\mathrm{M}_{d}(\mathbb{C})$, called reducing operator of $W$ over $E$ with respect to the exponent $p$, such that
$\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}\leq|\mathcal{W}_{E}\,e|\leq\sqrt{d}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}e|^{p}\,\mathrm{dm}(x)\right)^{1/p}$
for all $e\in\mathbb{C}^{d}.$ If $d=1$, i.e. $W$ is scalar-valued, then one can clearly take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/p},$ where we denote the average of $W$ over $E$ by
$\langle W\rangle_{E}=\frac{1}{|E|}\int_{E}W(x)\,\mathrm{dm}(x).$
Moreover, if $p=2$, then
$\frac{1}{|E|}\int_{E}|W(x)^{1/2}e|^{2}\,\mathrm{dm}(x)=\frac{1}{|E|}\int_{E}\langle W(x)e,e\rangle\,\mathrm{dm}(x),$
which by linearity is equal to
$\langle\langle W\rangle_{E}\,e,e\rangle=|\langle W\rangle_{E}^{1/2}e|^{2}$
for all $e\in\mathbb{C}^{d}.$ Thus in this special case one can take $\mathcal{W}_{E}\coloneq\langle W\rangle_{E}^{1/2}.$
Assume in addition now that the function $W^{\prime}\coloneq W^{-1/(p-1)}$ is also integrable over $E$. Then we let $\mathcal{W}^{\prime}_{E}$ be the reducing matrix of $W^{\prime}$ over $E$ corresponding to the exponent $p^{\prime}\coloneq p/(p-1)$, so that
$|\mathcal{W}^{\prime}_{E}\,e|\sim_{d}\left(\frac{1}{|E|}\int_{E}|W^{\prime}(x)^{1/p^{\prime}}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}=\left(\frac{1}{|E|}\int_{E}|W(x)^{-1/p}e|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p^{\prime}}$
for all $e\in\mathbb{C}^{d}.$ Note that one can take $\mathcal{W}^{\prime\prime}_{E}=\mathcal{W}_{E}$. Observe that
$|\mathcal{W}_{E}\mathcal{W}^{\prime}_{E}|\sim_{p,d}\left(\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)\right)^{1/p}.$
For a detailed exposition of reducing operators we refer for example to [DKPS2024]. Here we just state the following estimates for later convenience.
Lemma 3.
Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued function on $E$ taking a.e. positive-definite values, and such that $W$ and $W^{\prime}=W^{-1/(p-1)}$ are integrable over $E$ for some $1<p<\infty.$ Set
$C_{E}\coloneq\frac{1}{|E|}\int_{E}\left(\frac{1}{|E|}\int_{E}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x).$
We consider reducing operators of $W$ with respect to exponent $p$ and of $W^{\prime}$ with respect to exponent $p^{\prime}.$
  1. (1)
    There holds
    $|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{1/p}{\langle|W^{1/p}e|\rangle}_{E}$
    and
    $|\mathcal{W}_{E}^{-1}e|\leq|\mathcal{W}_{E}^{\prime}e|\lesssim_{p,d}C_{E}^{1/p}|\mathcal{W}_{E}^{-1}e|,$
    for all $e\in\mathbb{C}^{d}.$
  2. (2)
    There holds
    $|{\langle W^{1/p}\rangle}_{E}e|\leq|\mathcal{W}_{E}e|\lesssim_{p,d}C_{E}^{d/p}|{\langle W^{1/p}\rangle}_{E}e|,$
    for all $e\in\mathbb{C}^{d}.$
A proof of part (1) can be found, for example, in [DKPS2024]. A proof of part (2) can be found in [Isralowitz_Kwon_Pott2017].
Let $E,F$ be measurable subsets of $\mathbb{R}^{n},\mathbb{R}^{m}$ respectively with $0<|E|,|F|<\infty.$ Let $1<p<\infty.$ Let $W$ be a $\mathrm{M}_{d}(\mathbb{C})$-valued integrable function on $E\times F$ taking a.e. values in the set of positive-definite $d\times d$-matrices. For all $x_{1}\in E,$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in F.$ For a.e. $x_{1}\in E,$ denote by $\mathcal{W}_{x_{1},F}$ the reducing operator of $W_{x_{1}}$ over $F$ with respect to the exponent $p.$ It is proved in [DKPS2024] (see also [Cruz_Uribe_Bownik_Extrapolation]) that one can choose the reducing operator $\mathcal{W}_{x_{1},F}$ in a way that is measurable in $x_{1}.$
Set $W_{F}(x_{1})\coloneq\mathcal{W}_{x_{1},F}^{p},$ for a.e. $x_{1}\in E.$ Then $W_{F}\in\mathrm{L}^{1}(E;\mathrm{M}_{d}(\mathbb{C})).$ It is proved in [DKPS2024] that
$|\mathcal{W}_{F,E}e|\sim_{p,d}|\mathcal{W}_{E\times F}e|,\qquad\forall e\in\mathbb{C}^{d},$ (2.3)
where $\mathcal{W}_{F,E}$ is the reducing operator of $W_{F}$ over $E$ with respect to the exponent $p,$ and $\mathcal{W}_{E\times F}$ is the reducing operator of $W$ over $E\times F$ with respect to the exponent $p.$
2.6 Matrix $A_{p}$ weights
2.6.1 One-parameter matrix $A_{p}$ weights
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}$. We say that $W$ is a one-parameter $d\times d$-matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n})}\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all cubes $Q$ in $\mathbb{R}^{n}$. Note that if $W$ is a $(d\times d)$ matrix valued $A_{p}$ weight on $\mathbb{R}^{n}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n})}^{1/p^{\prime}}\sim_{p,d}[W]_{A_{p}(\mathbb{R}^{n})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n})}\sim_{p,d}\sup_{Q}|\mathcal{W}^{\prime}_{Q}\mathcal{W}_{Q}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\mathcal{D}$ is any dyadic grid in $\mathbb{R}^{n}$, we define
$[W]_{A_{p},\mathcal{D}}\coloneq\sup_{Q\in\mathcal{D}}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x),$
and we say that $W$ is a one-parameter $(d\times d)$ matrix valued $\mathcal{D}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\mathcal{D}}<\infty$.
2.6.2 Biparameter matrix $A_{p}$ weights
Let $W$ be a $(d\times d)$ matrix valued weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$. We say that $W$ is a biparameter $(d\times d)$ matrix valued $A_{p}$ weight if
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\coloneq\sup_{R}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x)<\infty,$
where the supremum is taken over all rectangles $R$ in $\mathbb{R}^{n}$ (with sides parallel to the coordinate axes). Note that if $W$ is a $(d\times d)$ matrix valued biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$, then $W^{\prime}\coloneq W^{-1/(p-1)}$ is a $(d\times d)$ matrix valued biparameter $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with $[W^{\prime}]_{A_{p^{\prime}}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p^{\prime}}\sim_{d,p}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}^{1/p}$, and
$[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}\sim_{p,d}\sup_{R}|\mathcal{W}^{\prime}_{R}\mathcal{W}_{R}|^{p},$
where the reducing matrices for $W$ correspond to exponent $p$, and those for $W^{\prime}$ correspond to exponent $p^{\prime}$.
If $\bm{\mathcal{D}}$ is any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, we define
$[W]_{A_{p},\bm{\mathcal{D}}}\coloneq\sup_{R\in\bm{\mathcal{D}}}\frac{1}{|R|}\int_{R}\left(\frac{1}{|R|}\int_{R}|W(x)^{1/p}W(y)^{-1/p}|^{p^{\prime}}\,\mathrm{dm}(y)\right)^{p/p^{\prime}}\,\mathrm{dm}(x);$
we say that $W$ is a biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p}$ weight if $[W]_{A_{p},\bm{\mathcal{D}}}<\infty$.
2.6.3 One-parameter restriction of biparameter matrix $A_{p}$ weights
Let $1<p<\infty.$ Let $W$ be a $(d\times d)$ matrix biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ set $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}.$ It is proved in [DKPS2024] that
$[W_{x_{1}}]_{A_{p}(\mathbb{R}^{m})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
Fix any cube $Q$ in $\mathbb{R}^{m}.$ For a.e. $x_{1}\in\mathbb{R}^{n},$ let $\mathcal{W}_{x_{1},Q}$ be the reducing operator of $W_{x_{1}}(x_{2})\coloneq W(x_{1},x_{2}),$ $x_{2}\in\mathbb{R}^{m}$ over $Q$ with respect to the exponent $p.$ Set $W_{Q}(x_{1})\coloneq\mathcal{W}_{x_{1},Q}^{p},$ for a.e. $x_{1}\in\mathbb{R}^{n}.$ It is shown in [DKPS2024] that
$[W_{Q}]_{A_{p}(\mathbb{R}^{n})}\lesssim_{p,d}[W]_{A_{p}(\mathbb{R}^{n}\times\mathbb{R}^{m})}.$ (2.4)
Of course, the dyadic version of this is also true. Moreover, both versions remain true if one “slices” with respect to the first coordinate instead of the second one.
2.7 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ In other words, the “symmetric convex sets” as we have defined them are precisely the balanced convex sets. We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite, we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [Cruz_Uribe_Bownik_Extrapolation, Cruz_Uribe_Extrapolation] and for a detailed exposition see [Aubin_2009].
Since through the article we will consider both functions taking values in $\mathbb{C}^{d}$ and functions taking values in $\mathcal{K}_{\mathrm{bcs}},$ we will use a typographic convention to avoid confusion. We will denote functions taking values in $\mathbb{C}^{d}$ with lowercase letters $f,g,h,\ldots$ On the other hand, we will use uppercase letters $F,G,H,\ldots$ to denote functions taking values in $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ In any case, we will also explicitly state the target space of the functions we use.
Given a set $K\subseteq\mathbb{C}^{d},$ let us define its norm by
$|K|\coloneq\sup\{|v|\colon v\in K\}.$
For a matrix $A$ we will denote its usual norm (given by its largest singular value) by $|A|.$ The action of matrix weights on convex body valued functions is given by the next definition. Given a convex body $K$ and a positive definite matrix $A,$ define the product
$AK\coloneq\{Au\colon u\in K\}$
and observe that $AK$ will also be a convex body.
We will say that a function $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ is measurable if for every open set $U\subseteq\mathbb{C}^{d}$ it holds that the set
$F^{-1}(U)\coloneq\{x\in\mathbb{R}^{n}\colon F(x)\cap U\neq\emptyset\}$
is measurable (in the sense of Lebesgue). A convex body valued function $F$ is measurable if and only if there exists a sequence $\{f_{k}\}_{k\geq 1}$ of measurable functions $f_{k}\colon\mathbb{R}^{n}\rightarrow\mathbb{C}^{d}$ such that
$F(x)=\overline{\{f_{k}(x)\colon k\geq 1\}}$ (2.5)
for every $x\in\mathbb{R}^{n}$ (see [Aubin_2009, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ Here we restrict to functions taking values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ due to the applications that we consider later. Nonetheless, the previous definitions and concepts apply verbatim to functions taking values on $\mathcal{K}(\mathbb{C}^{d}).$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9]). We include its proof for completeness, although the arguments of Bownik and Cruz-Uribe are equally valid for complex convex bodies in this case.
Lemma 4.
Consider a measurable function $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}).$ Then there exists $f\in S^{0}(F)$ such that
$|f(x)|=|F(x)|$
for all $x\in\mathbb{R}^{n}.$
Proof.
Through this proof, we identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$ in the usual way. Take the sequence $\{f_{k}\}$ of selection functions satisfying (2.5). Then, the function given by
$g_{0}(x)=\sup\{|v|\colon v\in F(x)\}=\sup\{|f_{k}(x)|\colon k\geq 1\}$
is also measurable. This allows us to define the function
$F_{0}(x)=\{v\in F(x)\colon|v|=g_{0}(x)\}=F(x)\cap\mathbf{{S}}_{\mathbb{C}^{{d}}}$
taking values on $\mathcal{K}(\mathbb{C}^{d}),$ where $\mathbf{{S}}_{\mathbb{C}^{{d}}}=\{v\in\mathbb{C}^{d}\colon|v|=1\}$ denotes the complex $(d-1)$-dimensional sphere. Since both $F$ and $g_{0}\mathbf{{S}}_{\mathbb{C}^{{d}}}$ are measurable, so is $F_{0}$ (see [Aubin_2009, Theorems 8.2.2 and 8.2.4]).
It is only left to choose $v(x)\in F_{0}(x)$ in a measurable way. This is done by choosing maximal vectors $v$ in every other real coordinate iteratively (in the $\mathbb{R}^{2d}$ sense). Let $P_{1}$ be the (continuous) projection onto the first complex coordinate and define
$g_{1}(x)=\sup_{k}|P_{1}(f_{k}(x))|.$
The function $g_{1}\colon\Omega\rightarrow[0,\infty)$ is then measurable. Next, use this function and the symmetry of $F(x)$ to define $F_{1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ as
$F_{1}(x)=\{v\in F_{0}(x)\colon P_{1}(v)=g_{1}(x)\}=F_{0}(x)\cap(\{g_{1}(x)\}\times\mathbb{C}^{d-1}).$
As before, $F_{1}$ is also a measurable function. Assume that we have defined $F_{j}$ measurable and taking values on $\mathcal{K}(\mathbb{C}^{d}).$ Define $F_{j+1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ by taking the set of points of $F_{j}(x)$ with maximal modulus of the $j+1$ complex coordinate, positive real part and zero imaginary part for the same coordinate. Also $F_{j+1}$ will be measurable because of the same reasons as before. Eventually, we get the measurable function $F_{d}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ and $F_{d}(x)$ must be a singleton for every $x\in\Omega,$ this is $F_{d}(x)=\{v(x)\}$ for some $v(x)\in\mathbb{C}^{d},$ because of the maximality of the modulus of each complex coordinate and the restriction of having each complex coordinate lying on the positive real axis. ∎
In order to define integrals of convex body valued functions, consider first the set
$S^{1}(F)\coloneq\{f\in S^{0}(F)\colon f\in\mathrm{L}^{1}(\mathbb{R}^{n})\}$
of integrable selection functions for $F.$ The Aumann integral of $F$ is then defined as
$\int_{\mathbb{R}^{n}}F(x)\,\,\mathrm{dm}(x)=\left\{\int_{\mathbb{R}^{n}}f(x)\,\,\mathrm{dm}(x)\colon f\in S^{1}(F)\right\}.$
In this work, we will restrict to integrably bounded convex body valued functions, that is to functions $F$ such that $|F(x)|$ is integrable. In particular, in this case $S^{0}(F)=S^{1}(F).$ If $|F(x)|$ is only locally integrable, we will say that $F$ is locally integrably bounded. For further convenience, given a cube $Q\subseteq\mathbb{R}^{n},$ we define the averaging operator $A_{Q}$ by
$A_{Q}F(x)\coloneq\frac{\mathbf{1}_{Q}(x)}{|Q|}\int_{Q}F(y)\,\,\mathrm{dm}(y),$
where $F$ is a locally integrably bounded convex body valued function.
Given $1\leq p<\infty,$ we define the Lebesgue space of convex body valued functions $\mathrm{L}^{p}(\mathbb{R}^{n},\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})),$ or just $\mathrm{L}^{p}$ when there is no ambiguity, as the set of functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that
$\left\|{F}\right\|_{{\mathrm{L}^{p}}}\coloneq\left(\int_{\mathbb{R}^{n}}|F(x)|^{p}\,\,\mathrm{dm}(x)\right)^{1/p}<\infty.$
The space of convex body valued functions $\mathrm{L}^{\infty}(\mathbb{R}^{n},\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})),$ or just $\mathrm{L}^{\infty},$ is defined as the set of convex body valued functions $F$ for which
$\left\|{F}\right\|_{{\mathrm{L}^{\infty}}}\coloneq\operatorname*{ess\,sup}\{|F(x)|\colon x\in\mathbb{R}^{n}\}<\infty.$
Given a $(d\times d)$-matrix weight $W,$ we define the weighted Lebesgue space $\mathrm{L}^{p}(W)$ of convex body valued functions as the set of functions $F$ for which $W(x)^{1/p}F(x)\in\mathrm{L}^{p}.$
3 Vector valued extensions of operators on convex body valued functions
The aim of this section is to prove vector valued estimates for convex body valued functions with matrix weights, as well as to apply it to vector valued estimates for the matrix weighted maximal operators for vector valued functions (see Subsection 3.1 for the definitions and details). This will be a consequence of an extrapolation result on spaces of functions with values on $\mathcal{K}_{\mathrm{bcs}}$ and it is a minor modification of the analogous result for matrix weights and vector valued functions due to Bownik and Cruz-Uribe (see [Cruz_Uribe_Bownik_Extrapolation, Section 9]). Before stating the extrapolation theorem, we introduce the concept of families of extrapolation pairs following the convention in [Book_Extrapolation]. A family of extrapolation pairs $\mathcal{F}$ is a collection of pairs $(F,G)$ of measurable functions taking values in $\mathcal{K}_{\mathrm{bcs}}$ such that neither $F$ nor $G$ is equal to $\{0\}$ almost everywhere. In addition, given such a family $\mathcal{F},$ we call each element $(F,G)\in\mathcal{F}$ an extrapolation pair. We are interested in inequalities of the form
$\left\|{F}\right\|_{{\mathrm{L}^{p}(W)}}\leq C\left\|{G}\right\|_{{\mathrm{L}^{p}(W)}},\qquad(F,G)\in\mathcal{F},$
in the sense that this holds for all pairs $(F,G)\in\mathcal{F}$ for which the left-hand side is finite and with the constant $C$ depending on the characteristic $[W]_{A_{p}}$ but not on the particular weight $W.$
Theorem A (Bownik, Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation]).
Consider a family of extrapolation pairs $\mathcal{F}.$ Suppose that for some $p_{0},$ $1\leq p_{0}\leq\infty,$ there exists an increasing function $C_{p_{0}}$ such that for every matrix weight $W\in A_{p_{0}}$ it holds that
$\left\|{F}\right\|_{{\mathrm{L}^{p_{0}}(W)}}\leq C_{p_{0}}([W]_{A_{p_{0}}})\left\|{G}\right\|_{{\mathrm{L}^{p_{0}}(W)}},\qquad(F,G)\in\mathcal{F}.$ (3.1)
Then, for all $1<p<\infty$ and for every matrix weight $W\in A_{p}$ it holds that
$\left\|{F}\right\|_{{\mathrm{L}^{p}(W)}}\leq C_{p}(p_{0},n,d,[W]_{A_{p}})\left\|{G}\right\|_{{\mathrm{L}^{p}(W)}},\qquad(F,G)\in\mathcal{F},$
where
$C_{p}(p_{0},n,d,[W]_{A_{p}})=C(p,p_{0})C_{p_{0}}\left(C(n,d,p,p_{0})[W]_{A_{p}}^{\max\left\{1,\frac{p_{0}-1}{p-1}\right\}}\right)$
if $p_{0}<\infty,$ and
$C_{p}(p_{0},n,d,[W]_{A_{p}})=C(p,p_{0})C_{p_{0}}\left(C(n,d,p,p_{0})[W]_{A_{p}}^{\frac{1}{p-1}}\right)$
if $p_{0}=\infty.$
The proof of this result is exactly the same as in [Cruz_Uribe_Bownik_Extrapolation, Theorem 9.1], up to the point that the calculations are done directly with the given convex body valued functions $F$ and $G,$ instead of taking $F(x)=\{\lambda f(x)\colon|\lambda|\leq 1\}$ and $G(x)=\{\lambda g(x)\colon|\lambda|\leq 1\}$ for given vector valued functions $f$ and $g$ (also note that in [Cruz_Uribe_Bownik_Extrapolation] $F(x)$ and $G(x)$ are segments with endpoints ±$\pm f(x)$ and ±$\pm g(x)$ because in that article real symmetric convex bodies are used, instead of complex symmetric). Since the proof in [Cruz_Uribe_Bownik_Extrapolation] follows closely that of the scalar case, we also refer the reader to [Book_Extrapolation, Chapter I.3] for a detailed exposition on the topic in the classical setting.
The main result of this section follows from Theorem A. This is a vector valued estimate for sequences $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ of a given family of extrapolation pairs $\mathcal{F}.$
Theorem 5 (Vector valued estimates for families of extrapolation pairs).
Consider a family of extrapolation pairs $\mathcal{F}.$ Suppose that for some $1\leq p_{0}\leq\infty$ there exists an increasing function $C_{p_{0}}$ such that for every matrix weight $W\in A_{p_{0}}$ inequalities (3.1) hold. Then, for any $p$ and $q,$ $1<p,q<\infty,$ and for every matrix weight $W\in A_{p}$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,p_{0},q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}G_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ and
$C(n,d,p_{0},p,q,[W]_{A_{p}})=$ $C(d,p,q)$
$C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q-1}\right\}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)$
if $p_{0}<\infty,$ and
$C(n,d,p_{0},p,q,[W]_{A_{p}})=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\frac{1}{q-1}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)$
if $p_{0}=\infty.$
Recall that one can define the Minkowski addition of $A,B\in\mathcal{K}_{\mathrm{bcs}}$ (and for subsets of $\mathbb{C}^{d}$ in general) by
$A+B\coloneq\{a+b\colon a\in A,b\in B\}.$
It is easy to check that if $A,B\in\mathcal{K}_{\mathrm{bcs}},$ then also $A+B\in\mathcal{K}_{\mathrm{bcs}}.$ One can extend the Minkowski addition of two convex body to that of $N$ convex bodies by induction, with $N\geq 2.$ We can also define the scalar multiplication, given $A\in\mathcal{K}_{\mathrm{bcs}}$ and $\lambda\in\mathbb{C},$ by
$\lambda A\coloneq\{\lambda a\colon a\in A\},$
and also $\lambda A\in\mathcal{K}_{\mathrm{bcs}}.$ Keep in mind that these operations do not define a vector space structure, since the Minkowski addition has no inverse.
Given $1\leq q<\infty$ and a sequence $\{K_{n}\}_{n=1}^{\infty}\subseteq\mathcal{K}_{\mathrm{bcs}}$ such that
$\sum_{n=1}^{\infty}|K_{n}|^{q}<\infty,$ (3.2)
we define the infinite $\ell^{q}$ Minkowski addition of $\{K_{n}\},$ denoted by $\Sigma_{q}(\{K_{n}\}),$ as
$\Sigma_{q}(\{K_{n}\})\coloneq\bigcup\left\{\sum_{n=1}^{\infty}a_{n}v_{n}\colon v_{n}\in K_{n}\text{ for }n\geq 1,\{a_{n}\}\in\ell^{q^{\prime}},\left\|{\{a_{n}\}}\right\|_{{\ell^{q^{\prime}}}}\leq 1\right\}.$ (3.3)
The sums in (3.3) are convergent due to the summability condition (3.2) and Hölder’s inequality. In particular, it holds that
$\left|\Sigma_{q}(\{K_{n}\})\right|\leq\left(\sum_{n=1}^{\infty}|K_{n}|^{q}\right)^{1/q}$
for any $1\leq q<\infty$ whenever the right-hand side of this expression is finite. Furthermore, given a $(d\times d)$ matrix $A,$ it also holds by linearity that
$A\Sigma_{q}(\{K_{n}\})=\Sigma_{q}(\{AK_{n}\}).$
For $1\leq q<\infty$ it also happens that $\Sigma_{q}(\{K_{n}\})$ is a convex symmetric bounded closed set, which is proved in next lemma.
Lemma 6.
Let $1\leq q<\infty$ and let $\{K_{m}\}^{\infty}_{m=1}$ be a sequence of sets in $\mathcal{K}_{\mathrm{bcs}}$ with
$\sum_{m=1}^{\infty}|K_{m}|^{q}<\infty.$
Then $K=\Sigma_{q}(\{K_{n}\})$ is a well-defined set in $\mathcal{K}_{\mathrm{bcs}}$ with
$|K|\sim_{d,q}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
Proof.
First of all, for all sequences $\{v_{m}\}_{m=1}^{\infty}$ with $v_{m}\in K_{m},$ for all $m=1,2,\ldots$ and for all complex numbers $a_{1},a_{2},\ldots$ with $\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\leq 1$ we have by Hölder’s inequality
$\sum_{m=1}^{\infty}|a_{m}v_{m}|\leq\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\left(\sum_{m=1}^{\infty}|v_{m}|^{q}\right)^{1/q}\leq\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}<\infty.$
This shows that $K$ is a well-defined bounded set with $|K|\leq\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
It is obvious that $K$ is symmetric. Moreover, a standard weak star compactness argument yields that $K$ is closed. Let us check that $K$ is convex. For any sequences $\{v_{m}\}_{m=1}^{\infty},$$\{u_{m}\}_{m=1}^{\infty}$ with $v_{m},u_{m}\in K_{m},$ for all $m=1,2,\ldots,$ any complex numbers $a_{1},a_{2},\ldots$ and $b_{1},b_{2},\ldots$ with $\left\|{\{a_{m}\}}\right\|_{{q^{\prime}}}\leq 1$ and $\left\|{\{b_{m}\}}\right\|_{{q^{\prime}}}\leq 1,$ and any $t\in(0,1),$ we can write
$t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}$
$=\sum_{m=1}^{\infty}(t|a_{m}|+(1-t)|b_{m}|)$
$\cdot\left(\frac{t|a_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{a_{m}}{|a_{m}|}v_{m}+\frac{(1-t)|b_{m}|}{t|a_{m}|+(1-t)|b_{m}|}\cdot\frac{b_{m}}{|b_{m}|}u_{m}\right),$
with the obvious modifications if $a_{m}=0$ or $b_{m}=0$ for some $m\geq 1,$ thus it is clear that $t\sum_{m=1}^{\infty}a_{m}v_{m}+(1-t)\sum_{m=1}^{\infty}b_{m}u_{m}\in K.$
We will now show that
$|K|\gtrsim_{d,q}\frac{1}{\sqrt{d}}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
We can pick a sequence $\{v_{m}\}_{m=1}^{\infty}$ with $v_{m}\in K_{m}$ and $|v_{m}|\geq\frac{1}{2}|K_{m}|,$ for all $m=1,2,\ldots.$ It is clear that there exists a positive integer $N$ with
$\sum_{m=1}^{N}|v_{m}|^{q}\geq\frac{1}{2}\sum_{m=1}^{\infty}|v_{m}|^{q}.$
For any vector $x\in\mathbb{C}^{d},$ we denote by $x^{1},\ldots,x^{d}$ its coordinates. Notice that
$\sum_{m=1}^{N}|v_{m}|^{q}\sim_{d,q}\sum_{m=1}^{N}\sum_{j=1}^{d}|v_{m}^{j}|^{q}=\sum_{j=1}^{d}\sum_{m=1}^{N}|v_{m}^{j}|^{q},$
therefore there exists $j\in\{1,\ldots,d\}$ with
$\sum_{m=1}^{N}|v_{m}^{j}|^{q}\gtrsim_{d,q}\sum_{m=1}^{N}|v_{m}|^{q}.$
Clearly, one can find complex numbers $a_{1},\ldots,a_{N}$ with $\left\|{\{a_{m}\}_{m=1}^{N}}\right\|_{{q^{\prime}}}\leq 1$ such that
$\sum_{m=1}^{N}a_{m}v_{m}^{j}=\left(\sum_{m=1}^{N}|v_{m}^{j}|^{q}\right)^{1/q}.$
It follows that
$\left|\sum_{m=1}^{N}a_{m}v_{m}\right|$ $\geq\left|\sum_{m=1}^{N}a_{m}v_{m}^{j}\right|=\left(\sum_{m=1}^{N}|v_{m}^{j}|^{q}\right)^{1/q}$
$\gtrsim_{d,q}\left(\sum_{m=1}^{N}|v_{m}|^{q}\right)^{1/q}\gtrsim_{q}\left(\sum_{m=1}^{\infty}|K_{m}|^{q}\right)^{1/q}.$
Since $\sum_{m=1}^{N}a_{m}v_{m}=\sum_{m=1}^{N}a_{m}v_{m}+\sum_{m=N+1}^{\infty}0\cdot 0\in K,$ the proof is complete. ∎
We are now ready to prove Theorem 5.
Proof of Theorem 5.
Fix first $1<q<\infty.$ We construct a new family $\mathcal{F}_{q}$ of convex body valued functions as follows. For each sequence $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ such that
$\sum_{n=1}^{\infty}|F_{n}(x)|^{q}<\infty\quad\text{ and }\quad\sum_{n=1}^{\infty}|G_{n}(x)|^{q}<\infty$ (3.4)
at almost every $x\in\mathbb{R}^{n},$ we define
$F(x)=\Sigma_{q}(\{F_{n}(x)\}),\qquad G(x)=\Sigma_{q}(\{G_{n}(x)\}).$
Observe that by lemma 6, both $F$ and $G$ are convex body valued functions with
$|W(x)^{1/p}F(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}F_{n}(x)|^{q}\right)^{1/q}$
and
$|W(x)^{1/p}G(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}G_{n}(x)|^{q}\right)^{1/q}$
at almost every $x\in\mathbb{R}^{n},$ for any $1<p<\infty.$ It is also easy to see that the convex body valued functions $F,G$ are measurable.
Next, note that the family of extrapolation pairs $\mathcal{F}_{q}$ that we have just constructed satisfies inequalities (3.1) with exponent $q$ for any matrix weight $W\in A_{q}.$ Indeed, for any given matrix weight $W\in A_{q},$ due to Lemma 6 it holds that
$\begin{split}\left\|{F}\right\|_{{\mathrm{L}^{q}(W)}}^{q}&=\int_{\mathbb{R}^{n}}|W^{1/q}F(x)|^{q}\,\,\mathrm{dm}(x)\\ &=\int_{\mathbb{R}^{n}}\left|\Sigma_{q}(\{W(x)^{1/q}F_{n}(x)\})\right|^{q}\,\,\mathrm{dm}(x)\\ &\leq C(d,q)\sum_{n=1}^{\infty}\int_{\mathbb{R}^{n}}|W(x)^{1/q}F_{n}(x)|^{q}\,\,\mathrm{dm}(x)\\ &\leq C_{q}([W]_{q})^{q}\sum_{n=1}^{\infty}\int_{\mathbb{R}^{n}}|W(x)^{1/q}G_{n}(x)|^{q}\,\,\mathrm{dm}(x)\\ &\leq C(d,q)C_{q}([W]_{q})^{q}\int_{\mathbb{R}^{n}}\left|\Sigma_{q}(\{W(x)^{1/q}G_{n}(x)\})\right|^{q}\,\,\mathrm{dm}(x)\\ &=C_{q}([W]_{q})^{q}\left\|{G}\right\|_{{\mathrm{L}^{q}(W)}}^{q}.\end{split}$
Here we have used Theorem A with the extrapolation paris $(F_{n},G_{n})\in\mathcal{F}$ and with exponent $q,$ so that we have
$C_{q}([W]_{A_{q}})=C(d,q,p_{0})C_{p_{0}}\left(C(n,d,q,p_{0})[W]_{A_{q}}^{\max\left\{1,\frac{p_{0}-1}{q-1}\right\}}\right).$
We have just seen that the family of extrapolation pairs $\mathcal{F}_{q}$ satisfies (3.1), therefore we can apply Theorem A and Lemma 6 to get that for any $1<p<\infty$ and any matrix weight $W\in A_{p}$ it holds that
$\begin{split}\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}F_{n}(x)|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}&\leq C(d,p)\left\|{W^{1/p}F}\right\|_{{\mathrm{L}^{p}}}\\ \leq C_{p,q}(&[W]_{A_{p}})\left\|{W^{1/p}G}\right\|_{{\mathrm{L}^{p}}}\\ \leq C_{p,q}(&[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}G_{n}(x)|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\end{split}$
for every sequence $\{(F_{n},G_{n})\}\subseteq\mathcal{F}$ satisfying (3.4) and where
$\begin{split}C_{p,q}([W]_{A_{p}})&=C(d,p)C_{p,q}\left(C(n,d,p,q)[W]_{A_{p}}^{\max\left\{1,\frac{q-1}{p-1}\right\}}\right)\\ &=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q-1}\right\}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)\end{split}$
if $p_{0}<\infty,$ and
$\begin{split}C_{p,q}([W]_{A_{p}})&=C(d,p)C_{p,q}\left(C(n,d,p,q)[W]_{A_{p}}^{\max\left\{1,\frac{q-1}{p-1}\right\}}\right)\\ &=C(d,p,q)C_{p_{0}}\left(C(n,d,p_{0},p,q)[W]_{A_{p}}^{\left(\frac{1}{q-1}\max\left\{1,\frac{q-1}{p-1}\right\}\right)}\right)\end{split}$
if $p_{0}=\infty,$ with the additional factor $C(d,p,q)$ being in both cases due to the various applications of Lemma 6. ∎
3.1 Fefferman–Stein vector valued inequalities for weighted maximal operators
Next, we give an application of Theorem 5. Fix $1<p<\infty.$ Given a $(d\times d)$ matrix valued weight $W$ on $\mathbb{R}^{n},$ one can define the pointwise matrix weighted maximal operator for vector valued functions by
$M_{W}f(x)\coloneq\sup_{Q}\frac{1}{|Q|}\int_{Q}|W(x)^{1/p}f(y)|\,\,\mathrm{dm}(y)\mathbf{1}_{Q}(x),$
where $f$ is a locally integrable function taking values on $\mathbb{C}^{d}$ and where the supremum is taken over all cubes with sides parallel to the coordinate axes. The maximal operator $M_{W}$ is also called Christ–Goldberg maximal operator, since it was studied by these two authors (see [Christ_Goldberg_2001] and [Goldberg_2003]). One can also define a weighted maximal operator with reducing operators as
$\widetilde{M}_{W}f(x)\coloneq\sup_{Q}\int_{Q}|\mathcal{W}_{Q}f(y)|\,\,\mathrm{dm}(y)\mathbf{1}_{Q}(x),$
where $f$ is a locally integrable function with values on $\mathbb{C}^{d},$ the supremum is taken over bounded cubes with sides parallel to the axes and where $\mathcal{W}_{Q}$ is the reducing operator of $W$ over $Q$ with exponent $p.$ We do not make the dependence on the exponent $p$ explicit since it will always be clear from the context.
Both weighted maximal operators play an important role in the theory of matrix weighted norm inequalities. In addition, whenever $W$ is a matrix $A_{p}$ weight, both maximal operators are bounded from weighted $\mathrm{L}^{p}(W)$ to unweighted $\mathrm{L}^{p}.$ Isralowitz and Moen [Isralowitz_Moen_2019] proved that
$\left\|{M_{W}}\right\|_{{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}}}\lesssim_{n,d,p}[W]_{A_{p}}^{1/(p-1)},$ (3.5)
while Isralowitz, Kwon and Pott [Isralowitz_Kwon_Pott2017] showed that
$\left\|{\widetilde{M}_{W}}\right\|_{{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}}}\lesssim_{n,d,p}[W]_{A_{p}}^{1/(p-1)}.$
We will use a trick due to Bownik and Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation] so as to use Theorem 5 (which is a statement about convex body valued functions) to show statements about $\mathbb{C}^{d}$ vector valued functions. In order to apply the previous results for convex body valued operators to these weighted maximal operators, we will also need to define the convex body valued analogue of $M_{W}.$ Both definitions will be related by the following correspondence between vector valued and convex body valued functions. Given a vector valued function $f,$ that is taking values on $\mathbb{C}^{d},$ we define the convex body valued function $F$ by
$F(x)\coloneq\{\lambda f(x)\colon|\lambda|\leq 1\}.$ (3.6)
By construction, $F$ takes values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and also $|F(x)|=|f(x)|.$ Thus, any estimate for operators depending on $|f(x)|$ can be studied through estimates for an analogous operator acting on $F(x)$ and depending as well on $|F(x)|$ uniquely. It is straightforward to check that, for a locally integrable vector valued function $f$ and a cube $Q,$ it holds that the averaging operator $A_{Q}$ applied to $F$ given by (3.6) can be computed as
$A_{Q}F(x)=\left\{\frac{1}{|Q|}\int_{Q}k(y)f(y)\,\,\mathrm{dm}(y)\colon k\in\mathrm{L}^{\infty},\left\|{k}\right\|_{{\mathrm{L}^{\infty}}}\leq 1\right\}\mathbf{1}_{Q}(x).$ (3.7)
For this reason, we will abuse notation and denote $A_{Q}f(x)=A_{Q}F(x).$ Also note that, by an appropriate choice of the function $k$ in (3.7), one can see that
$|A_{Q}f(x)|\sim_{d}\frac{1}{|Q|}\int_{Q}|f(y)|\,\,\mathrm{dm}(y),$ (3.8)
and more generally
$|A_{Q}F(x)|\sim_{d}\frac{1}{|Q|}\int_{Q}|F(y)|\,\,\mathrm{dm}(y)$
for convex body valued functions in general, which can be seen by choosing an appropriate selection function. For fixed $1<p<\infty$ and given a $(d\times d)$-matrix weight $W,$ the convex body valued analogue of $M_{W}$ acting on a locally integrably bounded convex body valued function $F$ is defined by
$M_{W}^{\mathcal{K}}F(x)\coloneq\overline{\mathrm{conv}\left(\bigcup_{Q}A_{Q}(W(x)^{1/p}F(x))\right)},$
where the union is taken over all bounded cubes with sides parallel to the axes. Moreover, given a locally integrable vector valued function $f,$ we denote $M_{W}^{\mathcal{K}}f(x)=M_{W}^{\mathcal{K}}F(x),$ where $F$ is the convex body valued function given by (3.6). In addition, if $W$ is a matrix weight and $f$ a vector valued function, linearity and (3.8) yield that
$|M_{W}^{\mathcal{K}}f(x)|=|W(x)^{1/p}M^{\mathcal{K}}f(x)|\sim_{d}M_{W}f(x),$ (3.9)
where $M^{\mathcal{K}}$ denotes the maximal operator weighted by the $(d\times d)$-identity matrix.
Theorem 1.
Consider a $(d\times d)$ matrix weight $W$ and a sequence of vector valued functions $\{f_{n}\}.$ Then, for each $1<p,q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Proof.
In order to apply Theorem 5, we need to construct a family $\mathcal{F}$ of extrapolation pairs for which the left-hand side of (3.1) is finite. To this end, we will restrict ourselves to vector valued functions $f\in\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ (compactly supported essentially bounded functions), and a density argument will yield the conclusion for $f\in\mathrm{L}^{p}(\mathbb{R}^{n};\mathbb{C}^{d}).$ For each vector valued function $f\in\mathrm{L}^{\infty}_{\mathrm{c}},$ we consider the extrapolation pair $(M^{\mathcal{K}}F,F),$ where $F$ is the convex body valued function defined by (3.6).
Consider the given $1<q<\infty.$ It is clear that the left-hand side of (3.1) is finite for every pair $(M^{\mathcal{K}}F,F)\in\mathcal{F}$ for $p_{0}=q,$ and that (3.1) is satisfied if we take $C_{p_{0}}(t)=C_{q}(t)=C(n,d,q)t^{\frac{1}{q-1}}.$ Indeed, because of (3.9), for any matrix $A_{q}$ weight $W$ we have that
$\left\|{M^{\mathcal{K}}F}\right\|_{{\mathrm{L}^{q}(W)}}\sim_{d}\left\|{M_{W}f}\right\|_{{\mathrm{L}^{q}}}\leq C(n,d,q)[W]_{A_{q}}^{1/(q-1)}\left\|{f}\right\|_{{\mathrm{L}^{q}(W)}}<\infty,$
where we have used estimate (3.5) and that $\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})\subseteq\mathrm{L}^{q}(\mathbb{R}^{n};\mathbb{C}^{d}).$
Now, Theorem 5 gives that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}M^{\mathcal{K}}F_{n}|^{q_{1}}\right)^{1/{q_{1}}}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p_{0},p,q_{1},[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q_{1}}\right)^{1/{q_{1}}}}\right\|_{{\mathrm{L}^{p}}}$ (3.10)
for any $1<p,q_{1}<\infty$ (with $q_{1}$ possibly different of $q=p_{0}$) and any matrix weight $W\in A_{p},$ where $\{(M^{\mathcal{K}}F_{n},F_{n})\}$ is a sequence in $\mathcal{F}$ and where
$C(n,d,p_{0},p,q_{1},[W]_{A_{p}})$
$=C(d,p)C_{p_{0}}\left(C(n,d,p_{0},p,q_{1})[W]_{A_{p}}^{\left(\max\left\{1,\frac{p_{0}-1}{q_{1}-1}\right\}\max\left\{1,\frac{q_{1}-1}{p-1}\right\}\right)}\right).$
In particular, if we restrict our attention to $q_{1}=p_{0}=q,$ we can take
$C(n,d,q,p,[W]_{A_{p}})=C(n,d,q,p)[W]_{A_{p}}^{\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Also observe that for each sequence $\{(M^{\mathcal{K}}F_{n},F_{n})\}\subseteq\mathcal{F},$ we can choose a sequence $\{f_{n}\}\subseteq\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ such that $F_{n}$ is obtained from $f_{n}$ using (3.6) for every $n\geq 1.$ On one hand observe that using that $|W^{1/p}F_{n}|=|W^{1/p}f_{n}|$ for the sequence $\{f_{n}\}\subseteq\mathrm{L}^{\infty}_{\mathrm{c}}(\mathbb{R}^{n};\mathbb{C}^{d})$ that we chose previously, we get that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}=\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$ (3.11)
On the other hand, using (3.9), we see that
$\left\|{\left(\sum_{n=1}^{\infty}|W^{1/p}M^{\mathcal{K}}F_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\sim_{d}\left\|{\left(\sum_{n=1}^{\infty}|M_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$ (3.12)
Finally, replacing (3.11) and (3.12) into (3.10), we get the desired vector valued estimates for the pointwise weighted maximal operator, as we wanted to show. ∎
The analogous result for the operator $\widetilde{M}_{W}$ is a consequence of Theorem 1.
Theorem 7.
Consider a $(d\times d)$ matrix weight $W$ and a sequence of vector valued functions $\{f_{n}\}.$ Then, for each $1<p,q<\infty$ it holds that
$\left\|{\left(\sum_{n=1}^{\infty}|\widetilde{M}_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(n,d,p,q,[W]_{A_{p}})\left\|{\left(\sum_{n=1}^{\infty}|W(x)^{1/p}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}},$
where
$C(n,d,p,q,[W]_{A_{p}})=C(n,d,p,q)[W]_{A_{p}}^{\frac{1}{p}+\max\left\{\frac{1}{q-1},\frac{1}{p-1}\right\}}.$
Proof.
Just observe that
$\begin{split}\widetilde{M}_{W}f(x)&=\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|\mathcal{W}_{Q}f(y)|\,\,\mathrm{dm}(y)\\ &\sim_{d}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|Q|}\int_{Q}|W(z)^{1/p}f(y)|^{p}\,\,\mathrm{dm}(z)\right)^{1/p}\,\,\mathrm{dm}(y)\\ &\leq C(d,p)[W]_{A_{p}}^{1/p}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}\frac{1}{|Q|}\int_{Q}|W(z)^{1/p}f(y)|\,\,\mathrm{dm}(z)\,\,\mathrm{dm}(y)\\ &\leq C(d,p)[W]_{A_{p}}^{1/p}\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}M_{W}f(z)\,\,\mathrm{dm}(z)\\ &=C(d,p)[W]_{A_{p}}^{1/p}M(M_{W}f)(x).\end{split}$
Therefore, we get that
$\left\|{\left(\sum_{n=1}^{\infty}|\widetilde{M}_{W}f_{n}|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}\leq C(d,p)[W]_{A_{p}}^{1/p}\left\|{\left(\sum_{n=1}^{\infty}|M(M_{W}f_{n})|^{q}\right)^{1/q}}\right\|_{{\mathrm{L}^{p}}}.$
An application of the classical Fefferman–Stein vector valued inequalities for the maximal operator followed by the use of Proposition 1 yields the desired result. ∎
4 Two matrix weighted biparameter product BMO
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$
Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We emphasize that $\mathcal{E}$ stands for the set of biparameter signatures, $\mathcal{E}=\mathcal{E}^{1}\times\mathcal{E}^{2},$ where $\mathcal{E}^{1}=\{0,1\}^{n}\setminus\{(1,\ldots,1)\}$ and $\mathcal{E}^{2}=\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$ We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2},$
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure, and all reducing operators are taken with respect to exponent $p.$ Note that
$|\mathcal{V}_{R}P\mathcal{U}^{-1}_{R}|=|\mathcal{U}^{-1}_{R}P^{\ast}\mathcal{V}_{R}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P^{\ast}(\mathcal{V}^{\prime}_{R})^{-1}|,\qquad\forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall P\in\mathrm{M}_{d}(\mathbb{C}).$
Therefore
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sim_{p,d}\|B^{\ast}\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(V^{\prime},U^{\prime},p^{\prime})}.$
4.1 $\mathrm{H}^{1}$-BMO duality
The main goal of this subsection is to prove Theorem 2. We split the proof in Propositions 8 and 9 below, each treating one direction of Theorem 2.
We define $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ as the set of all sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ such that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\coloneq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}<\infty.$
This is the direct biparameter analog of the one-parameter two matrix weighted $\mathrm{H}^{1}$ norm defined in [Isralowitz_2017]. It is easy to check that $(\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),\|\,\cdot\,\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)})$ is a Banach space.
Proposition 8.
Let $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p).$ Then, the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C},$
$\ell_{B}(\Phi)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$
is well-defined bounded with $\|\ell_{B}\|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}.$
Proof.
We adapt the first half of the proof of [Isralowitz_2017, Theorem 1.3]. Let $\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ be arbitrary. We show that the sum
$(\Phi,B)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})$
converges absolutely with
$|(\Phi,B)|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}.$
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})|=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathrm{tr}(\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R})|$
$\lesssim_{d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}|\leq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{U}_{R}^{-1}(B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}|$
$\lesssim_{p,d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|.$
Set now
$F\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}$
and define
$\Omega_{k}\coloneq\{F>2^{k}\},\qquad k\in\mathbb{Z},$
$\mathcal{B}_{k}\coloneq\left\{R\in\bm{\mathcal{D}}:\leavevmode\nobreak\ |R\cap\Omega_{k+1}|\leq\frac{1}{2}|R|<|R\cap\Omega_{k}|\right\},\qquad k\in\mathbb{Z},$
$\widetilde{\Omega}_{k}\coloneq\left\{M_{\bm{\mathcal{D}}}(\mathbf{1}_{\Omega_{k}})>\frac{1}{2}\right\},\qquad k\in\mathbb{Z}.$
Clearly $\Omega_{k}\subseteq\widetilde{\Omega}_{k}$ up to a set of zero measure, and in fact $|\widetilde{\Omega}_{k}|\sim|\Omega_{k}|,$ for all $k\in\mathbb{Z}.$ It is also obvious that $R\subseteq\widetilde{\Omega_{k}},$ for all $R\in\mathcal{B}_{k},$ for all $k\in\mathbb{Z}.$ If $R\in\bm{\mathcal{D}}$ with $\Phi_{R}^{\varepsilon}\neq 0$ for some $\varepsilon\in\mathcal{E}$ and $|R|\geq 2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow-\infty$ we deduce
$|R|\geq 2|R\cap\{F>0\}|=2|R|,$
since $F>0$ a. e. on $R,$ contradiction. If $R\in\bm{\mathcal{D}}$ with $|R|<2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow\infty$ we deduce
$|R|\leq 2|R\cap\{F=\infty\}|=0,$
since by assumption $F\in\mathrm{L}^{1}(\mathbb{R}^{n+m}),$ contradiction. It follows that for every $R\in\bm{\mathcal{D}}$ with $\Phi_{R}\neq 0$ for some $\varepsilon\in\mathcal{E}$ there exists $k\in\mathbb{Z}$ such that $R\in\mathcal{B}_{k}.$ Thus, we have
$|(\Phi,B)|$ $\lesssim_{p,d}\sum_{k\in\mathbb{Z}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|$
$\leq\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2}$
$\leq\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\text{sh}(\mathcal{B}_{k})|^{1/2}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\widetilde{\Omega}_{k}|^{1/2}.$
We will now prove that
$\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{4/p}2^{2k}|\widetilde{\Omega}_{k}|,\qquad\forall k\in\mathbb{Z}.$ (4.1)
This will be enough to conclude the proof, because assuming it we will get
$\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\widetilde{\Omega}_{k}|^{1/2}\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\widetilde{\Omega}_{k}|$
$\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\Omega_{k}|\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|F\|_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}.$
Fix now $k\in\mathbb{Z}.$ We begin by noticing that
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\leq 2^{2k+2}|\widetilde{\Omega}_{k}\setminus\Omega_{k+1}|\leq 2^{2k+2}|\widetilde{\Omega}_{k}|,$
by the definition of $\Omega_{k+1}.$ Moreover, we have
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\geq\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\,\mathrm{dm}(x)$
$=\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\,\mathrm{dm}(x)$
$\sim_{d}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\,\mathrm{dm}(x)$
$\geq\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{|R\setminus\Omega_{k+1}|}{|R|}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}.$
Let $v\in\mathbb{C}^{d}$ be arbitrary. Consider the function
$w\coloneq|V^{-1/p}v|^{p^{\prime}}.$
Then, $w$ is a scalar $\bm{\mathcal{D}}$-dyadic biparameter $A_{p^{\prime}}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with
$[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}\lesssim_{p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{p^{\prime}/p},$
because $V^{-1/p}$ is a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight (see for example Lemma 3.2 in [DKPS2024]). It is then well-known that
$[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}$
and
$\langle w\rangle_{R}^{1/p^{\prime}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}\langle w^{1/p^{\prime}}\rangle_{R},\quad\forall R\in\bm{\mathcal{D}}$
(see for example Subsection 2.3.3 in [Kakaroumpas_Soler_2022]). Using Jensen’s inequality and the definition of the $A_{2}$ characteristic, it follows that
$\frac{w^{1/p^{\prime}}(R\setminus\Omega_{k+1})}{w^{1/p^{\prime}}(R)}\geq[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}^{-1}\frac{|R\setminus\Omega_{k+1}|}{|R|}\geq[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}^{-1}\cdot\frac{1}{2},$
so
$\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}v|\,\mathrm{dm}(x)\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-1/p}\int_{R}|V(x)^{-1/p}v|\,\mathrm{dm}(x).$
Thus
$\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-2/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{p^{\prime}}\,\mathrm{dm}(x)\right)^{2/p^{\prime}}$
$\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\sim_{d}[V]_{A_{p},\bm{\mathcal{D}}}^{-4/p}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2},$
concluding the proof. ∎
Before we proceed to the second half of Theorem 2, recall that the strong dyadic Christ–Goldberg maximal function corresponding to a weight $W$ on $\mathbb{R}^{n+m}$ (and exponent $p$) is defined as
$M_{\bm{\mathcal{D}},W}f(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|W(x)^{1/p}f|\rangle_{R}\mathbf{1}_{R}(x),\qquad x\in\mathbb{R}^{n+m},\leavevmode\nobreak\ f\in\mathrm{L}^{1}_{\mathrm{loc}}(\mathbb{R}^{n+m};\mathbb{C}^{d}).$
Because of [Vuorinen_2023, Theorem 1.3] we have that the operator $M_{\bm{\mathcal{D}},W}$ is bounded when acting on $L^{p}(W)\rightarrow L^{p}(\mathbb{R}^{n+m})$ (note that the target space is unweighted) provided that $[W]_{A_{p},\bm{\mathcal{D}}}<\infty,$ specifically one has the bound
$\|M_{\bm{\mathcal{D}},W}\|_{\mathrm{L}^{p}(W)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p}}^{2/(p-1)},$ (4.2)
for $1<p<\infty.$
Proposition 9.
Let $\ell$ be any bounded linear functional on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$ Then, there exists a unique $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with
$\ell(\Phi)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon}),\quad\forall\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p).$
Moreover, there holds
$\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}\|\ell\|.$
Proof.
We partially adapt the second half of the proof of [Isralowitz_2017, Theorem 1.3]. We denote by $\mathcal{H}$ the set of all $\mathrm{M}_{d}(\mathbb{C})$-valued $\mathrm{L}^{2}$ functions on $\mathbb{R}^{n+m}$ with finitely many nonzero biparameter cancellative Haar coefficients, and we identify such functions with finitely supported sequences $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C})$ in the obvious way. Then, for all $\Phi\in\mathcal{H},$ we have
$\Phi=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}$
and therefore
$\ell(\Phi)$ $=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}\ell(h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)E_{ij})=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\sum_{i,j=1}^{d}h_{R}^{\varepsilon}\Phi_{R}^{\varepsilon}(i,j)\ell(E_{ij})$
$=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})=(\Phi,B),$
where $E_{ij}$ is the $d\times d$-matrix with 1 at the $(i,j)$-entry and 0 at all other entries, $A(i,j)$ is the $(i,j)$-entry of a matrix $A,$ and
$B_{R}^{\varepsilon}(i,j)\coloneq\overline{\ell(h_{R}^{\varepsilon}E_{ji})},\qquad\forall i,j=1,\ldots,d,\leavevmode\nobreak\ \forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall\varepsilon\in\mathcal{E}.$
By Proposition 8 and since $\mathcal{H}$ is dense in $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),$ it suffices only to prove that the defined sequence $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ is in $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)$ with
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}\|\ell\|.$ (4.3)
Note that by the scalar, unweighted BMO equivalences we have
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}$
$\sim_{n,m,p,d}\sup_{\Omega}\frac{1}{|\Omega|^{1/p^{\prime}}}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m})}=:\mathcal{C},$
so it suffices only to prove that $\mathcal{C}\lesssim_{n,m,p,d}\|\ell\|.$ By the Monotone Convergence Theorem, we can without loss of generality assume that $B$ has only finitely many nonzero terms.
[…]
$\langle A,B\rangle\coloneq\mathrm{tr}(B^{\ast}A).$
[…]
$\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
$\sim_{n,m,p,d}\left\{\left|\int_{\mathbb{R}^{n+m}}\langle F(x),G(x)\rangle\,\mathrm{dm}(x)\right|:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\right\}$
=$=\bigg{\{}\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((G^{\varepsilon}_{R})^{\ast}F_{R}^{\varepsilon})\bigg{|}:\leavevmode\nobreak\ G\in\mathcal{H},\leavevmode\nobreak\ \|G\|_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}=1\bigg{\}},$ (4.4)
for all $F\in\mathcal{H}.$ We recall also the usual (unweighted) dyadic Littlewood–Paley estimates
$\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
[…]
Fix now any Lebesgue-measurable subset $\Omega$ of $\mathbb{R}^{n+m}$ with finite nonzero measure. Then, we have
$\bigg{(}\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p^{\prime}/2}\,\mathrm{dm}(x)\bigg{)}^{1/p^{\prime}}$
(4.1)${\sim}_{n,m,p,d}\bigg{\|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}h_{R}^{\varepsilon}\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
[…]
$=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}}\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}\left((B_{R}^{\varepsilon})^{\ast}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\right)\bigg{|}$
$=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|(\hat{A},B)|=\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}|\ell(\hat{A})|$
$\leq\|\ell\|\sup_{A\in\mathcal{H}\setminus\{0\}}\frac{1}{\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))}}\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\hat{A}^{\varepsilon}_{R}\coloneq\begin{cases}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1},\text{ if }R\in\bm{\mathcal{D}}(\Omega)\\ \\ 0,\text{ otherwise}\end{cases},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{2};\mathrm{M}_{d}(\mathbb{C}))},$
for all $A\in\mathcal{H}.$ We set
$N_{\Omega}(x)\coloneq\sup_{R\in\bm{\mathcal{D}}(\Omega)}|V(x)^{-1/p}\mathcal{V}_{R}|\mathbf{1}_{R}(x),\quad x\in\mathbb{R}^{n+m}.$
We define now functions $\widetilde{N}_{\Omega}^{(k)}$ by applying the strong dyadic Christ–Goldberg maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$). That is, we define
$\widetilde{N}_{\Omega}^{(k)}(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|V(x)^{-1/p}V^{1/p}\mathbf{1}_{\Omega}e_{k}|\rangle_{R}\mathbf{1}_{R}(x).$
If we consider $\widetilde{N}_{\Omega}\coloneq\sup_{k=1,\ldots,d}\widetilde{N}_{\Omega}^{(k)},$ this satisfies
$\widetilde{N}_{\Omega}(x)\gtrsim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{-1/p}N_{\Omega}(x).$
Indeed, just observe that
$N_{\Omega}(x)$ $=\sup_{R\in\bm{\mathcal{D}}(\Omega)}|V(x)^{-1/p}\mathcal{V}_{R}|\mathbf{1}_{R}(x)=\sup_{R}|\mathcal{V}_{R}V(x)^{-1/p}|\mathbf{1}_{R}(x)$
$\sim_{d}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V^{1/p}V(x)^{-1/p}|^{p}\rangle^{1/p}\mathbf{1}_{R}(x)$
$\lesssim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V^{1/p}V(x)^{-1/p}|\rangle_{R}\mathbf{1}_{R}(x)$
$=[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V(x)^{-1/p}V^{1/p}|\rangle_{R}\mathbf{1}_{R}(x),$
because the scalar weight $|V^{1/p}M|^{p}$ is uniformly in the biparameter $\bm{\mathcal{D}}$-dyadic Muckenhoupt $A_{p}$ class for every positive definite matrix $M$ (see for example [DKPS2024, Lemma 3.4]). Next, using the comparability between the matrix norm and the supremum of norms of matrix columns, we get
$N_{\Omega}(x)$ $\lesssim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\sup_{k=1,\ldots,d}\langle|V(x)^{-1/p}V^{1/p}e_{k}|\rangle_{R}\mathbf{1}_{R}(x)$
$\leq[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\widetilde{N}_{\Omega}(x).$
Finally, using the boundedness (4.2) of the strong dyadic Christ–Goldberg maximal operator, we obtain
$\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}.$
Thus, for all $A\in\mathcal{H}$ we have
$\|\hat{A}\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}$ $=\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\mathcal{V}_{R}A_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$=\int_{\Omega}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\mathcal{V}_{R}A_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$\leq\int_{\Omega}N_{\Omega}(x)\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|A_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}\,\mathrm{dm}(x)$
$\lesssim_{n,m,p,d}\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))}$
$\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))},$
concluding the proof of (4.3).
Uniqueness of $B$ follows immediately by testing $\ell$ on sequences in $\mathcal{H}.$
Remark 10.
It is clear that the proofs of Propositions 8, 9 work also in the one-parameter setting.
4.2 Two-matrix weighted bounds for paraproducts
Let $B:\mathbb{R}^{n+m}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|\bigg{)}^{1/2},$
where $\{B_{R}^{\varepsilon}\}$ is the sequence of biparameter Haar coefficients of $B$. Following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “pure” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
$\Pi_{\bm{\mathcal{D}},B}^{(11)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}h^{\varepsilon}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R},\quad\Pi_{\bm{\mathcal{D}},B}^{(00)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{\mathbf{1}_{R}}{|R|}B^{\varepsilon}_{R}f^{\varepsilon}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{1\oplus\varepsilon\oplus\delta}B_{R}^{\varepsilon}f_{R}^{\delta}.$
Here we denote
$1\oplus 1=0\oplus 0=0,\quad 1\oplus 0=0\oplus 1=1$
and extend these operations in the obvious entrywise fashion to $\mathcal{E}^{1},$ $\mathcal{E}^{2}$ and $\mathcal{E}.$ Moreover, abusing notation we denote $(1,\ldots,1)$ (where the number of entries is always clear from the context) by just $1$ Notice that $1\oplus\varepsilon_{i}\oplus\delta_{i}\neq 1,$ for all $\varepsilon,\delta\in\mathcal{E}$ with $\varepsilon_{i}\neq\delta_{i},$ $i=1,2.$
Clearly $(\Pi_{\bm{\mathcal{D}},B}^{(00)})^{\ast}=\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense. Observe also that a change of summation variables yields
$\Gamma_{\bm{\mathcal{D}},B}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}h_{R}^{\varepsilon}B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\delta},$
therefore $(\Gamma_{\bm{\mathcal{D}},B})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense.
Proposition 11.
Let $d,p,U,V$ and $B$ be as above.
  1. (a)
    There holds
    $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\sim\|\Pi_{\bm{\mathcal{D}},B}^{(00)}f\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
  2. (b)
    There holds
    $\|\Gamma_{\bm{\mathcal{D}},B}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
    where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
Proof.
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,p,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt part of the proof of [Isralowitz_2017, Theorem 2.2]. Note that by the John–Nirenberg inequalities for (unweighted) scalar dyadic product BMO we have
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}$
$\sim_{p,d,n,m}\sup_{\Omega}\frac{1}{|\Omega|^{1/p}}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\coloneq\mathcal{C},$
where the supremum is taken over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{2}$ of nonzero finite measure. Therefore, it suffices to prove that
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\gtrsim\mathcal{C}$
and
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
Let us first see that $\mathcal{C}\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$ Let $\Omega$ be any Lebesgue-measurable subset of $\mathbb{R}^{2}$ of nonzero finite measure. Let also $e\in\mathbb{C}^{d}\setminus\{0\}$ be arbitrary. We test $\Pi_{\bm{\mathcal{D}},B}^{(11)}$ on the function $f\coloneq\mathbf{1}_{\Omega}U^{-1/p}e.$ Using Lemma 5.3 of [DKPS2024] we obtain
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\gtrsim\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\geq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}.$
Note also that $\|f\|_{\mathrm{L}^{p}(U)}=|\Omega|^{1/p}\cdot|e|.$ Therefore, we see that
$\frac{1}{|\Omega|^{1/p}}\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}e|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}|e|,$
for all $e\in\mathbb{C}^{d}.$ In view of (2.3) we deduce
$\frac{1}{|\Omega|^{1/p}}\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle U^{-1/p}\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\lesssim\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}.$
By part (2) of Lemma 3 we have
$|P\langle U^{-1/p}\rangle_{R}|=|\langle U^{-1/p}\rangle_{R}P^{\ast}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P|\sim_{p,d}|\mathcal{U}^{-1}_{R}P|=|P\mathcal{U}^{-1}_{R}|,\quad\forall P\in\mathrm{M}_{d}(\mathbb{C}),$
yielding the desired result.
We now prove that $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ Using Lemma 5.3 of [DKPS2024] we have
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\sim\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\leq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}|\mathcal{U}_{R}\langle f\rangle_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\leq\bigg{(}\int_{\mathbb{R}^{n+m}}\bigg{(}\sum_{R\in\bm{\mathcal{D}}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}(\langle\widetilde{M}_{\bm{\mathcal{D}},U}f\rangle_{R})^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{p/2}\,\mathrm{dm}(x)\bigg{)}^{1/p}$
$\sim_{p,n,m}\|\Pi_{\bm{\mathcal{D}},b}^{(11)}(\widetilde{M}_{\bm{\mathcal{D}},U}f)\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})},$
where $b=(b_{R}^{\varepsilon})_{\begin{subarray}{c}R\in\bm{\mathcal{D}}(\Omega)\\ \varepsilon\in\mathcal{E}\end{subarray}}$ is the sequence given by
$b_{R}^{\varepsilon}\coloneq|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|,\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E},$
and
$\widetilde{M}_{\bm{\mathcal{D}},U}f\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|\mathcal{U}_{R}f|\rangle_{R}\mathbf{1}_{R}.$
By the well-known unweighted bounds for paraproducts in the scalar setting (see e. g. [Blasco_Pott_2005]) we have
$\|\Pi_{\bm{\mathcal{D}},b}^{(11)}\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{p,n,m}\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}.$
By definition, $\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}=\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ From [DKPS2024, Proposition 4.1] we also have
$\|\widetilde{M}_{\bm{\mathcal{D}},U}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim 1.$
It follows that
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\widetilde{M}_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|f\|_{\mathrm{L}^{p}(U)}.$
Finally, by duality we obtain
$\|\Pi_{\bm{\mathcal{D}},B}^{(00)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}$ $=\|\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}\|_{L^{p^{\prime}}(V^{\prime})\rightarrow L^{p^{\prime}}(U^{\prime})}\sim\|B^{\ast}\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(V^{\prime},U^{\prime},p^{\prime})}$
$\sim_{p,d}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We have
$(\Gamma_{\bm{\mathcal{D}},B}f,g)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\langle B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\varepsilon},g_{R}^{\delta}\rangle.$
Observe that if $A\in\mathrm{M}_{d}(\mathbb{C})$ and $x,y\in\mathbb{C}^{d},$ then
$\langle Ax,y\rangle$ $=x^{T}A^{T}\bar{y}=\mathrm{tr}(x^{T}A^{T}\bar{y})=\overline{\mathrm{tr}(A^{\ast}y\bar{x}^{T})}.$
It follows that
$|(\Gamma_{\bm{\mathcal{D}},B}f,g)|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}\text{tr}((B_{R}^{\varepsilon})^{\ast}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T})\bigg{|}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R|}}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\bigg{(}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\bigg{)}\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}|^{2}\cdot|\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$=2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
$\leq 2^{n+m}S_{\bm{\mathcal{D}},V^{\prime}}g(x)\widetilde{S}_{\bm{\mathcal{D}},U}f(x),$
for all $x\in\mathbb{R}^{n+m},$ where
$S_{\bm{\mathcal{D}},V^{\prime}}g\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2},\quad\widetilde{S}_{\bm{\mathcal{D}},U}f\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \delta\in\mathcal{E}\end{subarray}}|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}}{|R|}\bigg{)}^{1/2}.$
Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq 2^{n+m}\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemma 5.2 and Corollary 5.4] that
$\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},\qquad\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})},$
yielding the desired bound. ∎
Next, following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “mixed” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
$\Pi_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(h_{R_{1}}^{\varepsilon_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}},\quad\Pi_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\varepsilon_{2}}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}h_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{1\oplus\varepsilon_{2}\oplus\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\\ \\ delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}h_{R}^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{1\oplus\varepsilon_{1}\oplus\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{\delta_{1},\varepsilon_{2}}_{R}.$
It is clear that $(\Pi_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(10)}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense. Observe also that a change of summation variables yields
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}_{R},$
therefore $(\Gamma_{\bm{\mathcal{D}},B}^{(10)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(10),\ast}$ as well as $(\Gamma_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(01),\ast}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense.
Two-matrix weighted bounds for the above mixed paraproducts can be easily deduced from two-matrix weighted bounds for the mixed type operators considered in [DKPS2024, Section 8]. Although at the time of [DKPS2024] only a rather incomplete treatment of the latter bounds was possible, they all follow now readily from Theorem 1 and Theorem 7. We give only one example below, the proof for the other operators being similar. Note that the parts of the proof that do not rely on the matrix weighted extension of the Fefferman–Stein vector valued inequalities were already carried out in [DKPS2024, Section 8]. Nevertheless, for the reader’s convenience we include full details.
Lemma 12.
Let $1<p<\infty$ and let $W$ be a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ For (suitable) functions $f:\mathbb{R}^{n+m}\to\mathbb{C},$ let
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
for all $x=(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Then, we have
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\beta}\|f\|_{\mathrm{L}^{p}(W)},$
where
$\beta=\begin{cases}1+\frac{2}{p}+\frac{1}{p-1},\text{ if }p\leq 2\\ \\ \frac{1}{2}+\frac{1}{p}+\frac{2}{p-1},\text{ if }p>2\end{cases}.$
To prove Lemma 12, we need a technical observation already present implicitly in the proof of [DKPS2024, Lemma 5.3].
Lemma 13.
Let $1<p<\infty$ and let $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ be a family of nonnegative measurable functions on $\mathbb{R}^{m}.$ Then, it holds
$\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\lesssim_{m,p}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}.$
Proof.
We use duality. Note that through an application of the monotone convergence theorem we may assume without loss of generality that only finitely many of the functions $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ are not identically equal to zero. Then, by the dyadic (unweighted) Littlewood–Payley estimates we have
$\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\sim_{m,p}\|F\|_{\mathrm{L}^{p}(\mathbb{R}^{m})},$
where
$F\coloneq\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle_{P}h_{P}^{\varepsilon}.$
Let $g\in L^{p^{\prime}}(\mathbb{R}^{m})$ be arbitrary. Then, we have
$\int_{\mathbb{R}^{m}}|F(x)g(x)|\,\mathrm{dm}(x)\leq\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\langle f_{P,\varepsilon}\rangle\cdot|g_{P}^{\varepsilon}|$
$=\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}\int_{\mathbb{R}^{m}}f_{P,\varepsilon}(x)h_{P}^{\varepsilon}(x)g_{P}^{\varepsilon}h_{P}^{\varepsilon}(x)\,\mathrm{dm}(x)$
$\leq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|g_{P}^{\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{m})}$
$\sim_{m,p}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\|g\|_{L^{p^{\prime}}(\mathbb{R}^{m})}.$
An appeal to the Riesz representation theorem concludes the proof. ∎
We now prove Lemma 12.
Proof of Lemma 12.
First of all, we have
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}^{p}=\int_{\mathbb{R}^{n}}A(x_{1})\,\mathrm{dm}(x_{1}),$
where
$A(x_{1})\coloneq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{2}}|\mathcal{W}_{R_{1}\times R_{2}}(f_{R_{2}}^{\varepsilon_{2},2})_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}),$
for all $x_{1}\in\mathbb{R}^{m}.$
Fix $x_{1}\in\mathbb{R}^{m}.$ For a.e. $x_{2}\in\mathbb{R}^{n},$ we denote by $\mathcal{W}_{x_{2},R_{1}}$ the reducing operator of the weight $W_{x_{2}}(y)\coloneq W(y,x_{2}),$ $y\in\mathbb{R}^{n}$ over any $R_{1}\in\mathcal{D}^{1}$ with respect to the exponent $p.$ For fixed $R_{1}\in\mathcal{D}^{1},$ we define $W_{R_{1}}(x_{2})\coloneq\mathcal{W}_{x_{2},R_{1}}^{p},$ for a.e. $x_{2}\in\mathbb{R},$ and denote by $\mathcal{W}_{R_{1},R_{2}}$ the reducing operator of $W_{R_{1}}$ over any $R_{2}\in\mathcal{D}^{2}$ with respect to the exponent $p.$ Applying now first (2.3), then part (1) of Lemma 3, and finally (2.4), we obtain
$\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R_{1}\times R_{2}}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\lesssim_{p,d}\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R_{1},R_{2}}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})$
$\lesssim_{p,d}\sup_{R_{1}\in\mathcal{D}^{1}}[W_{R_{1}}]_{A_{p},\mathcal{D}^{2}}^{\frac{1}{p}}\left\langle|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\right\rangle_{R_{2}}\mathbf{1}_{R_{1}}(x_{1})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\frac{1}{p}}\bigg{\langle}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{\rangle}_{R_{2}},$
for all $\varepsilon_{2}\in\mathcal{E}^{2}$ and $R_{2}\in\mathcal{D}^{2}.$ Observe that in the last $\lesssim_{p,d},$ we used (2.4), as well as the fact that the supremum of the integrals is dominated by the integral of the supremum. So
$A(x_{1})\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}B(x_{1}),$
where
$B(x_{1})\coloneq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{\langle}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{\rangle}_{R_{2}}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}).$
Since $x_{1}$ is fixed, by Lemma 13 we obtain
$B(x_{1})$ $\leq\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})$
$=\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{2}).$
Thus, by Fubini–Tonelli we have
$\int_{\mathbb{R}^{n}}B(x_{1})\,\mathrm{dm}(x_{1})$
$\leq\int_{\mathbb{R}^{m}}\bigg{(}\int_{\mathbb{R}^{n}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})\bigg{)}\,\mathrm{dm}(x_{2}).$
For a.e. $x_{2}\in\mathbb{R}^{m},$ using Theorem 7 in the first step and (2.6.3) in the second step, we obtain
$\int_{\mathbb{R}^{n}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}\bigg{(}\widetilde{M}_{W_{x_{2}},\mathcal{D}^{1}}(f_{R_{2}}^{\varepsilon_{2},2})(x_{1})\bigg{)}^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$\lesssim_{m,p,d}[W_{x_{2}}]_{A_{p},\mathcal{D}^{1}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}|W_{x_{2}}^{1/p}(x_{1})f_{R_{2}}^{\varepsilon_{2},2}(x_{1})|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}|W(x_{1},x_{2})^{1/p}f_{R_{2}}^{\varepsilon_{2},2}(x_{1})|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{p/2}\,\mathrm{dm}(x_{1})$
$=[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{1}),$
where we denote $W_{x_{1}}\coloneq W(x_{1},x_{2}),$ $x_{1}\in\mathbb{R}^{n}.$ Thus, applying Fubini–Tonelli again we get
$\int_{\mathbb{R}^{n}}B(x_{1})\,\mathrm{dm}(x_{1})$
$\lesssim_{m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{1+\max\{p,p^{\prime}\}}\int_{\mathbb{R}^{n}}\bigg{(}\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})\bigg{)}\,\mathrm{dm}(x_{1}).$
For a.e. $x_{1}\in\mathbb{R}^{n},$ applying first the matrix weighted bounds for the one-parameter dyadic square function from [Isralowitz_2020] and then (2.6.3), we get
$\int_{\mathbb{R}^{m}}\bigg{(}S_{\mathcal{D}^{2},W_{x_{1}}}(f(x_{1},\,\cdot\,))(x_{2})\bigg{)}^{p/2}\,\mathrm{dm}(x_{2})$
$\lesssim_{n,p,d}[W_{x_{1}}]_{A_{p},\mathcal{D}_{2}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W_{x_{1}}(x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2})$
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\}}\int_{\mathbb{R}^{m}}|W(x_{1},x_{2})^{1/p}f(x_{1},x_{2})|^{p}\,\mathrm{dm}(x_{2}).$
Putting the above estimates together, we finally deduce
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},W}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}^{p}\lesssim_{n,m,p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\alpha}\int_{\mathbb{R}^{n+m}}|W(x)^{1/p}f(x)|^{p}\,\mathrm{dm}(x),$
where
$\alpha=1+1+\max\{p,p^{\prime}\}+\max\left\{p^{\prime},\frac{p}{2}+\frac{1}{p-1}\right\},$
concluding the proof. ∎
We now prove two-matrix weighted bounds for the mixed paraproducts.
Proposition 14.
Let $d,p,U,V$ and $B$ be as above. If $\Pi$ is any of the above defined mixed biparameter paraproducts, then there holds
$\|\Pi\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}\lesssim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$
where all implied constants depend only on $n,m,d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
Proof.
Throughout the proof ,$\lesssim$,$\gtrsim$,$\sim$ , ≳ , ∼ mean that all implied inequality constants depend only on $n,m,d,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We treat as an example $\Gamma_{\bm{\mathcal{D}},B}^{(10)},$ the proof for the other mixed paraproducts being similar or following by duality. We have
$\left|(\Gamma_{\bm{\mathcal{D}},B}^{(10)}f,g)\right|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\langle B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},g_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}\rangle\bigg{|}$
$=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})\bigg{|}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)},$
where
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
It suffices now to prove that
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$ (4.6)
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\left(\frac{1}{\sqrt{|R_{2}|}}\sum_{\delta_{2}\in\mathcal{E}^{2}\setminus\{\varepsilon_{2}\}}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\right)\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R}\overline{\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R_{1}|\cdot|R_{2}|^{2}}$
$\leq 2^{m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}|V(x)^{-1/p}g^{\varepsilon}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}}|^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}$
$\leq 2^{m}S_{\bm{\mathcal{D}},V^{\prime}}g(x)[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x),$
where
$S_{\bm{\mathcal{D}},V^{\prime}}g(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\bigg{)}^{1/2}$
and
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{U}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
for all $x=(x_{1},x_{2})\in\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Therefore
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemmas 5.2] that
$\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
Moreover, Lemma 12 yields
$\|[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},$
proving (4.6). ∎
4.3 Two matrix weighted upper bounds for bicommutators
Two matrix weighted upper bounds for biparameter paraproducts lead naturally to two matrix weighed upper bounds for bicommutators. Here we overview very briefly the simplest case: that of bicommutators with Haar multipliers on $\mathbb{R}.$ The argument is a straightforward adaptation of the one in the (unweighted) $\mathrm{L}^{2}$ scalar valued case from [Blasco_Pott_2005].
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be a biparameter dyadic grid on $\mathbb{R}^{2},$ and let $\sigma_{1}=\{\sigma_{1}(I)\}_{I\in\mathcal{D}^{1}}$ and $\sigma_{2}=\{\sigma_{2}(J)\}_{J\in\mathcal{D}^{2}}$ be (finitely supported) sequences in $\{-1,0,1\}.$ For any function $f\in\mathrm{L}^{2}(\mathbb{R};\mathbb{C}^{d})$ we define
$T_{\sigma_{1}}f\coloneq\sum_{I\in\mathcal{D}^{1}}\sigma_{1}(I)h_{I}f_{I},\quad T_{\sigma_{2}}f\coloneq\sum_{J\in\mathcal{D}^{2}}\sigma_{2}(J)h_{J}f_{J}.$
We can then consider the operators $T^{1}_{\sigma_{1}}$ and $T^{2}_{\sigma_{2}}$ acting on functions $f\in\mathrm{L}^{2}(\mathbb{R}^{2};\mathbb{C}^{d})$ by
$T^{1}_{\sigma_{1}}f(x_{1},x_{2})\coloneq T_{\sigma_{1}}(f(\,\cdot\,,x_{2}))(x_{1}),\quad T^{2}_{\sigma_{2}}f(x_{1},x_{2})\coloneq T_{\sigma_{2}}(f(x_{1},\,\cdot\,))(x_{2}),$
for a.e. $(x_{1},x_{2})\in\mathbb{R}^{2}.$
Let $B:\mathbb{R}^{2}\to\mathrm{M}_{d}(\mathbb{C})$ be a locally integrable function. For scalar valued locally integrable functions $b$ on $\mathbb{R}^{2}$ it is shown in [Blasco_Pott_2005] that
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},b]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{b}]],$ (4.7)
where $\Lambda_{b}$ is the so-called symmetrized paraproduct given by
$\Lambda_{b}f\coloneq\Pi^{(11)}_{b}f+\Pi^{(10)}_{b}f+\Pi^{(01)}_{b}f+\Pi^{(00)}_{b}f.$
Applying (4.7) entrywise we deduce
$[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},B]]=[T^{1}_{\sigma_{1}},[T^{2}_{\sigma_{2}},\Lambda_{B}]],$
where the symmetrized paraproduct $\Lambda_{B}$ is given by
$\Lambda_{B}f\coloneq\Pi^{(11)}_{B}f+\Pi^{(10)}_{B}f+\Pi^{(01)}_{B}f+\Pi^{(00)}_{B}f.$
Let now $1<p<\infty$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}\times\mathbb{R}.$ In the following estimates, all implied constants depend only on $d,p,[U]_{A_{p},\bm{\mathcal{D}}}$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Using (2.6.3) and the well-known two matrix weighted bounds in the one parameter setting we deduce
$\|T^{j}_{\sigma_{j}}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1,\quad j=1,2.$
Observe also that Proposition 11 and Proposition 14 immediately yield
$\|\Lambda_{B}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1.$
Thus, we obtain
$\|[T_{1},[T_{2},B]]\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}=\|[T_{1},[T_{2},\Lambda_{B}]]\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}$
$\leq\|T_{1}T_{2}\Lambda_{B}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}+\|T_{1}\Lambda_{B}T_{2}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}$
$+\|T_{2}\Lambda_{B}T_{1}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}+\|\Lambda_{B}T_{1}T_{2}\|_{\mathrm{L}^{p}(U)\to\mathrm{L}^{p}(V)}\lesssim 1.$
Appendix A Appendix
Through the article we have used the results of Bownik and Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation] in the complex setting. The results in that paper are stated in the real setting and it is not immediate that they work in the complex setting. Some modifications are necessary and some steps require proper justification. For instance, given a family of norms measurably parametrized (over $\mathbb{C}^{d}$), it is not obvious that one can assign them a reducing matrix and a complex John ellipsoid in a measurable way. This is shown in detail in [DKPS2024, Appendix A]. We devote this appendix to briefly discussing the necessary modifications and where to find the details to get the results of [Cruz_Uribe_Bownik_Extrapolation] for complex valued matrix weights.
A.1 Convex sets and seminorms in the complex setting
We begin with the required changes in [Cruz_Uribe_Bownik_Extrapolation, Section 2]. In both the present work and that of Bownik and Cruz-Uribe, one considers symmetric sets. The difference is that we substitute real symmetric sets $E\subseteq\mathbb{R}^{d},$ that is sets $E$ such that $-u\in E$ for every $u\in E,$ by complex symmetric sets $E\subseteq\mathbb{C}^{d},$ i.e. sets $E$ for which $\lambda u\in E$ for every $u\in E$ and for any $\lambda\in\mathbb{C}$ with $|\lambda|=1.$ Let us note that the “symmetric convex sets” as we have defined them are precisely the balanced convex sets, meaning that $\lambda u\in E$, for all $u\in E$ and $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1.$ Besides this, the rest of the definitions concerning convex bodies, including that of the Minkowski addition of sets, is the same. Of course, the definitions for seminorms are also the same, only that we consider them to be defined on the vector space $\mathbb{C}^{d}$ and to be complex homogeneous functions, which means that if $\rho$ is a seminorm on $\mathbb{C}^{d},$ then $\rho(\lambda v)=|\lambda|\rho(v)$ for any $v\in\mathbb{C}^{d}$ and $\lambda\in\mathbb{C}.$ Also, the fact that seminorms are in a one-to-one correspondence with absorbing convex (complex) symmetric bodies (the analogous result to [Cruz_Uribe_Extrapolation, Theorem 2.4]) also holds with the same proof (see [Rudin_Functional_Analysis_1991, Theorems 1.34 and 1.35]).
Dual seminorms, their properties and their relation to the polar of convex (complex) symmetric sets follow the same exposition in the vector space $\mathbb{C}^{d}$ as in [Cruz_Uribe_Bownik_Extrapolation]. Here one only needs to keep in mind that we substitute the real Euclidean product by the complex Hermitian product of $\mathbb{C}^{d}.$ For this reason, in this context the support function $h_{K}$ of a set $K\in\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ has to be defined as
$h_{K}(v)\coloneq\sup_{w\in K}|\langle v,w\rangle|,$
(cf. [Cruz_Uribe_Bownik_Extrapolation, Definition 2.12]). With this definition, the proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 2.11], which covers the properties of seminorms and their relation to convex bodies, needs to be adapted. For instance, to show that
$p_{(K_{1}+K_{2})^{\circ}}=p_{K_{1}^{\circ}}+p_{K_{2}^{\circ}},$
one first shows that
$p_{(K_{1}+K_{2})^{\circ}}\leq p_{K_{1}^{\circ}}+p_{K_{2}^{\circ}}$
by the triangle inequality. Then, the reverse inequality is proved by taking $\varepsilon>0$ arbitrary and $w_{1}\in K_{1}$ and $w_{2}\in K_{2}$ with $\langle v,w_{j}\rangle\geq h_{K_{j}}(v)-\varepsilon/2$ for $j=1,2$ (we can omit the moduli here because of complex symmetry). This argument has to be repeated through the rest of the statements of [Cruz_Uribe_Bownik_Extrapolation, Theorem 2.11]. The properties of weighted geometric means of norms follow the same explanation for norms defined on $\mathbb{C}^{d}.$
For the rest of the section, the facts about positive definite matrices follow by the same arguments, since the facts that the authors of [Cruz_Uribe_Bownik_Extrapolation] use hold in the complex setting (see [Bhatia_Positive_definite_matrices_2007, Chapter 6]). Here, one only needs to take into account that we consider hermitian matrices instead of symmetric ones and that we use unitary matrices instead of orthogonal ones.
A.2 Convex-set valued functions in the complex setting
In this subsection we explain the necessary adaptations that have to be performed in Section 3 of [Cruz_Uribe_Bownik_Extrapolation] in the complex setting. In some places $\mathbb{R}^{d}$ needs to be replaced by $\mathbb{R}^{2d}$, which as a topological space, as a metric space, as a measure space and as a real vector space is the same as $\mathbb{C}^{d}$. In some other places, $\mathbb{R}^{d}$ needs to be replaced directly by $\mathbb{C}^{d}$ and “complex” versions of the ingredients of statements or proofs are necessary. We lay out the details below.
First of all, measurable maps $F:\Omega\to\mathcal{K}(\mathbb{C}^{d})$ from a positive, $\sigma$-finite, complete measure space $\Omega$ into the set $\mathcal{K}(\mathbb{C}^{d})$ of closed subsets of $\mathbb{C}^{d}$ are the same as measurable maps $F:\Omega\to\mathcal{K}(\mathbb{R}^{2d})$. The various characterizations of measurability for maps $F:\Omega\to\mathcal{K}(\mathbb{R}^{d})$ given in [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.2] with $\mathbb{R}^{d}$ as the underlying space continue to hold with $\mathbb{R}^{d}$ replaced by $\mathbb{R}^{2d}$, and thus $\mathbb{C}^{d}$ as the underlying space. Let us observe in particular that a measurable selection function $f:\Omega\to\mathbb{C}^{d}$ of a map $F:\Omega\to\mathcal{K}(\mathbb{C}^{d})$ is the same as a measurable selection function $f:\Omega\to\mathbb{R}^{2d}$ of the map $F:\Omega\to\mathcal{K}(\mathbb{R}^{2d})$.
Similarly, the measurability of closed convex hulls of countable unions and the measurability of countable intersections of convex body valued functions [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.3] hold in $\mathbb{R}^{2d}$ and therefore also in $\mathbb{C}^{d}.$ A similar remark applies to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5], which states the equivalence between the definition of closed set valued functions and the measurability as a mapping with respect to the Hausdorff topology.
Referring to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.4] proving the measurability of the polar of a measurable map, one replaces $\mathbb{R}^{d}$ directly by $\mathbb{C}^{d}$. This entails that complex symmetric convex sets are considered, and that the real Euclidean product on $\mathbb{R}^{d}$ needs to be replaced by the complex Hermitian product on $\mathbb{C}^{d}$. A similar change applies to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6]. This result applies to measurable set valued functions $F$ taking values on the set of linear subspaces of $\mathbb{C}^{d},$ and it states that if $P(x)$ is the orthogonal projection from $\mathbb{C}^{d}$ to $F(x),$ the mapping $P\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ is also measurable. The necessary change is to replace $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$, so that orthogonal projections and the Gram–Schmidt orthonormalization process are considered with respect to the Hermitian product on $\mathbb{C}^{d}$. Real linear subspaces are replaced by complex linear subspaces. At the last step of the proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6], a countable dense subset of $\mathbb{C}^{d}$ is needed, one can take $\mathbb{Q}(i)^{d}$ for example.
Next, one needs the existence and properties of John ellipsoids for bounded complex symmetric convex subsets of $\mathbb{C}^{d}$. These are thoroughly established in [DKPS2024, Appendix A] and the distinction between complex and real ellipsoids is made precise. In fact, it is implicit in [DKPS2024, Subsection A.4.7] that a real ellipsoid in $\mathbb{R}^{2d}$ is a complex ellipsoid in $\mathbb{C}^{d}$ if and only if it is invariant under the real orthogonal map $\widetilde{L}_{z}$ for all $z\in\mathbb{C}$ with $|z|=1$. For the reader’s convenience we lay out the details below.
We identify $\mathbb{R}^{2d}$ with $\mathbb{C}^{d}$ in the natural way, namely by the map $R:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ given by
$R(x_{1}+ix_{2},\ldots,x_{2d-1}+ix_{2d}):=(x_{1},x_{2},\ldots,x_{2d-1},x_{2d}).$
We denote by $\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ the closed unit ball in $\mathbb{R}^{2d}$ and by $\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ the closed unit ball in $\mathbb{C}^{d}$, so $\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=R(\mathbf{\overline{B}}_{\mathbb{C}^{d}})$.
A real ellipsoid (more precisely, a nondegenerate centrally symmetric ellipsoid) in $\mathbb{R}^{2d}$ is by definition a subset $E$ of $\mathbb{R}^{2d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ for some invertible real linear map $A:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. In this case, the polar decomposition yields $E=(AA^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$. It follows that for each ellipsoid $E$ in $\mathbb{R}^{2d}$ there is a unique real positive definite linear map $A:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ with $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$, which we denote by $A:=M_{\mathbb{R},2d}(E)$.
A complex ellipsoid (more precisely, a nondegenerate centrally symmetric ellipsoid) in $\mathbb{C}^{d}$ is by definition a subset $E$ of $\mathbb{C}^{d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ for some invertible complex linear map $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$. In this case, the polar decomposition yields $E=(AA^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{C}^{d}}$. It follows that for each ellipsoid $E$ in $\mathbb{C}^{d}$ there is a unique complex positive definite linear map $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$ with $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$, which we denote by $A:=M_{\mathbb{C},d}(E)$.
If $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is any complex linear map, then under our identification of $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$, $A$ corresponds to the real linear map $RAR^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. Concretely, $RAR^{-1}$ is obtained by $A$ through replacing each complex entry $a_{ij}$ of $A$ with
$H(a_{ij})=\begin{bmatrix}\mathrm{Re}(a_{ij})&-\mathrm{Im}(a_{ij})\\ \mathrm{Im}(a_{ij})&\mathrm{Re}(a_{ij})\end{bmatrix}.$
Observe that the map $H:\mathbb{C}\to\mathbb{R}^{2\times 2}$ is a ring monomorphism. We write $H(A):=RAR^{-1}$. It is then easy to see that if $A$ is complex hermitian, respectively unitary, respectively positive definite, then $H(A)$ is real symmetric, respectively orthogonal, respectively positive definite. In particular, if for each $z\in\mathbb{C}$, $L_{z}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ denotes multiplication with $z$, then we can consider the corresponding linear map $\widetilde{L}_{z}:=RL_{z}R^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. So for each $z\in\mathbb{C}$ and for each nonempty subset $E$ of $\mathbb{C}^{d}$ we have $zE=E$ if and only if $\widetilde{L}_{z}(R(E))=R(E)$. Moreover, if $z\in\mathbb{C}$ with $|z|=1$, then $\widetilde{L}_{z}$ is easily seen to be a real orthogonal map.
The precise relation between real and complex ellipsoids is explained in the following lemma.
Lemma 15.
Let $\mathcal{C}_{2d}(\mathbb{R})$ be the set of all real ellipsoids in $\mathbb{R}^{2d}$ and let $\mathcal{C}_{d}(\mathbb{C})$ be the set of all complex ellipsoids in $\mathbb{C}^{d}$. Then we have
$\mathcal{C}_{d}(\mathbb{C})=\{R^{-1}(E):\leavevmode\nobreak\ E\in\mathcal{C}_{2d}(\mathbb{R}^{2d})\text{ and }\widetilde{L}_{z}(E)=E\}.$ (A.1)
In particular, the complex ellipsoids in $\mathbb{C}^{d}$ form a closed subset of the real ellipsoids in $\mathbb{R}^{2d}$.
Proof.
First of all, let $E$ be a real ellipsoid in $\mathbb{R}^{2d}$ with $\widetilde{L}_{z}(E)=E$. Then, for each $z\in\mathbb{C}$ with $|z|=1$ we have
$E$ $=\widetilde{L}_{z}E=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=(\widetilde{L}_{z}M_{\mathbb{R},2d}(E)(\widetilde{L}_{z}M_{\mathbb{R},2d}(E))^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}},$
where we used the fact that $\widetilde{L}_{z}$ is an orthogonal linear map. Since $\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1}$ is positive definite, by the uniqueness of $M_{\mathbb{R},2d}(E)$ it follows that
$M_{\mathbb{R},2d}(E)=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1},$
thus also
$M_{\mathbb{R},2d}(E)=\widetilde{L}_{z}^{-1}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}.$
Now we set $\widetilde{E}:=R^{-1}(E)$ and compute
$\widetilde{E}=R^{-1}M_{\mathbb{R},2d}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=R^{-1}M_{\mathbb{R},2d}(E)R\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
In order to show that $\widetilde{E}\in\mathcal{C}_{d}(\mathbb{C})$, it suffices to show that the invertible real linear map $R^{-1}M_{\mathbb{R},2d}(E)R:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is in fact complex linear. To this end, it suffices to prove that
$L_{z}R^{-1}M_{\mathbb{R},2d}(E)R=R^{-1}M_{\mathbb{R},2d}(E)RL_{z},$
for all $z\in\mathbb{C}$ in $|z|=1$. For such $z$ we compute
$L_{z}R^{-1}M_{\mathbb{R},2d}(E)RL_{z}^{-1}=L_{z}R^{-1}\widetilde{L}_{z}^{-1}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}RL_{z}^{-1}$
$=L_{z}R^{-1}RL_{z}^{-1}R^{-1}M_{\mathbb{R},2d}(E)RL_{z}R^{-1}RL_{z}^{-1}$
$=R^{-1}M_{\mathbb{R},2d}(E)R.$
This proves the inclusion $\supseteq$ in (A.1).
We now show the inclusion $\subseteq$ in (A.1). Let $\widetilde{E}$ be any complex ellipsoid in $\mathbb{R}^{2d}$. We set $E:=R(\widetilde{E})$. Then, we compute
$E=RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}.$
The map $RM_{\mathbb{C},d}(\widetilde{E})R^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ is invertible real linear, thus $E$ is a real ellipsoid in $\mathbb{R}^{2d}$. Moreover, for all $z\in\mathbb{C}$ with $|z|=1$ we have
$\widetilde{L}_{z}E$ $=\widetilde{L}_{z}RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RL_{z}R^{-1}RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=RL_{z}M_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C},d}(\widetilde{E})L_{z}R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C},d}(\widetilde{E})L_{z}\mathbf{\overline{B}}_{\mathbb{C}^{d}}$
$=RM_{\mathbb{C},d}(\widetilde{E})\mathbf{\overline{B}}_{\mathbb{C}^{d}}=E,$
where we used the fact that $L_{z}M_{\mathbb{C},d}(\widetilde{E})=M_{\mathbb{C},d}(\widetilde{E})L_{z}$ due to the complex linearilty of $M_{\mathbb{C},d}(\widetilde{E})$ as well as the fact that the closed unit ball of $\mathbb{C}^{d}$ remains invariant under multiplication with $z$.
In particular, the complex ellipsoids in $\mathbb{C}^{d}$ form a closed subset of the real ellipsoids in $\mathbb{R}^{2d}$, for orthogonal maps induce isometries with respect to the Hausdorff metric. ∎
The previous exposition is related to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.7]. The idea behind this result is the fact that for any measurable convex body valued function $F,$ the function $G(x)$ defined as the John ellipsoid of the convex body $F(x)$ is also measurable. For the reader’s convenience, we include the precise statement in the context of complex convex body valued functions.
Theorem 16 (Theorem 3.7 in [Cruz_Uribe_Bownik_Extrapolation]).
Suppose that $F\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ is measurable. Then there exists a measurable matrix-valued mapping $W\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ such that
  1. (i)
    the columns of $W(x)$ are mutually orthogonal,
  2. (ii)
    for every $x\in\Omega,$ it holds that
    $W(x)\mathbf{\overline{B}}_{\mathbb{C}^{d}}\subseteq F(x)\subseteq\sqrt{d}W(x)\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
The proof of this result is based on [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.8], [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9] and [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.10], in a way analogous to that of the real convex body context. Since this is one of the crucial points to get Theorem A, we include the statements of these lemmata and their proofs.
Lemma 17 (Lemma 3.8 in [Cruz_Uribe_Bownik_Extrapolation]).
Given a measurable convex body valued function $F\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that $F(x)$ has nonempty interior for every $x\in\Omega,$ there exists a measurable convex body valued function $G\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that $G(x)$ is an ellipsoid with nonempty interior and with
$G(x)\subseteq F(x)\subseteq\sqrt{d}G(x)$ (A.2)
for every $x\in\Omega.$
Proof.
We follow the same approach as in [Cruz_Uribe_Bownik_Extrapolation]. For every $x\in\Omega,$ define $G(x)$ as the unique John ellipsoid of $F(x),$ which has nonempty interior. This already satisfies (A.2). It is only left to prove the measurability of $G.$
To this end, consider a dense sequence $P_{1},P_{2},\ldots$ of invertible matrices in $\mathrm{M}_{{d}}({\mathbb{C}}).$ In particular, we have that for any ellipsoid $E$ with nonempty interior, we can express $E=P\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}$ with $P\in\mathrm{M}_{{d}}({\mathbb{C}})$ and for any $\varepsilon>0$ there exists $n\geq 1$ such that
$P_{n}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq E\subseteq(1+\varepsilon)P_{n}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
We construct by induction a sequence of measurable convex body valued functions $G_{n}\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that for every $n\geq 1,$ $G_{n}(x)$ is either an ellipsoid with nonempty interior or $G_{n}(x)=\{0\}$ for every $x\in\Omega.$ First, define
$G_{1}(x)=\begin{cases}P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}&\text{ if }P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x),\\ \{0\}&\text{ otherwise. }\end{cases}$
The values of $G_{1}$ are either an ellipsoid with nonempty interior or $\{0\}$ by construction, while it is easy to see that it is also measurable since the set
$\{x\in\Omega\colon P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\nsubseteq F(x)\}=\{x\in\Omega\colon F(x)\cap P_{1}(\mathbb{C}^{d}\setminus\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\neq\emptyset\}$
is measurable due to $F$ being measurable. Assume now that we have defined measurable functions $G_{1},\ldots,G_{n}$ and define
$G_{n+1}(x)=\begin{cases}P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}&\text{if }P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x)\text{ and }m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})>m_{d}(G_{n}(x)),\\ G_{n}(x)&\text{ otherwise.}\end{cases}$
Again, by construction $G_{n+1}(x)$ is either an ellipsoid with nonempty interior or just the trivial convex body $\{0\}.$ We need to see that $G_{n+1}$ is also measurable. To do so, remember that the volume functional $K\mapsto m_{d}(K)$ is a continuous mapping from $\mathcal{K}_{b}(\mathbb{C}^{d})$ to $[0,\infty),$ since this is the same as the volume functional $K\mapsto m_{2d}(K)$ from $\mathcal{K}_{b}(\mathbb{R}^{2d})$ to $[0,\infty).$ This and the measurability of $G_{n}$ implies that $m_{d}(G_{n}(x))\colon\Omega\rightarrow[0,\infty)$ is also a measurable mapping (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5]). Now, for any given open set $U\subseteq\mathbb{C}^{d},$ we have that
$\{x$ $\in\Omega\colon G_{n+1}(x)\cap U\neq\emptyset\}$
=($=(\{x\in\Omega\colon m_{d}(G_{n}(x))\geq m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\}$
$\cap\{x\in\Omega\colon G_{n}(x)\cap U\neq\emptyset\})$
($\cup(\{x\in\Omega\colon m_{d}(G_{n}(x))<m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\}$
$\cap\{x\in\Omega\colon P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x)\text{ and }P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\cap U\neq\emptyset\}).$
By the previous considerations, each of the sets appearing on the right-hand side is measurable, so $G_{n+1}$ is measurable as well.
The last step of the proof is to show that the sequence $G_{n}(x)$ converges to $G(x)$ in the Hausdorff metric. This will imply that $G$ is measurable with respect to the Hausdorff topology, which is equivalent to the definition that we have used of measurability of convex body valued functions (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5]). Assume this is not the case for a given $x\in\Omega.$ By our construction, the density of the sequence $P_{1},P_{2},\ldots$ and the maximality of the John ellipsoid $G(x),$ we have that $m_{d}(G_{n})\to m_{d}(G(x))$ as $n\to\infty.$ The Blaschke selection theorem [Schneider_Convex_Bodie_1993, Theorem 1.8.7] asserts that any bounded sequence of convex bodies has a subsequence that converges to a convex body, so we have a subsequence $G_{n_{k}}(x)$ converging to an ellipsoid $E^{\prime}$ with $m_{d}(E^{\prime})=m_{d}(G(x))$ but with $E^{\prime}\neq G(x),$ which contradicts the fact that the John ellipsoid is unique. Note that $E^{\prime}$ is a complex ellipsoid, since real ellipsoids are closed in the Hausdorff topology and complex ellipsoids are a closed subset of the real ellipsoids. Thus $G_{n}(x)\to G(x)$ and the last function is also measurable. ∎
Remark 18.
Another approach consists in noting that the map sending each convex body with nonempty interior to its John ellipsoid is continuous with respect to the Hausdorff distance. In the real case, [DKPS2024, Subsection A.4.3] appeals directly to [Mordhorst_2017] for this result. An extension to the complex setting is performed in detail in [DKPS2024, Subsection A.4.7]. Combining this with the complex version of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5], we immediately deduce the complex version of [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.8], since the composition of measurable maps remains measurable.
The statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9] for complex convex bodies correspond to Lemma 4 in the present work. The proof included in Section 2.7 already covers the case of complex convex bodies.
Finally, [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.10] relates measurable complex ellipsoid valued mappings $G$ to measurable matrix valued mappings with mutually orthogonal columns.
Lemma 19 (Lemma 3.10 in [Cruz_Uribe_Bownik_Extrapolation]).
Consider a measurable mapping $G\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that $G(x)$ is a complex ellipsoid for every $x\in\Omega$ (possibly with empty interior). Then there exists a measurable mapping $W\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ such that
  1. (i)
    the columns of the matrix $W(x)$ are mutually orthogonal,
  2. (ii)
    it holds that $G(x)=W(x)\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}$ for every $x\in\Omega.$
Proof.
The proof consists in constructing the columns $v_{1},\ldots,v_{d}$ of the matrix $W(x)$ as measurable mappings $\Omega\rightarrow\mathbb{C}^{d}$ and being mutually orthogonal (with some of them possibly null at a given $x\in\Omega$). Given our measurable mapping $G,$ take the measurable mapping $v_{1}\colon\Omega\rightarrow\mathbb{C}^{d}$ such that $|G(x)|=|v_{1}(x)|$ at every $x\in\Omega$ given by Lemma 4. Then define the mappings $J_{1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ and $J_{1}^{\perp}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ given by $J_{1}(x)=\mathrm{span}\{v_{1}(x)\}$ and $J_{1}^{\perp}(x)=(J_{1}(x))^{\perp},$ respectively. By considering the field of Gaussian rationals $\mathbb{Q}(i)=\{p+iq\colon p,q\in\mathbb{Q}\},$ which is a dense set of $\mathbb{C},$ we get the sequence of selection functions $\{\lambda v_{1}(x)\}_{\lambda\in\mathbb{Q}(i)}$ for $J_{1},$ yielding immediately that the latter is measurable. A standard argument using the characterisations of measurability of closed set valued functions (see [Aubin_2009, Theorem 8.1.4]) shows that $J_{1}^{\perp}$ is also measurable.
Assume now that $v_{1},\ldots,v_{k}$ have already been defined for some $1\leq k<d.$ As before, define the mappings
$J_{k}(x)=\mathrm{span}\{v_{1}(x),\ldots,v_{k}(x)\},\qquad J_{k}^{\perp}(x)=(J_{k}(x))^{\perp},$
which are again measurable by the same standard arguments. Next, define the measurable mapping $G_{k}\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ taking values on the set of complex ellipsoids given by $G_{k}(x)=G(x)\cap J_{k}^{\perp}(x)$ (it might happen that $G_{k}(x)$ is the trivial ellipsoid). Applying Lemma 4, one gets a new measurable vector valued function $v_{k+1}$ such that it is orthogonal to $v_{1},\ldots,v_{k}.$
The vectors $v_{1}(x),\ldots,v_{d}(x)$ constructed in this way are the semiaxes of an ellipsoid, each one taken to be maximal in $G(x)$ at each step and given in decreasing order. Thus, these vectors are precisely the semiaxes of $G(x)$ and, if we define the measurable matrix valued function $W(x)$ whose columns are the previous vectors, we get the desired equality $G(x)=W(x)\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
The proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.7] follows by replacing $\mathbb{R}^{d}$ by $\mathbb{C}^{d}.$ Also, the real inner product and associated notions are replaced by the Hermitian product and respective notions. We include the details for completeness.
Proof of Theorem 3.7 in [Cruz_Uribe_Bownik_Extrapolation].
Note that if $F(x)$ is a convex body in $\mathbb{C}^{d}$ with nonempty interior for every $x\in\Omega,$ then Lemmata 17 and 19 yield the result. If that is not the case, one needs to divide $\Omega$ into sets $\Omega_{0},\ldots,\Omega_{d}$ such that $F(x)$ is contained in a linear subspace of dimension $k$ for every $x\in\Omega_{k}.$
Given the measurable mapping $F\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}),$ define $J\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ by
$J(x)=\mathrm{span}F(x)=\overline{\mathrm{conv}}\left(\bigcup_{\lambda\in\mathbb{Q}(i)}\lambda F(x)\right),$
which is measurable by a standard argument using the density of the Gaussian rationals in $\mathbb{C}.$ Also, the mapping $P\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ with $P(x)$ being the orthogonal projection from $\mathbb{C}^{d}$ to $J(x)$ at every $x\in\Omega$ is measurable as well (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6], which also holds in the complex setting by the previous discussion). Then, the sets
$\Omega_{k}=\{x\in\Omega\colon\mathrm{rank}(P(x))=k\},\qquad 0\leq k\leq d,$
are also measurable since the rank can be computed taking the determinants of minors. Now one only needs to construct the restrictions $W|_{\Omega_{k}}$ satisfying the conclusions of the theorem to get the desired function.
Let us fix $0\leq k\leq d.$ One can find measurable functions $v_{1},\ldots,v_{k}\colon\Omega_{k}\rightarrow\mathbb{C}^{d}$ such that $v_{1}(x),\ldots,v_{k}(x)$ are an orthonormal basis of $J(x)$ for every $x\in\Omega_{k}$ (see [Helson_Spectral_Theorem_1986, Theorem 2 in Section 1.3], which applies to complex Hilbert spaces). Denote the set of $s\times t$-matrices by $\mathrm{M}_{{s\times t}}({\mathbb{C}}).$ The matrix $M_{k}(x)\in\mathrm{M}_{{d\times k}}({\mathbb{C}})$ whose columns are $v_{1}(x),\ldots,v_{k}(x)$ is an isometry of $\mathbb{C}^{k}$ onto $J(x)$ for every $x\in\Omega_{k}$ with inverse given by $M_{k}^{\ast}(x)=(M_{k}(x))^{\ast}$ its conjugate transpose. Thus, we can consider the measurable function $F_{k}\colon\Omega_{k}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{k})$ given by $F_{k}(x)=M_{k}^{\ast}(x)F(x),$ which satisfies that $F_{k}(x)$ is a convex body with nonempty interior in $\mathbb{C}^{k}.$ Hence, Lemmata 17 and 19 yield a matrix valued function $W_{k}\colon\Omega_{k}\rightarrow\mathrm{M}_{{k}}({\mathbb{C}})$ such that its columns are mutually orthogonal and with
$W_{k}(x)\mathbf{\overline{B}}_{\mathbb{C}^{{k}}}\subseteq F_{k}(x)\subseteq\sqrt{d}W_{k}(x)\mathbf{\overline{B}}_{\mathbb{C}^{{k}}}$
for every $x\in\Omega_{k}.$
The actual mapping $W$ restricted to $\Omega_{k}$ is obtained as follows. Take $P_{k}$ the coordinate projection from $\mathbb{C}^{d}$ to $\mathbb{C}^{k},$ so $P_{k}\in\mathrm{M}_{{d}}({\mathbb{C}}).$ Then define $W(x)=M_{k}(x)W_{k}(x)P_{k}(x)$ for $x\in\Omega_{k}.$ This mapping is measurable and its columns are mutually orthogonal by construction. It is easy to check that it also satisfies the desired inclusions when applied to $\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
We now turn our attention to the second part of [Cruz_Uribe_Bownik_Extrapolation, Section 3], which concerns integrals of convex-set valued maps. The definitions of the Aumann integral and integrable bounded functions there are valid with $\mathbb{R}^{d}$ replaced by $\mathbb{R}^{2d}$ and thus also by $\mathbb{C}^{d}$.
Lemma 3.13 in [Cruz_Uribe_Bownik_Extrapolation] yields, for a given measurable real vector valued function $f,$ a measurable real convex body valued function $F$ with $|F(x)|=|f(x)|$ at every $x\in\Omega.$ For the complex version, we consider a measurable complex vector valued function $f$ and we construct the corresponding measurable complex convex body valued function $F$ with $|F(x)|=|f(x)|$ at every point in the domain. This is achieved performing one major change that is a recurring theme in the passage from the real to the complex case. Namely, given $f\in\mathrm{L}^{1}(\Omega,\mathbb{C}^{d})$, one defines the convex-set valued map $F$ by
$F(x):=\{zf(x):\leavevmode\nobreak\ z\in\mathbb{C}\text{ with }|z|\leq 1\},\quad x\in\Omega.$
This map is measurable just as in the real case in [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.13] because also the closed unit disk in the complex plane has a dense countable subset.
The validity of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.14], which states that the Aumann integral of a closed set valued function with respect to a nonatomic measure is a convex (not necessarily closed) set, is obvious in the complex case, because this theorem concerns only the topological and real vector space $\mathbb{R}^{d}$, remaining valid for $\mathbb{R}^{2d}$ and thus also $\mathbb{C}^{d}$.
To obtain a complex version of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.15], which states that the closure of the Aumann integral of a convex body valued function is also a convex body, we need once again to directly replace $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$. The notion of real symmetric sets is accordingly replaced by the notion of complex symmetric sets. The Dunford–Pettis theorem and Mazur’s lemma hold equally well with the complex numbers as the underlying field of scalars.
The statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.16] remain true verbatim in the complex case with $\mathbb{R}^{d}$ replaced by $\mathbb{C}^{d}$, because as mentioned Mazur’s lemma remains true in complex Banach spaces.
It is obvious that [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.17, Corollary 3.18], which deal with linearity and monotonicity of Aumann integrals of convex body valued functions, and their proofs remain both true verbatim in the complex case with $\mathbb{R}^{d}$ replaced by $\mathbb{C}^{d}$.
Observe again that the definition of the Aumann integral makes use only of the topological space and real vector space structures of $\mathbb{R}^{d}$. Note also that complex symmetric sets are in particular real symmetric sets. Since [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.19], which states that the Aumann integral of a convex body valued function $F$ equals $\{0\}$ only if $F(x)=0$ almost everywhere, holds equally well for $\mathbb{R}^{2d}$, we deduce that it remains true also for $\mathbb{C}^{d}$.
Finally, in the complex versions of [Cruz_Uribe_Bownik_Extrapolation, Proposition 3.20, Proposition 3.21], which are respectively analogues of Hölder’s inequality and Minkowski’s inequality for Aumann integrals of convex body valued functions, one replaces $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and then considers complex-symmetric sets instead of real-symmetric ones and complex homogeneous norms on $\mathbb{C}^{d}$ instead of real homogeneous norms on $\mathbb{R}^{d}$. The statements and proofs remain in both cases otherwise completely unchanged, because complex symmetric sets are in particular real symmetric sets, complex homogeneous norms are in particular real homogeneous, and as remarked above, [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.17] extends obviously to the complex case.
A.3 Seminorm functions
In this last section, we cover the adaptation of [Cruz_Uribe_Bownik_Extrapolation, Section 4] to the complex setting. As mentioned in Subsection A.1, the definition of seminorm functions itself only requires substituting the vector space $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and real homogeneity by complex homogeneity.
For the statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 4.2], that gives a one-to-one correspondence between seminorm functions and convex body valued functions, one only needs the obvious modifications. This is, one needs to substitute $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ consider the Borel $\sigma$-algebra $\mathcal{B}$ on $\mathbb{C}^{d},$ use countable dense sets of $\mathbb{C}^{d}$ such as $\mathbb{Q}(i)^{d}$ and use the complex version of Hahn-Banach Theorem.
Regarding [Cruz_Uribe_Bownik_Extrapolation, Lemma 4.4], which gives that every measurable convex body valued mapping is the pointwise limit of simple measurable mappings (in the Hausdorff topology), both its statement and its proof follow verbatim under the substitution of $\mathbb{R}^{d}$ by $\mathbb{C}^{d}.$ Note that the characterization of the convergence of convex bodies used in [Cruz_Uribe_Bownik_Extrapolation] can be used in the same way in our context because it actually applies to general nonempty compact convex sets (see [Schneider_Convex_Bodie_1993, Theorem 1.8.7]).
To get [Cruz_Uribe_Bownik_Extrapolation, Theorem 4.5], which characterises integrably boundedness of a measurable convex body valued function in terms of the corresponding seminorm function, one needs to follow the same argument replacing $\mathbb{R}^{d}$ by $\mathbb{C}^{d}$ and keeping in mind that we have changed the definition of function $h_{F(x)}(v)$ to
$h_{F(x)}(v)\coloneq\sup_{w\in F(x)}|\langle v,w\rangle|$
Next, all results about $\mathrm{L}^{p}$ spaces of convex-set valued functions follow with the replacement of $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ since these are based on the topology given by the Hausdorff metric.
Finally, all considerations about matrix weights and seminorms also hold. Here, in addition to the substitution of $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ it is also necessary to replace orthogonal matrices by unitary ones and symmetric matrices by those that are hermitian.
A.4 Main results including the Extrapolation Theorem with matrix weights
In the previous subsections we have explained how the theory developed to deal with real valued matrix weights in [Cruz_Uribe_Bownik_Extrapolation, Sections 2–4] can be modified to be applied to complex valued matrix weights. Once these tools have been conveniently adapted, they can be used to get the complex version of the main results in that article without further changes. In other words, the exposition of [Cruz_Uribe_Bownik_Extrapolation, Sections 5–9] holds for the complex setting by using the results explained in the current appendix. In particular, Theorem A holds.
[…]
Spyridon Kakaroumpas
Würzburg Mathematics Center of Communication and Interaction,
Julius-Maximilians-Universität Würzburg,
Campus Hubland Nord,
Emil-Fischer-Straße 41,
97074 Würzburg, Germany
E-mail address:
spyridon.kakaroumpas@uni-wuerzburg.de
Odí Soler i Gibert
Universitat Politècnica de Catalunya - BarcelonaTech (UPC),
Pavelló I,
Diagonal 647,
08028 Barcelona, Catalunya
E-mail address:
odi.soler@upc.edu